A Complete Bibliography of Publications in Foundations of Chemistry

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

17 October 2024 Version 1.29

Title word cross-reference

#504 [145].

2n^{2} [19]. $30 [154]. 3d [401, 545]. 4 [412, 505, 504]. 4s [401, 545]. _{2} [610, 137, 254]. _{6} [610]. _{N} [278]. d [150]. n [336]. n + 10 [336].

-Shells [150].

//doi.org/10.1007/s10698 [557]. //doi.org/10.1007/s10698-021-09398-4 [557].

0 [145]. 0-465-02775-X [154]. 0-7388-4196-X [155].

1 [278]. 18th [594, 593]. 1914 [33]. 1929 [263, 67].

2 [145, 278]. 2-922686-00-0 [145]. 2007 [214, 221, 228]. 2008 [250]. 2012 [298, 299]. 2015 [468]. 2020 [548]. 2022 [625, 586]. 21st [179]. 2nd [304, 303, 337, 536].

3 [539]. 3-center-2-electron [610]. 30 [372]. 30-second [678, 679, 680, 694, 372]. 35 [325]. 3rd [323].

4 [557]. 41 [362]. 4E7 [145].

5 [146, 324].

6 [613, 614].

7P4 [145].

90 [146]. 90-5867-063-5 [146]. 978 [613, 614]. 978-1-5011-8130-6 [613, 614].

A. [67]. above [354]. absence [579]. absorption [268]. accommodation [524]. account [483, 202, 517]. acid [639]. Acidity [638, 640, 225]. acids [588, 611, 642]. actinium [375]. active [232]. Activity [150]. actual [592]. Ada [535]. additional [584]. Adequacy [369]. Ado [280, 673]. Advances [388]. advent [226]. adventure [287, 676]. Aesthetic [163]. Aesthetics [101]. aether [216]. Affinity [120, 630]. affordance [497]. after [499, 202]. against [386, 364]. Age [90, 652]. agent [283]. agriculture [585]. Ahmed [412]. air [501]. air-related [501]. airs [479]. Alain [145, 337]. Alan [271]. Albert [155]. alchemist [628]. Alchemy [145, 659]. Alcindo [374]. aldohexoses [446]. Alexander [304]. Alexei [304]. algorithmic [543]. algorithms [578]. alkali [364]. all-in-one [639]. allegory [447]. Allen [177]. along [551]. alternative [438]. America [265]. among [285]. Ana [403]. analysis [494, 646, 407, 458, 546]. Anastasios [127]. ancient [391]. Andrea [506]. anions [575]. Anne [287]. Anne-Marie [287]. answer [461]. antiquity [594, 593]. Antonio [496]. any [401]. apart [503]. appear [485]. application [45]. applications [198, 219, 536, 388]. appraisal [578]. approach [589, 603, 508, 438, 427]. approximation [260]. arcane [441]. Architect [97, 687]. Argon [50]. arguing [197]. argument [364]. arguments [365, 354]. arithmetic [608]. art [431, 226]. Artist [97, 687]. aspect [311]. Aspects [192, 340, 258]. assessment [321]. assignment [563]. associated [546]. Athel [13]. Atkins [20]. Atom [216, 667, 21, 88, 469, 271, 351, 650, 401, 344, 632, 556]. Atomes [261]. Atomic [596, 19, 145, 32, 659, 536, 513, 632, 458, 573, 448]. atomic-number [448]. Atomistic [32]. Atoms [475, 341, 349, 667, 21, 150, 548, 340, 500, 232, 352, 271]. atop [364]. Audiences [67, 686]. Aufbau [355]. Austere [345, 346]. Author [82, 105, 130, 156]. Authors [22, 40, 61, 48, 70, 84, 92, 107, 132, 158, 8]. authorship [631]. autonomic [411]. Autonomy [142, 165, 335, 275]. Ave [630]. Avogadro [320]. Axel [21].

B [682, 305, 325, 179, 67, 610]. back [195]. Banchetti [548]. Banchetti-Robino [548]. base [639]. based [448]. bases [611]. Basic [154, 446, 316, 532]. Basis [124, 345, 333, 346]. Basolo [128]. Bassam [324]. Be [354, 342, 466]. beer [683]. before [596, 499]. beginnings [494]. beings [427]. Belousov [44]. Benfey [97]. Bensaude [296, 67]. Bensaude-Vincent [296, 67]. Bernadette [296]. Bernard [21]. Berthollet [624]. Berzelian [43]. Better [124, 461]. Between [146, 17, 168, 282, 259, 520, 425, 431, 613, 614, 72, 500, 586, 451, 592, 535, 654]. Beyond [670, 343, 461, 449, 337]. BHC [454]. Bhushan [66]. biao [680]. Bibilography [53]. bibliometric [646]. Bifurcation [638, 626]. Bifurcations [641]. binary [604, 482]. Binódic [508]. Biochemical [13, 683]. Biochemistry [113, 72, 18, 589, 603, 377]. Biological [44]. Biology [72, 495, 356]. biosignaling [589, 603]. birth [630]. Blind [57]. blocks [369, 411]. boats [461]. Bob [305]. Bohmian [425]. boiling [402]. Bombs [145, 659]. Bond [635, 610, 431, 267, 185, 217]. bond- [185]. Bonding [81, 610, 656, 212, 607]. bonds [349, 481]. Book [103, 428, 81, 178, 237, 126, 271, 21, 154, 186, 613, 146, 280, 90, 128, 145, 304, 296, 305, 297, 303, 324, 337, 338, 323, 325, 361, 371, 372, 370, 179, 239, 581, 270, 121, 403, 76, 13, 238, 33, 104, 66, 180, 380, 67, 155, 263, 560, 29, 97, 193, 287, 122, 536, 549, 187, 262, 261, 439]. Books [154, 208]. Bookstore [145]. bottle [683]. boundary [259]. bounded [341]. Bowden [13]. Boyle [548, 310, 469, 590, 496]. Brad [625]. Brakel [146, 239, 308]. branch [379]. Brazilian [374]. Breaking [197]. Brenner [127]. Brick [571]. bridge [500]. brief [557]. Brock [237]. Buchner [13, 683]. build [461]. building [567]. Burns [682, 687, 97]. butterfly [462].

Cabral [374]. Caldin [409, 417]. Call [7, 34, 68, 69, 118, 36]. caloric [284]. came [573]. can [342]. Can.$70.00 [145]. Canada [145]. Canham [549]. cannot [466]. Carbon [74]. carbonyls [333]. Carl [90]. Case [74, 174, 471, 589, 603, 579, 620, 521, 599, 80, 622, 206, 623, 586, 245, 631]. Catalytic [160]. Cathy [103]. cations [575]. Causal [251]. causality [283]. cause [587]. cause-effect [587]. Celebrating [630]. cellular [601]. Centenary [638]. center [610]. centers [446]. Centuries [21, 606, 553]. Century [79, 282, 216, 462, 594, 593, 605, 359, 329, 196, 631, 43, 179, 297]. Chain [367]. Challenge [177]. challenges [212, 343, 376]. Chalmers [271]. Chang [588, 611]. change [462]. changes [644]. changing [430]. chapter [475]. characteristics [378, 575]. charge [592]. Charles [536, 241]. Charley [55]. Chemical [321, 96, 126, 430, 656, 135, 203, 293, 153, 142, 26, 622, 127, 205, 97, 17, 53, 133, 191, 162, 648, 173, 137, 39, 427, 587, 169, 208, 579, 310, 408, 469, 590, 382, 368, 349, 292, 548, 433, 465, 608, 311, 197, 415, 322, 251, 269, 525, 602, 267, 212, 209, 599, 80, 323, 276, 284, 239, 505, 504, 481, 165, 331, 248, 474, 594, 593, 605, 606, 383, 359, 555, 94, 662, 671, 564, 483, 189, 95, 630, 379, 199, 689, 247, 470, 607, 185, 482, 512, 226, 217, 503, 546, 643]. Chemical [81, 324]. Chemie [29]. Chemist [33, 330, 684, 222]. chemistries [543]. Chemistry [31, 265, 298, 152, 36, 665, 428, 74, 443, 124, 538, 527, 569, 146, 98, 110, 72, 128, 145, 151, 38, 43, 37, 75, 78, 357, 26, 5, 33, 454, 409, 417, 532, 67, 557, 2, 424, 57, 87, 286, 519, 163, 55, 162, 490, 88, 16, 102, 86, 144, 407, 112, 168, 111, 169, 10, 253, 471, 559, 651, 475, 266, 282, 657, 685, 567, 456, 394, 633, 596, 465, 658, 436, 422, 528, 499, 520, 547, 229, 342, 510, 634, 574, 419, 423, 425, 461, 431, 494, 692, 511, 322, 620, 210]. chemistry [530, 659, 231, 283, 577, 531, 236, 345, 366, 400, 80, 688, 669, 290, 684, 626, 291, 246, 260, 312, 677, 410, 403, 365, 686, 500, 383, 184, 480, 618, 268, 380, 206, 576, 261, 487, 553, 472, 495, 566, 332, 672, 230, 308, 623, 116, 690, 449, 612, 438, 346, 648, 585, 578, 175, 619, 344, 516, 534, 512, 388, 611, 301, 275, 356, 451, 645, 458, 631, 535, 376, 654, 47, 183, 296, 551, 155, 303, 324, 103, 625, 66, 338, 403, 238, 569]. Chemists [6, 12, 395, 248, 560, 329, 573]. Chief [325]. chimica [661]. chimie [653, 76]. chiral [446, 385]. chirality [385, 232]. chlorine [521]. choice [456]. choices [624]. Christie [104]. Christies [144]. Chromium [174]. chymical [469, 548]. Circulation [15]. Cité [145]. civilization [659, 145]. Clashing [645]. class [634]. classical [290]. Classification [137, 39, 374]. classifications [313]. classify [573]. classifying [291]. clear [256]. Clericuzio [496]. Climate [462]. Closed [150]. closer [297]. closure [183]. Clothes [171]. C’Nano [393]. Co [631, 254]. Co-authorship [631]. coauthors [451]. Cobb [103]. Coello [128, 657]. cognitive [395, 577]. cold [613, 614]. collaboration [596]. Collected [690, 412, 238]. collection [682, 560]. collective [393]. colliding [639]. colors [530]. colourful [431]. column [364]. Combination [126, 127]. Comment [164, 346, 506, 529, 557, 274, 451]. Commentary [177, 117, 607]. Commercialization [237, 666]. Common [112, 111, 624]. Communicating [686, 67]. communities [602]. community [199]. comparison [456, 356]. complex [481, 604]. Complexity [490, 102, 617, 427]. composition [479, 540, 539]. compound [292]. compounds [604, 402, 482]. Computational [38, 419]. Computer [377, 472]. computers [578]. comunità [661]. Concept [88, 169, 579, 292, 574, 639, 640, 383, 497, 94, 623, 185, 222, 451]. Conception [153, 212, 607]. Concepts [15, 191, 608, 511, 251, 677, 350, 95, 512, 380]. Conceptual [553, 512, 574, 494, 495]. Concerning [316, 149, 161, 446, 407]. concrete [185]. conditions [377]. conduit [459]. conference [299]. configuration [446, 544, 623]. configurations [565]. Confirmation [518]. confusion [512]. Connecting [410, 644, 586]. connection [454]. Connections [168, 574]. conservation [474]. constant [533]. Constitutive [101]. construct [344]. construction [643]. constructionist [517]. constructive [522]. Constructivism [170, 168, 169, 322, 171]. Content [423]. Contents [41, 62, 83, 106, 131, 157, 23]. context [197, 224, 424].

Contingent [171]. continuity [465]. contribute [294]. Contributes [52]. contribution [393, 374]. controversy [202, 624]. conundrum [545]. conventions [598]. Cornish [13]. Cornish-Bowden [13]. Corporation [155]. corpuscles [310]. Correction [603, 614, 467, 505, 594, 540, 359]. correlated [608]. Corrigendum [123, 129]. counterarguments [365]. Courses [124, 383]. created [579]. creative [443]. credit [498]. critical [335]. Critique [20, 162, 625, 313, 607]. Crookes [237, 666]. crystal [547, 535]. crystalline [486]. crystallography [529]. cubic [505, 504]. current [343]. Curriculum [88, 512]. Cusco [299].

D [154, 505, 412, 504]. D-cubic [505, 504]. D. [179, 33]. Dalton [624]. David [280]. deal [489]. dealing [392]. Debate [144]. Deciphering [636]. decisive [610]. decoherence [418]. deductions [533]. Deductive [204]. Deductive-Nomological [204]. deep [518]. defense [621]. define [355]. definition [520]. Deliberating [393]. Deltete [127]. demonstrations [324, 648]. densities [656, 81]. Density [564, 599]. Department [183]. depend [283]. dependent [276]. Describe [11]. describing [493]. description [318]. Design [79, 423, 317]. designing [600]. determination [235]. determining [546]. Developing [73, 497]. development [440, 246, 225]. Diagnostics [419]. Dialectics [400, 4]. did [329]. Dietmar [338]. difference [613, 614]. different [503, 624]. differential [644]. difficulties [546]. diffraction [668, 305]. dimensional [396, 668, 305]. Dioxin [113]. Direct [25]. disagreements [407]. disappearing [674, 270]. Discipline [36]. discourses [269]. discoveries [358, 459]. Discovery [224, 100, 86, 498, 279, 486, 503, 595]. Discussion [149]. Disparities [574]. display [248]. dispositional [590]. dissemination [643]. Distinctness [2]. distributed [145]. division [485]. Djerassi [90]. Dmitri [258]. Dmitrii [154, 663, 196]. Do [283, 278, 255, 618, 395]. Does [485, 437, 351, 57, 294]. doi.org [557]. doing [634]. domain [197]. domain-specificity [197]. Donder [644, 630]. dong [680]. don’t [489]. doubled [485]. Dr. [625]. Draper [64]. driven [601]. Duhem [126]. During [73, 553]. dynamics [538, 527].

each [372, 678, 679, 694]. Early [528, 43, 454, 268, 533]. earth [268]. economy [670, 337]. ed [372, 370, 13, 380, 263, 91]. Edited [126]. edition [337, 536]. Editor [325, 360, 352, 616, 27, 185]. Editor-in-Chief [325]. Editorial [421, 327, 1, 9, 14, 24, 30, 35, 49, 42, 54, 63, 71, 77, 85, 93, 99, 108, 119, 125, 133, 141, 147, 159, 166, 172, 181, 188, 194, 200, 207, 213, 220, 227, 234, 240, 249, 257, 264, 272, 281, 288, 300, 306, 314, 326, 339, 348, 353, 363, 373, 381, 389, 397, 405, 413, 420, 429, 434, 442, 450, 455, 484, 463, 476, 491, 502, 514, 526, 537, 550, 561, 572, 582, 591, 615, 627, 637, 274, 435, 289, 468, 515, 390, 250, 214, 221, 228, 362]. Editors [179]. edn [304, 303, 323]. Eds [303, 337, 338, 239, 33, 66, 67, 560, 97]. Eduard [683, 13]. Education [169, 559, 567, 197, 322, 569, 80, 470]. educators [560]. Edward [297, 409, 417]. effect [462, 587]. eighteenth [462]. Einstein [613, 614]. Eka [415]. Eka-elements [415]. Elaborating [153]. electrolyte [376]. Electron [81, 11, 610, 656, 599, 544]. Electronegativity [426, 592, 471, 316, 513, 632]. Electronic [20, 350, 333, 623]. Electronism [638]. element [229, 521, 94, 560, 294]. elemental [316]. elementary [283]. Elements [120, 510, 136, 276, 149, 53, 49, 191, 39, 208, 579, 609, 415, 411, 485, 372, 674, 504, 481, 541, 331, 248, 594, 593, 605, 606, 268, 202, 285, 647, 678, 679, 680, 681, 694, 173, 482, 503, 573, 457, 505, 371, 270]. Elena [560]. Ellen [287, 676]. else [278]. elucidation [446]. emancipation [301]. Embodiment [78]. Emergence [312, 78, 5, 223, 384, 500, 622, 548]. emerging [416]. Émile [474]. Empirical [101, 215, 472, 624]. Encapsulating [25]. end [594, 593]. energetic [198, 219]. Energetics [182, 198, 219]. energies [635]. engagement [586]. enlarged [337]. entangled [511]. entities [599]. Entropic [350]. Entropy [598, 562, 553]. entry [436]. Environmental [357]. Enzyme [160]. Epistemological [88, 168, 94, 483]. epistemology [331, 645]. equation [636, 644, 533]. equilibrium [555, 487, 483, 533, 546]. equivalent [478]. era [518]. Eric [117, 361, 371, 372, 238, 560, 193, 570, 611, 262]. Erling [370]. Ernst [170, 211]. Erratum [445, 404, 362]. errors [459]. Essays [127, 263, 560, 126]. Essences [135]. essential [229]. essentialism [621]. Essentials [382]. Ethics [37]. Eugene [45]. Europe [33, 684]. evaluating [472]. everywhere [649, 675, 304]. evolution [633, 551, 223]. evolving [596]. Examples [18, 525]. exemplifying [512]. Exeter [183]. exist [255]. existence [266, 392]. expectations [474]. Experimental [38, 474]. Experimental/Theoretical/Computational [38]. Experiments [672, 377, 175, 338]. Explain [115]. Explained [117, 372, 678, 679, 694]. Explaining [247, 143, 544]. explains [613, 614]. Explanation [204, 113, 236, 191, 112, 143, 111, 469, 640, 210, 283]. Explanations [51, 349, 556]. Explanatory [112, 111]. extensions [553]. extent [383]. externalism [322]. Extractive [145].

faces [342, 426]. facie} [618]. Factory [428]. Failed [667, 271]. failure [464]. Fathi [145]. February [414]. ferment [408]. fifth [433]. filling [341]. fills [401]. First [79, 51, 143, 310, 364, 606, 383, 354, 250, 401]. first-order [310]. first-year [383]. Fisher [280]. fission [279]. Fit [50]. Five [470]. Flores [374]. following [633]. force [276, 232]. form [316]. formal [494, 583, 592]. formation [236]. formats [283]. Formula [19]. formulae [333]. Formulas [43, 210]. formulation [369]. Fortin [451]. fortunate [258]. foundation [483]. Foundations [31, 170, 557, 376, 487, 624, 625]. fourth [465]. Foy [145]. Frantisek [263]. Fred [128]. Fridge [613, 614]. Friend [616]. Fukui [564]. function [310, 469, 225, 517]. functional [564]. functions [564]. Fundamental [215, 88, 608, 369, 411, 233]. Further [486, 354]. future [549]. fuzzy [185].

G [337]. G. [538, 527]. G1K [145]. G1X [145]. Gain [667, 271]. Galen [651, 155]. gallium [252]. game [594, 593, 605, 606]. gao [680]. Gas [56, 337, 670]. gases [673, 196, 280]. Gavroglu [403]. Gene [4]. general [574, 383]. generalized [529]. genesis [45]. genius [241]. Geoff [549]. Geoffrey [545]. geometric [608]. Geometry [81, 656]. George [337, 268]. Gestalt [163]. Getting [478]. Ghibaudi [560]. Giant [649, 675, 304]. Gibbs [636]. Gillespie [81]. Giovanni [490]. given [259]. Glasersfeld [170, 211]. Glauber [496, 628]. Gleditsch [287, 676]. glossaries [574]. glue [253]. Goeppert [337]. goods [430]. Gordin [154]. Gorter [535]. GPCRs [589, 603]. grammars [269]. graphic [411]. Grapí [501]. Greek [640]. Green [151, 626, 357, 454, 338, 246, 410, 672, 585, 551]. Greene [323]. Grosberg [304]. grounds [600]. Group [336, 540, 45, 242, 539]. Growth [13, 683]. Guest [421, 435, 289, 468, 515, 390]. guide [355, 323, 662, 671].

H [237, 610, 412, 179, 33, 137]. Habashi [145]. habit [568]. hadronic [388]. half [372, 678, 679, 694]. handbook [324, 648]. handle [436]. happened [573]. hardbound [145]. Hargittai [297]. harmonic [608]. hasn’t [117]. Hasok [588, 611]. Head [55]. heat [398, 630, 439]. Hegel [59]. Heisenberg [267]. helium [457, 629]. helix [492]. Helmont [408]. Help [191]. helps [311]. her [287, 676]. Herbert [338]. here [649, 675, 304]. Herman [535]. Hevesy [268]. high [633]. Highes [31]. Hints [583]. historians [560]. historical [633, 184, 483]. historical/epistemological [483]. histories [320]. History [103, 428, 21, 174, 280, 145, 65, 26, 33, 97, 465, 658, 422, 378, 311, 279, 673, 461, 692, 659, 674, 684, 626, 270, 501, 403, 594, 593, 605, 606, 454, 487, 553, 650, 449, 254, 441]. honour [682]. horizons [229]. hot [613, 614]. https [557]. hua [680]. Human [21, 650]. humanities [586]. Hund [355]. Hurmuzescu [302]. hybrid [408]. Hydrogen [161, 629, 143, 364, 190, 460]. Hypotheses [32].

Ibuprofen [454]. idealisations [493]. Idealization [169]. ideas [483, 470]. Identity [137, 620]. IdF [393]. II [398, 219, 527, 364]. Ilya [176]. Image [146, 163, 654]. imaging [273]. Immaculate [90, 652]. impact [268]. Implementation [218]. implications [357, 340]. Imponderable [64]. impossibility [334]. impossible [367]. Improving [559, 577]. impure [665, 296]. incompatibilities [662, 671, 323]. Incompatible [471]. incomplete [546]. inconsistencies [639, 318]. independent [459]. Index [40, 61, 82, 105, 130, 156, 22, 513]. individuals [436, 319]. Industrial [454]. inertness [385]. Influence [211, 408]. Influences [176]. influential [297]. Ingold [31]. initio [116]. inner [244]. innovation [218]. Innovative [151]. Inorganic [128, 4, 657, 402, 325]. inquiry [290]. insider [183]. inspired [583, 585]. Inspiring [501]. Instructions [8, 48, 70, 84, 92, 107, 132, 158]. Instrument [17]. Instrumental [98]. instruments [633]. insularity [461, 449]. integrated [551]. integrity [525]. Interaction [643, 587]. intercultural [308]. Interface [72, 535]. Internal [335, 575, 485].

International [47, 299, 265, 298]. Interplay [160]. interpretation [408, 473, 622, 488, 643]. Interpreting [610]. Interpretive [26, 602]. Interview [170, 597, 570]. intimacy [525]. Introducing [633, 399]. Introduction [109, 211, 126, 134, 307, 361]. Introductory [124]. intuitive [618]. invention [454]. inventors [138]. investigate [521]. invisible [600]. ISBN [154, 613, 614, 146, 145, 155]. isms [167]. Isodiagonality [277]. Isoelectronic [233]. isolated [493]. isomeric [446]. Isomerism [418, 473]. isotopes [498, 595]. isotopy [596]. ISPC [298, 468, 250, 214, 221, 228]. Issue [49, 133, 468, 340, 561]. issued [506]. issues [235]. István [297]. Italia [661, 122]. italiana [661]. Itself [112, 111].

J [428, 81, 303, 97]. Jaap [239, 146, 308]. Janet [241]. Jay [586]. Jean [380]. Jean-Pierre [380]. Jeffrey [461]. Joachim [29]. Johann [628]. John [64]. Jonathan [296]. Judging [297]. July [414]. Just [116, 554]. Justification [149].

kainosymmetry [575]. Kaji [414]. Kant [331, 230]. Kantian [645]. Kean [270]. Kennepohl [338]. Kenneth [303]. key [483]. Khokhlov [304]. kind [276]. Kinds [135, 133, 525, 602, 489, 642]. kinetic [533]. kinetics [382]. Kingdom [20]. Klabunde [303]. Klaus [239, 263]. Klein [292, 91]. Knight [33, 189, 177]. knowing [330]. Knowledge [25, 567, 667, 368, 683, 271, 295, 512, 643, 13]. Kostas [403]. Kouso [439]. Kragh [33]. Kubanek [287]. Kuhn [321, 209]. Kuhnian [458].

L [81]. Labarca [335, 164]. Labinger’s [586]. Laboratory [91, 259, 428]. Lagerkvist [370]. Lame [57]. landmark [261]. Language [162, 494, 322, 583, 482]. languages [601]. Lanthanides [148, 444, 445, 467]. lanthanum [375]. Laszlo [76]. Latin [265]. lattice [505, 504]. Laval [145]. Lavoisier [386, 398, 136, 661, 122]. Law [73, 369, 100, 52, 192, 197, 553, 411]. lawrencium [375]. Laws [96, 86]. Layer [655, 104]. layperson [568]. Learning [171, 169, 518, 601]. Lectures [155, 651]. legacy [267]. legitimate [556]. Leibniz [527, 538]. less [244, 287, 676]. Let [461]. Letter [27, 185, 352, 360]. Leuven [146, 146]. level [633]. Lewis [81, 404, 656, 583, 387]. life [529, 287, 676, 258]. Lifetime [128, 657]. light [378, 486, 529, 331, 439]. like [185]. limitations [357]. Limits [102, 293]. linear [446]. lingua [286]. linkage [336]. Linking [366, 365]. Liquid [547, 535]. Literature [586, 535]. living [427]. Llored [380]. location [540, 539]. logic [355]. Logics [543]. Lombardi [366, 335, 164, 597]. look [297]. love [674, 270]. Löwdin [177]. Lundgren [67]. lutetium [375].

M [303, 33]. Mackay [486]. macro [494]. macro-chemistry [494]. macromolecules [245]. macrostructure [517]. madness [674, 270]. Magick [658, 103]. magnetochemistry [302]. Magnum [16]. make [477]. Making [33, 558, 684]. man [258]. Manifest [146, 654]. manifestations [575]. many [309]. Mapping [660, 121]. Marcus [378]. Maria [630]. Marie [287]. Marina [548]. mark [243, 155]. Markovnikov [65]. Marks [557, 334, 557]. Masanori [414]. mass [474]. material [430, 290, 623]. materials [303, 688, 669]. Mathematical [152, 192, 508]. mathematician [634]. Mathematics [38]. mathematize [332]. matrix [258]. Matter [428, 412]. Matters [256]. Maureen [104]. Mavericks [103, 658]. maximum [636]. may [480]. Mayhem [103, 658]. McCaw [536]. me [680]. mean [329]. meaning [529, 636]. means [608, 273]. measurable [273]. measurement [554, 426]. Mechanical [90, 394, 652, 488, 556]. mechanicism [548]. Mechanics [176, 79, 117, 45, 520, 425, 384, 473, 466, 345, 500, 294, 438, 346, 356]. mechanism [282, 278]. mechanisms [556]. Mechanistic [204]. medical [579]. medieval [579]. medium [316]. mei [680]. melting [402]. memorize [395]. memory [414]. Mendeleev [154, 252, 73, 506, 663, 136, 100, 195, 196, 464, 86, 258, 477]. Mendeleiev [364, 354]. Mendeleyev [542, 557]. mereological [344]. Mereologies [269]. Mereology [260, 328]. Mess [58]. Metal [150, 364]. Métallurgie [145]. Metallurgy [145, 659]. Metals [244, 242, 364, 354, 507]. metaphysical [253]. Metaphysics [203, 160]. methanol [337, 670]. Methodism [167]. methodologies [677, 380]. Meyerson [474]. Miao [680]. Michael [154, 212, 268]. Michele [616]. Microstructuralism [245, 525]. microstructure [469]. Minds [66, 685]. Mineral [489]. mineralogical [579]. mineralogists [489]. minute [678, 679, 694, 372]. Miroir [653, 76]. mis [525]. misbehavior [489]. Misconception [90, 623, 652]. missed [370, 693]. misunderstanding [359]. Mixture [126, 56, 127]. Mixtures [140]. Modality [140]. Model [495, 88, 144, 394, 416]. Modeling [44]. modelling [80]. Models [601, 143, 471, 80, 556, 523]. Modern [31, 97, 387, 404, 611]. Modes [91]. mole [315, 320]. Molecular [81, 79, 163, 101, 523, 554, 511, 656, 466, 447, 260, 500, 497, 522, 273, 222]. molecule [437, 493, 344]. Molecules [328, 66, 341, 687, 685, 349, 649, 675, 304, 340, 500, 600, 232, 352, 475, 97]. monograph [462]. moonlighting [245]. moral [410]. Morris [428, 97]. most [297, 372, 678, 679, 694]. Motivations [182, 343]. Move [189]. Moving [449, 461]. Much [673, 57, 280]. multiple [596, 426, 459]. Muspratt [569]. must [470].

Naïve [167]. Nalini [66]. Name [594, 593, 605, 606]. Naming [202, 594, 593, 605, 606]. nan [680]. nano [392]. nanomaterials [436]. Nanoscale [688, 669, 303]. Nanoscience [393]. Nanotechnology [391, 163]. Narrative [120]. Natural [602, 355, 640, 525, 489, 642, 441]. Nature [4, 59, 86, 590, 398, 610, 599, 239, 309, 184, 689, 313, 588, 512, 217]. Need [57, 395, 184, 253]. Needham [126, 165, 32]. Neglect [10]. neglected [311]. Neither [692, 403]. neon [457]. Nernst [639]. network [481, 604, 482]. Neuss [545]. Newlands [248]. ni [680]. Nineteenth [359, 216, 605, 553, 329, 43]. Nineteenth-century [359, 216, 329, 43]. nipponium [563]. nitrate [310]. No [139, 520, 253]. Nobel [370, 693, 196]. Noble [280, 150, 673, 196]. Nomological [204]. Non [584, 162, 266, 341, 384]. non-existence [266]. Non-observables [162]. non-overlapping [341]. Non-periodic [584]. non-reducibility [384]. nonmetals [507]. nor [692, 403]. Norrby [370]. Notation [32]. Note [616, 212, 335]. Nothing [280, 676, 673, 287]. Novel [155, 651]. nowadays [391]. nuclear [596, 279, 509, 232]. nucleophilicity [513]. number [596, 623, 320, 235, 458, 573, 448]. Numbers [11]. Numerical [39].

O [97, 137]. Objects [290, 424]. Observables [162]. Obstacle [88]. Occasion [144]. occur [369, 278]. October [625]. Off [6, 12]. Office [393]. Ogawa [563]. oil [337, 670]. Olah [337]. old [683, 224, 275]. Olimpia [597]. One [330, 64, 11, 639, 297, 401]. One-Electron [11]. only [503]. Ontological [282, 142, 164, 516, 310, 384, 392, 165, 335]. ontologies [284]. ontology [589, 603, 408, 430, 231]. open [341]. Operations [392]. optical [473]. optically [232]. Opus [16, 331]. orbital [260, 273, 545, 451]. orbitals [355, 255, 565, 536]. order [310, 635]. Ordered [154, 663]. Organic [31, 531, 419, 210, 618, 552]. Organising [507]. Organized [167]. orientations [525]. Origin [74, 100, 163, 292, 225, 232]. Origins [182, 246]. orthodox [343]. Orthogonality [565]. oscillating [400]. oscillations [382]. Ostwald [182, 198, 219]. Other [167, 674, 270]. Other-isms [167]. Our [36]. Outline [16]. Overcoming [497]. overlapping [341]. overview [246]. Oxford [548]. Oxidation [235, 592]. Ozone [655, 104].

P [45, 323, 263]. P. [97]. panenmentalist [367]. Paneth [230, 331]. Paola [548]. Paper [117, 43]. Papers [34, 68, 69, 118, 7, 682, 238, 690, 691, 262]. Paradoxes [137]. parallel [266]. part [594, 593, 605, 606, 364, 198, 219]. particle [195]. particular [401]. parts [411]. past [549]. patches [215]. patterns [336]. Paul [81, 126, 613, 614, 32]. Pauli [355]. Peculiar [86]. pedagogic [558]. Pedagogical [168]. Pedagogy [65]. Pegniochemistry [379]. Peirce [55]. Peircean [643]. Perception [163]. Pere [501]. Perfect [56]. period [485]. Periodic [89, 73, 117, 50, 154, 242, 369, 575, 100, 179, 51, 148, 333, 52, 149, 193, 336, 192, 161, 20, 53, 49, 664, 191, 39, 299, 499, 197, 609, 663, 509, 452, 584, 375, 361, 370, 674, 195, 693, 190, 270, 316, 374, 541, 201, 233, 277, 647, 401, 508, 691, 247, 313, 294, 549, 524, 646, 241, 453, 262, 540, 539, 416, 448, 568, 457, 400, 411, 460]. periodicities [584]. Periodicity [317, 333, 447, 485, 584, 575, 248]. periods [369]. Periphery [45]. Persia [579]. persistence [426]. Perspective [655, 425, 427, 643, 104]. Perspectives [66, 685, 645]. Peru [299]. Peter [428]. phenomena [199]. Phenomenological [143]. Philosopher [667, 271]. philosophers [560]. philosophica [286]. Philosophical [124, 176, 340, 66, 600, 133, 252, 685, 496, 486, 601, 629, 407]. Philosophische [29]. Philosophy [47, 265, 298, 36, 667, 655, 74, 6, 12, 525, 72, 80, 5, 104, 59, 308, 18, 16, 102, 10, 654, 559, 310, 469, 548, 436, 271, 461, 677, 410, 238, 380, 230, 690, 449, 611, 441, 146]. phlogiston [386, 398, 440, 284]. Photophysics [174]. Photophysics- [174]. Physical [103, 216, 658, 636, 256, 487]. Physics [52, 57, 191, 475, 596, 425, 692, 366, 195, 403, 365, 184, 268, 553, 175, 451, 253]. Piece [50]. Pierre [76, 380, 126]. pipe [568]. pivotal [451]. place [201, 354]. placeholder [621]. placement [364]. platinum [242, 561]. platinum-group [242]. Plato [528, 576]. pluralism [293, 407]. pluralistic [519]. plus [145]. poetry [547, 609, 535]. Pohanish [323]. points [402]. Polanyi [268]. Policy [113]. Polyelectronic [11]. polymorphism [553]. Popelier [81]. Posidonius [206]. Position [161, 457, 460, 629]. positions [375]. possibilities [367, 358, 415, 486]. possibility [334]. postage [145]. postumum [331]. potassium [310]. potentials [644]. Power [160]. Power- [160]. powerful [383]. Pp [548, 154, 146, 145, 155]. Practical [75, 409, 417]. Practically [280, 673]. practice [406, 525, 602, 175]. practices [368, 677, 380]. Prague [263]. praise [580]. Prakash [337]. Predicting [604]. Prediction [113, 524]. Predictions [89, 252, 506, 464, 477]. prehistory [254]. present [549]. Press [548, 146, 263]. Presuppositions [203]. Priestley [479]. Prigogine [176, 309]. prima [618]. principle [518]. Principles [120, 446, 585]. Prins [535]. priori [474]. Prize [693, 370]. prizes [196]. problem [475, 554, 473, 392, 335, 544, 173, 407]. problematic [512]. Problems [133, 440, 578, 275]. Process [160, 589, 603, 454]. processes [494, 487]. product [430]. production [295]. Professionalism [37]. Professor [414]. Program [60]. Programme [171, 347]. progress [609, 569, 495]. progressing [401]. Progressive [171]. promoting [197]. propensity [619]. properties [252, 590, 565, 632]. property [316, 233]. proposal [633]. Prospective [585]. prospects [343]. protein [517]. proteins [245]. Protonism [638]. Prototypical [174]. Proust [624]. provides [592]. Public [113]. published [625]. Pullman [21]. Pure [358, 415, 486]. purity [295]. put [190]. puzzle [579, 311].

qi [680]. QSAR [518]. QTAIM [340, 347, 343]. qualitative [530]. Quality [454]. Quantitative [575]. Quantum [79, 117, 384, 149, 11, 191, 578, 556, 341, 45, 520, 554, 692, 473, 466, 236, 345, 599, 260, 312, 403, 340, 500, 622, 268, 472, 623, 116, 294, 438, 346, 645, 475]. quantum-chemical [599]. quasicrystals [334]. Québec [145, 145]. quest [452, 432]. question [334, 243, 294]. Questions [191, 168].

R [21, 304, 361, 179, 193, 611, 262]. R. [682]. Radical [170]. Rallying [36]. range [235]. Rapoport [226]. rare [268]. rather [452]. Rauchfuss [325]. ray [668, 305]. Rayner [549]. Rayner-Canham [549]. Rays [64]. Re [472, 321, 436, 563]. re-assessment [321]. re-assignment [563]. re-entry [436]. Re-evaluating [472]. Reaction [44, 383, 546]. Reactions [203, 153, 657, 367, 400, 474, 278, 439, 318, 128]. Reactivity [619, 114, 564, 587]. Reading [25, 569]. Realism [422, 406, 46, 335, 522, 523].

Realismus [29]. really [352]. reappraisal [524]. Reasoning [75]. Reception [100]. recipes [628]. Recommended [191]. reconciled [342]. redintegration [310, 496]. Redox [318]. reduced [466]. reducibility [384]. Reduction [164, 475, 520, 622]. Reductionism [3]. reductive [345, 346]. Reflections [152, 36, 386, 398, 569, 601]. regarding [574]. Reisinger [21]. Related [126, 127, 479, 501, 541, 402]. Relating [632]. Relation [39, 355]. relational [590, 517]. relational-constructionist [517]. relations [482]. Relationship [17, 169, 425, 189, 451, 592]. Relationships [285]. relativity [399, 566]. remembered [682]. remix [492]. Remnants [3]. reply [496, 347, 366]. Reporting [503]. Repositioned [148]. Representation [121, 91, 101, 368, 396, 423, 620, 531, 660, 411, 448, 643]. representations [601, 453]. represented [316]. representing [512]. reproduction [652, 90]. requires [522]. Research [72, 121, 481, 17, 646, 171, 168, 169, 660, 347, 601]. resolution [624]. resonance [267]. respiration [601]. Response [96, 625, 179, 127, 360, 28, 177, 545, 112, 165]. responsibility [393]. Retrodictions [89]. Reveal [112, 111]. revelation [452]. reverse [432]. reversible [487]. Review [103, 428, 81, 548, 237, 126, 271, 21, 154, 186, 146, 280, 90, 128, 145, 304, 296, 305, 297, 303, 324, 337, 338, 323, 325, 361, 371, 372, 370, 239, 581, 270, 121, 403, 76, 13, 238, 33, 104, 66, 180, 380, 67, 155, 263, 560, 29, 97, 193, 287, 122, 536, 549, 187, 262, 178, 609, 613, 623]. Revised [148]. revision [518]. Revisited [135, 417, 557, 248, 542, 409]. Revolution [321, 98, 26, 433, 465, 452, 209, 458]. Revolutionary [535]. Revolutions [612]. rhenium [563]. Richard [323]. Richards [303, 595, 631]. riddle [224]. right [478]. rightful [201]. Rise [31, 302]. Ritter [302]. rivalry [606]. rivoluzione [661]. Road [191]. Robert [548, 310, 590, 682, 687, 127, 97]. Robino [548]. Roesky [338]. Role [38, 162, 143, 247, 493]. Ronald [81]. Roots [454]. Rosenfeld [66]. Routledge [586]. royal [628]. Rudolph [628]. rue [145]. Rule [65, 355]. rule-set [355]. rupture [465]. Ruthenberg [239, 263]. Ryan [303].

S [536, 278]. s10698-021-09398-4 [557]. Sainte [145]. Salt [139]. saltpeter [496]. Sam [270]. Saving [199]. scale [513]. scandium [252, 544]. Scerri [361, 371, 372, 238, 560, 193, 262, 117, 570, 611]. school [633]. Schummer [29]. Science [667, 655, 113, 6, 12, 104, 490, 160, 171, 144, 665, 666, 216, 378, 443, 271, 620, 395, 577, 501, 409, 417, 532, 379, 459, 612, 407, 676, 441, 535, 237, 586, 296, 287]. Sciences [97, 91, 521, 586]. Scientific [620, 146, 46, 191, 86, 520, 609, 358, 199, 175, 407, 523, 654]. scientifica [661]. Scientist [667, 271]. scientists [297, 606, 202]. screening [632]. Scrutinizing [86]. Sea [139]. Search [26]. second [310, 553, 678, 679, 680, 694, 221, 372]. second-order [310]. Secondary [53, 485, 278]. secret [628]. Seeman [461]. Seimika [439]. Selected [691, 262]. semantics [322]. semi [472]. semi-empirical [472]. semiotics [643]. Sen [613, 614]. sense [558]. separating [175]. serendipity [224].

Series [97, 233]. set [355]. seventeenth [282]. Sex [90, 652]. Shadow [154, 663]. Shahbazian [347]. Shakhashiri [324]. shape [511, 351]. Sharp [640]. Shechtman [334]. Shell [19]. Shell-Structure [19]. Shells [150]. shifts [229]. short [361, 320]. should [498]. side [256]. sign [598]. Significance [664, 252, 486, 529, 189, 193]. significant [372, 678, 679, 694]. Signs [112, 111]. Similarity [368, 518]. Simões [403]. Simon [296]. Simple [153, 555]. simulation [578]. simulations [377, 472]. simultaneous [459]. since [569]. single [355]. sixteenth [282]. Skepticism [167, 497]. Sketch [60]. Sketches [18]. Slater [632]. Social [33, 684]. Society [47, 265, 298, 393]. sociology [461, 449]. Solid [153]. solvolysis [278]. Some [176, 203, 20, 378, 608, 358, 313, 185]. something [278]. somewhat [258]. Sonya [226]. Sources [53]. space [341, 396]. space-filling [341]. speaking [482]. Special [340, 566, 49, 133, 468]. specificity [197]. Specimen [17]. spectra [536]. Spectrometer [25]. spectroscopic [633]. spectroscopy [633]. Spectrum [121, 143, 660]. speculations [313]. spherical [453]. spin [276]. spin-dependent [276]. Spinoza [496]. Spiral [411, 647]. spirals [196]. spirit [408]. Spirited [103, 658]. spoon [674, 270]. Stability [114]. standards [596]. Stanley [323]. Start [55]. State [153, 592]. Statement [137]. states [565]. Statistical [176, 488, 553]. Status [169, 94, 343, 646, 516]. step [500]. Stoffen [29]. stoically [206]. stoichiometric [482]. Stone [667, 271]. Story [193, 664]. Strain [115, 114]. Straining [115]. Strategies [26, 112, 111]. strategy [446]. strictly [466]. striking [358]. strong [384]. Structural [210, 469, 212, 607]. Structure [511, 19, 20, 402, 114, 596, 311, 554, 466, 350, 437, 497, 522, 222]. Structure-related [402]. Structured [356]. structures [583]. Stuart [66]. student [577]. Studies [46, 609]. study [383, 623, 427]. Stuff [319, 689]. Stuff. [239]. Subatomic [51]. subsets [485]. subshells [401]. substance [229, 574, 430, 552]. substances [293, 290, 392, 689, 239]. substrates [278]. Succeeded [667, 271]. success [577, 464]. successful [252]. Summer [298, 265]. supervenience [205]. supporting [483]. Supramolecular [163]. Surya [337]. Sustainability [357]. sustainable [338, 626, 672, 585]. Switch [163]. symbolism [226]. Symmetry [396, 385]. Symposium [298, 265]. synergetics [400]. Syntheses [325]. Synthesis [115, 218]. System [50, 51, 53, 49, 191, 39, 640, 508, 241, 453, 356, 448, 411]. Systemic [490, 427]. Systems [87, 11, 44, 341, 385, 95, 617]. systems-theoretical [95].

T [428, 97]. Table [89, 117, 242, 457, 364, 179, 148, 333, 149, 193, 161, 664, 299, 499, 609, 663, 509, 452, 485, 584, 375, 361, 370, 674, 195, 693, 190, 270, 316, 201, 354, 277, 401, 691, 247, 313, 294, 549, 524, 646, 539, 416, 568, 154, 400, 460, 540, 262]. tables [575, 541, 196]. take [206]. tale [371, 681]. tales [674, 270, 495]. tangible [259]. target [512]. taste [640]. teach [480, 545]. teacher [559]. teachers [633, 324, 648]. Teaching [499, 88, 168, 169, 355, 660, 601, 623, 121]. Techniques [121, 660]. Technochemistry [330]. Technology [151]. Teller [297]. Telltale [112, 111]. telluric [492]. temperature [488]. tensions [282]. terms [513]. test [610]. tests [474]. tetrad [584]. Tetrahedral [74, 453]. text [558]. textbook [574]. Textbooks [67, 686, 623]. texts [579]. Their [143, 215, 686, 67, 318, 573]. Theodore [595, 631]. Theoretical [110, 38, 75, 144, 143, 528, 486, 95, 611, 624]. Theories [96, 217, 86, 215, 440, 520, 618]. Theory [31, 475, 456, 87, 160, 386, 45, 198, 219, 610, 231, 236, 385, 340, 500, 350, 359, 387, 404, 399, 268, 564, 483, 566, 329, 459, 482, 611, 645]. there [139, 649, 675, 304, 352]. thermal [487]. Thermodynamic [487, 644, 533]. thermodynamics [382, 553]. Thing [154, 663]. thinking [530]. third [276, 228, 299]. Thomas [209, 325]. Thomistic [622]. thought [538, 527, 424, 650, 175, 21]. Three [541, 396, 334]. till [594, 593]. Timaeus [528]. time [391, 516]. Tithonic [64]. tomographic [273]. Tools [43, 91]. topics [541]. topological [351]. topology [511]. total [218]. transfer [495]. transformations [383]. transition [244, 285]. Translated [32, 126, 21]. translation [439, 432]. treatment [555]. triads [247, 580, 448]. Trichotomy [38]. triggered [569]. Trivial [171]. Trope [231]. true [674, 270, 460, 217]. truth [416]. turbulent [605]. turn [406, 631]. Turned [6, 12]. twentieth [297, 606, 553, 631]. Twenty [79, 606]. twenty-first [606]. Two [668, 500, 520, 342, 485, 485, 305]. Two-dimensional [668, 305]. Two-step [500]. typical [244].

U [91]. U.S.$50.00 [145]. Udagawa [439]. Ulf [370]. uncertainty [392]. uncompensated [630]. underpinnings [521]. Understanding [562, 522, 52, 382, 184, 644]. unfragmented [355]. unified [611]. unifying [438]. universals [173]. universe [609, 613, 614]. universes [266]. Universitaire [145]. University [548, 146, 145, 80, 183]. unknown [604]. unlikely [626]. unrecognized [241]. Unruly [58]. Untersuchungen [29]. update [375]. updated [337]. uranium [243]. Usanovich [639]. Use [11, 525]. Useful [18]. Using [355]. usually [316]. UV [633]. UV-visible [633].

vague [534]. Valencies [444, 445, 467]. valency [610, 387, 404]. variants [485]. variations [369]. vernaculars [251]. version [387, 404]. versus [204, 525, 319, 474]. very [361]. view [183, 481, 604, 95, 320, 313, 619, 568]. Viewing [291]. Views [211, 479]. Vihalemm [96]. viii [145, 242]. Villani [490]. Vincent [296, 67]. visible [633]. vision [551]. Visual [121, 660]. visualization [412, 317]. vitalism [282]. vitro [377]. vivo [377]. vivo-in [377]. voices [309]. Void [21]. Volume [23, 40, 41, 61, 62, 82, 83, 105, 106, 130, 131, 156, 157, 324, 325]. vs [638].

W [538, 527, 338, 631]. Wald [263]. was [258]. waste [430]. Watch [530]. water [479, 137]. Wave [394]. Wave-mechanical [394]. ways [330, 291]. weak [232]. Weidler [287]. Weisberg [313, 607]. Weisberg’s [212]. Well [154, 663]. Well-Ordered [154, 663]. Where [190]. Which [401]. Whom [498]. Wigner [45]. Wiley [323, 662, 671]. Wilhelm [182, 198, 219]. William [64, 666, 237]. Wissenschaft [29]. within [440, 201]. Woodward [97, 682, 687]. Woody [506]. Work [73, 124, 378, 636, 618, 268, 451, 258]. working [503]. works [412]. Workshop [638]. World [142, 97, 58, 687, 674, 270, 165]. worlds [503]. Would [124, 50]. Wray [625]. Writings [20]. wrong [478].

X [154, 305, 155, 548, 668]. X-ray [305, 668]. xiang [680]. xiv [146]. Xlibris [155]. xue [680]. xx [154]. XYZ [138].

year [383]. years [551, 503]. York [154, 586]. Youan [439]. yuan [680].

Z [324]. Zewail [412]. Zhabotinsky [44]. zhou [680]. Zinjafr [579]. Zone [145].

References

Scerri:1999:Ea

[1] Eric R. Scerri. Editorial 1. Foundations of Chemistry, 1(1):1–5, March 1999. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

     (electronic). URL http://link.springer.com/article/10.1023/A:
     1009976108288.

Plesch:1999:DC

[2] Peter H. Plesch. On the distinctness of chemistry. Foundations of Chemistry, 1(1):6–15, March 1999. CODEN FOCHFL. ISSN 1386-4238 (print),

     1572-8463 (electronic). URL http://link.springer.com/article/10.
     1023/A:1009980225127.

Vemulapalli:1999:RR

[3] G. Krishna Vemulapalli and Henry Byerly. Remnants of reductionism. Foundations of Chemistry, 1(1):17–41, March 1999. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

     springer.com/article/10.1023/A:1009984310105.

Mackay:1999:DNI

[4] Alan L. Mackay. From “the dialectics of nature” to the inorganic gene. Foundations of Chemistry, 1(1):43–56, March 1999. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

     springer.com/article/10.1023/A:1009924107380.

McIntyre:1999:EPC

[5] Lee McIntyre. The emergence of the philosophy of chemistry. Foundations of Chemistry, 1(1):57–63, March 1999. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1023/A:1009932309197.

Good:1999:WCTa

[6] Robert J. Good. Why are chemists ‘turned off’ by philosophy of science? Foundations of Chemistry, 1(1):65–95, March 1999. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1023/A:1009931623309.

Anonymous:1999:CP

[7] Anonymous. Call for papers. Foundations of Chemistry, 1(1):97, March 1999. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1023/A:1017230830409.

Anonymous:1999:IA

[8] Anonymous. Instructions for authors. Foundations of Chemistry, 1(1): 99–106, March 1999. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1017259931318.

Scerri:1999:Eb

[9] Eric R. Scerri. Editorial 2. Foundations of Chemistry, 1(2):107– 109, June 1999. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1009917310399.

vanBrakel:1999:NPC

[10] J. van Brakel. On the neglect of the philosophy of chemistry. Foundations of Chemistry, 1(2):111–174, June 1999. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1023/A:1009936404830.

Richman:1999:UOE

[11] Robert M. Richman. The use of one-electron quantum numbers to describe polyelectronic systems. Foundations of Chemistry, 1(2):173– 181, June 1999. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1009988526035.

Good:1999:WCTb

[12] Robert J. Good. Why are chemists ‘turned off’ by philosophy of science? Foundations of Chemistry, 1(2):185–215, June 1999. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1023/A:1009934808874.

Malmstrom:1999:BRA

[13] Bo G. Malmström. Book review: Athel Cornish-Bowden, ed., Eduard Buchner and the Growth of Biochemical Knowledge. Foundations of Chemistry, 1(2):217–219, June 1999. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1023/A:1009998112333.

Scerri:1999:Ec

[14] Eric R. Scerri. Editorial 3. Foundations of Chemistry, 1(3):221–223, October 1999. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1009974614720.

Laszlo:1999:CC

[15] Pierre Laszlo. Circulation of concepts. Foundations of Chemistry, 1(3): 225–238, October 1999. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1009964503013.

VanCik:1999:OMO

[16] Hrvoj VanCik. Opus magnum: An outline for the philosophy of chemistry. Foundations of Chemistry, 1(3):239–254, October 1999. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1023/A:1009989903921.

Rothbart:1999:RBI

[17] Daniel Rothbart. On the relationship between instrument and specimen in chemical research. Foundations of Chemistry, 1(3):255–268, October 1999. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1023/A:1009902306647.

Strand:1999:TUP

[18] Roger Strand. Towards a useful philosophy of biochemistry: Sketches and examples. Foundations of Chemistry, 1(3):269–292, October 1999. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1023/A:1009994005739.

Harcourt:1999:ASS

[19] Richard D. Harcourt. The atomic shell-structure formula 2n^{2}. Foundations of Chemistry, 1(3):293–294, October 1999. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

accesspage/article/10.1023/A:1009906525864.

Scerri:1999:CAP

[20] Eric R. Scerri. A critique of Atkins’ periodic kingdom and some writings on electronic structure. Foundations of Chemistry, 1(3):295–303, October 1999. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1009906513673.

Everdell:1999:BRC

[21] William R. Everdell. Book review: 25 centuries of atoms and void. Pullman, Bernard, The Atom in the History of Human Thought, translated by Axel R. Reisinger. Foundations of Chemistry, 1(3):305–309, October 1999. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1023/A:1009929123489.

Anonymous:1999:AI

[22] Anonymous. Authors index. Foundations of Chemistry, 1(3):311, October 1999. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL http://link.springer.com/accesspage/article/
10.1023/A:1017294916963.

Anonymous:1999:VC

[23] Anonymous. Volume contents. Foundations of Chemistry, 1(3):312– 314, October 1999. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1017201816055.

Scerri:2000:Ea

[24] Eric R. Scerri. Editorial 4. Foundations of Chemistry, 2(1):1–4, ???? 2000. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1023/A:1009984100928.

Baird:2000:EKD

[25] Davis Baird. Encapsulating knowledge: The direct reading spectrometer. Foundations of Chemistry, 2(1):5–46, ???? 2000. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1023/A:1009958604987.

McEvoy:2000:SCR

[26] John G. McEvoy. In search of the chemical revolution: Interpretive strategies in the history of chemistry. Foundations of Chemistry, 2(1): 47–73, ???? 2000. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1009968823008.

Nesbet:2000:LE

[27] Robert K. Nesbet. Letter to the Editor. Foundations of Chemistry, 2(1): 75–76, ???? 2000. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL http://link.springer.com/accesspage/article/
10.1023/A:1009907629118.

Scerri:2000:R

[28] Eric Scerri. Response. Foundations of Chemistry, 2(1):77–78, ???? 2000. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1023/A:1009963900271.

Ramsey:2000:BRJ

[29] Jeffry L. Ramsey. Book review: Joachim Schummer, Realismus und Chemie: Philosophische Untersuchungen der Wissenschaft von den Stoffen. Foundations of Chemistry, 2(1):79–84, ???? 2000. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/article/10.1023/A:1009905108906.

Scerri:2000:Eb

[30] Eric R. Scerri. Editorial 5. Foundations of Chemistry, 2(2):95–98, May 2000. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL http://link.springer.com/article/10.1023/A:
1009930610570.

Akeroyd:2000:FMO

[31] F. Michael Akeroyd. The foundations of modern organic chemistry: The rise of the Highes and Ingold theory from 1930–1942. Foundations of Chemistry, 2(2):99–125, May 2000. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1023/A:1009942227721.

Needham:2000:ANA

[32] Paul Needham. Atomic notation and atomistic hypotheses translated by Paul Needham. Foundations of Chemistry, 2(2):127–180, May 2000. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1023/A:1009964831496.

Morrel:2000:BRD

[33] Jack Morrel. Book review: D. M. Knight and H. Kragh (eds.): The Making of the Chemist: The Social History of Chemistry in Europe, 1789–1914. Foundations of Chemistry, 2(2):181–185, May 2000. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1023/A:1009906827722.

Anonymous:2000:CP

[34] Anonymous. Call for papers. Foundations of Chemistry, 2(2):187–188, May 2000. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL http://link.springer.com/accesspage/article/
10.1023/A:1017261913227.

Scerri:2000:Ec

[35] Eric R. Scerri. Editorial 6. Foundations of Chemistry, 2(3):189–194, October 2000. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1009959528934.

Benfey:2000:RPC

[36] Theodor Benfey. Reflections on the philosophy of chemistry and a rallying call for our discipline. Foundations of Chemistry, 2(3):195– 205, October 2000. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1009946914577.

Kovac:2000:PEC

[37] Jeffrey Kovac. Professionalism and ethics in chemistry. Foundations of Chemistry, 2(3):207–219, October 2000. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1023/A:1009919031335.

King:2000:RME

[38] R. Bruce King. The role of mathematics in the experimental/theoretical/ computational trichotomy of chemistry. Foundations of Chemistry, 2(3): 221–236, October 2000. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1009989423317.

Sneath:2000:NCC

[39] P. H. A. Sneath. Numerical classification of the chemical elements and its relation to the periodic system. Foundations of Chemistry, 2(3):237– 263, October 2000. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1009933705492.

Anonymous:2000:AIV

[40] Anonymous. Authors index volume 2. Foundations of Chemistry, 2(3): 265, October 2000. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL http://link.springer.com/accesspage/article/
10.1023/A:1017261930265.

Anonymous:2000:CV

[41] Anonymous. Contents of volume 2. Foundations of Chemistry, 2(3):267– 268, October 2000. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL http://link.springer.com/accesspage/article/
10.1023/A:1017246613426.

Scerri:2001:Ea

[42] Eric R. Scerri. Editorial 7. Foundations of Chemistry, 3(1):1–5, ???? 2001. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1023/A:1011400124413.

Klein:2001:BFP

[43] Ursula Klein. Berzelian formulas as paper tools in early Nineteenth-Century chemistry. Foundations of Chemistry, 3(1):7–32, ???? 2001. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1023/A:1011460318779.

Shanks:2001:MBS

[44] Niall Shanks. Modeling biological systems: The Belousov–Zhabotinsky reaction. Foundations of Chemistry, 3(1):33–53, ???? 2001. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1023/A:1011434929814.

Chayut:2001:PGE

[45] Michael Chayut. From the periphery: the genesis of Eugene P. Wigner’s application of group theory to quantum mechanics. Foundations of Chemistry, 3(1):55–78, ???? 2001. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1023/A:1011431408763.

Psillos:2001:SSR

[46] Stathis Psillos. Studies in scientific realism. Foundations of Chemistry, 3 (1):79–86, ???? 2001. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1011403012572.

Anonymous:2001:ISP

[47] Anonymous. The International Society for the Philosophy of Chemistry. Foundations of Chemistry, 3(1):87–88, ???? 2001. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A:1017377925321.

Anonymous:2001:IA

[48] Anonymous. Instructions for authors. Foundations of Chemistry, 3(1): 89–96, ???? 2001. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1017328226350.

Scerri:2001:ESI

[49] Eric R. Scerri. Editorial 8 — special issue on the periodic system of the elements. Foundations of Chemistry, 3(2):97–104, May 2001. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1023/A:1011443411806.

Giunta:2001:APS

[50] Carmen J. Giunta. Argon and the periodic system: the piece that would not fit. Foundations of Chemistry, 3(2):105–128, May 2001. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1023/A:1011464516139.

Kragh:2001:FSE

[51] Helge Kragh. The first subatomic explanations of the periodic system. Foundations of Chemistry, 3(2):129–143, May 2001. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/article/10.1023/A:1011448410646.

Ostrovsky:2001:WHP

[52] V. N. Ostrovsky. What and how physics contributes to understanding the periodic law. Foundations of Chemistry, 3(2):145–181, May 2001. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1023/A:1011476405933.

Scerri:2001:BSS

[53] Eric R. Scerri and Jacob Edwards. Bibilography of secondary sources on the periodic system of the chemical elements. Foundations of Chemistry, 3(2):183–195, May 2001. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1023/A:1011458803337.

Scerri:2001:Eb

[54] Eric R. Scerri. Editorial 9. Foundations of Chemistry, 3(3):197–199, October 2001. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1012989418752.

Seibert:2001:CPH

[55] Charles Seibert. Charley Peirce’s head start in chemistry. Foundations of Chemistry, 3(3):201–226, October 2001. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1023/A:1012901029626.

LeVent:2001:WPG

[56] S. Le Vent. What is a perfect gas mixture? Foundations of Chemistry, 3 (3):227–239, October 2001. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1023/A:1012972025303.

Psarros:2001:LBH

[57] Nikos Psarros. The lame and the blind, or how much physics does chemistry need? Foundations of Chemistry, 3(3):241–249, October 2001. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1023/A:1012982928838.

vanBrakel:2001:WUM

[58] J. van Brakel. The world: An unruly mess. Foundations of Chemistry, 3 (3):251–262, October 2001. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1023/A:1012926808391.

Ramsey:2001:HPN

[59] Jeffry L. Ramsey. Hegel’s philosophy of nature. Foundations of Chemistry, 3(3):263–268, October 2001. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1023/A:1012932023412.

Laszlo:2001:SP

[60] Pierre Laszlo. A sketch of a program. Foundations of Chemistry, 3(3): 269–271, October 2001. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1012958923716.

Anonymous:2001:AIV

[61] Anonymous. Authors index volume 3. Foundations of Chemistry, 3(3): 273, October 2001. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL http://link.springer.com/accesspage/article/
10.1023/A:1017408801856.

Anonymous:2001:CV

[62] Anonymous. Contents of volume 3. Foundations of Chemistry, 3(3):275– 276, October 2001. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL http://link.springer.com/accesspage/article/
10.1023/A:1017317629985.

Scerri:2002:Ea

[63] Eric R. Scerri. Editorial 10. Foundations of Chemistry, 4(1):1–4, ???? 2002. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL http://link.springer.com/article/10.1023/A:
1014457521660.

Hentschel:2002:WOM

[64] Klaus Hentschel. Why not one more imponderable? John William Draper’s tithonic rays. Foundations of Chemistry, 4(1):5–59, ???? 2002. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1023/A:1014443008465.

Kerber:2002:MRH

[65] Robert C. Kerber. Markovnikov’s rule in history and pedagogy. Foundations of Chemistry, 4(1):61–72, ???? 2002. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1023/A:1014479921278.

Nye:2002:BRN

[66] Mary Jo Nye. Book review: Nalini Bhushan and Stuart Rosenfeld, eds: Of Minds and Molecules: New Philosophical Perspectives on Chemistry. Foundations of Chemistry, 4(1):73–77, ???? 2002. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/article/10.1023/A:1014496426735.

Pence:2002:BRL

[67] Harry E. Pence. Book review: A. Lundgren and B. Bensaude-Vincent, eds: Communicating Chemistry: Textbooks and their Audiences, 1789– 1929. Foundations of Chemistry, 4(1):79–81, ???? 2002. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1023/A:1014479318540.

Anonymous:2002:CPa

[68] Anonymous. Call for papers (1). Foundations of Chemistry, 4(1):83, ???? 2002. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL http://link.springer.com/accesspage/article/
10.1023/A:1017450326453.

Anonymous:2002:CPb

[69] Anonymous. Call for papers (2). Foundations of Chemistry, 4(1):84, ???? 2002. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL http://link.springer.com/accesspage/article/
10.1023/A:1017457710523.

Anonymous:2002:IAa

[70] Anonymous. Instructions for authors. Foundations of Chemistry, 4(1): 85–92, ???? 2002. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1017353113146.

Scerri:2002:Eb

[71] Eric R. Scerri. Editorial 11. Foundations of Chemistry, 4(2):93–96, May 2002. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL http://link.springer.com/article/10.1023/A:
1016055204995.

Jacob:2002:PBR

[72] Claus Jacob. Philosophy and biochemistry: Research at the interface between chemistry and biology. Foundations of Chemistry, 4(2): 97–125, May 2002. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1016052605025.

Brooks:2002:DPL

[73] Nathan M. Brooks. Developing the periodic law: Mendeleev’s work during 1869–1871. Foundations of Chemistry, 4(2):127–147, May 2002. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1023/A:1016022129746.

Cintas:2002:OTC

[74] Pedro Cintas. On the origin of tetrahedral carbon: A case for philosophy of chemistry? Foundations of Chemistry, 4(2):149–161, May 2002. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1023/A:1016063004499.

Kovac:2002:TPR

[75] Jeffrey Kovac. Theoretical and practical reasoning in chemistry. Foundations of Chemistry, 4(2):163–171, May 2002. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1023/A:1016035726186.

Luna:2002:BRP

[76] Fernando J. Luna. Book review: Pierre Laszlo: Miroir de la Chimie. Foundations of Chemistry, 4(2):173–177, May 2002. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1023/A:1016041703950.

Scerri:2002:Ec

[77] Eric R. Scerri. Editorial 12. Foundations of Chemistry, 4(3):179– 182, October 2002. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1020610630274.

Luisi:2002:ECC

[78] Pier Luigi Luisi. Emergence in chemistry: Chemistry as the embodiment of emergence. Foundations of Chemistry, 4(3):183–200, October 2002. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1023/A:1020672005348.

Eberhart:2002:QMM

[79] Mark Eberhart. Quantum mechanics and molecular design in the twenty first century. Foundations of Chemistry, 4(3):201–211, October 2002. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1023/A:1020655631654.

Justi:2002:PCU

[80] Rosária S. Justi and John K. Gilbert. Philosophy of chemistry in university chemical education: The case of models and modelling. Foundations of Chemistry, 4(3):213–240, October 2002. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1023/A:1020608215725.

Bloor:2002:BRR

[81] John E. Bloor. Book review: Ronald J. Gillespie and Paul L. A. Popelier: Chemical Bonding and Molecular Geometry: From Lewis to Electron Densities. Foundations of Chemistry, 4(3):241–247, October 2002. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1023/A:1020691330843.

Anonymous:2002:AIV

[82] Anonymous. Author index volume 4. Foundations of Chemistry, 4(3):249, October 2002. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL http://link.springer.com/accesspage/article/
10.1023/A:1020643221655.

Anonymous:2002:CV

[83] Anonymous. Contents of volume 4. Foundations of Chemistry, 4(3):251– 252, October 2002. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL http://link.springer.com/accesspage/article/
10.1023/A:1020699623472.

Anonymous:2002:IAb

[84] Anonymous. Instructions for authors. Foundations of Chemistry, 4(3): 253–260, October 2002. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1020647605725.

Scerri:2003:Ea

[85] Eric R. Scerri. Editorial 13. Foundations of Chemistry, 5(1):1–6, ???? 2003. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL http://link.springer.com/article/10.1023/A:
1021949109810.

Vihalemm:2003:LNS

[86] R. Vihalemm. Are laws of nature and scientific theories peculiar in chemistry? Scrutinizing Mendeleev’s discovery. Foundations of Chemistry, 5 (1):7–22, ???? 2003. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/

A:1021980526951. See response [96].

Reiher:2003:STC

[87] Markus Reiher. A systems theory for chemistry. Foundations of Chemistry, 5(1):23–41, ???? 2003. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1023/A:1021995611796.

Taber:2003:ACC

[88] Keith S. Taber. The atom in the chemistry curriculum: Fundamental concept, teaching model or epistemological obstacle? Foundations of Chemistry, 5(1):43–84, ???? 2003. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1023/A:1021995612705.

Akeroyd:2003:PRP

[89] Michael Akeroyd. Predictions, retrodictions and the periodic table. Foundations of Chemistry, 5(1):85–88, ???? 2003. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1023/A:1021965027473.

Kauffman:2003:BRC

[90] George B. Kauffman and Laurie M. Kauffman. Book review: Carl Djerassi: An Immaculate Misconception: Sex in an Age of Mechanical Reproduction. Foundations of Chemistry, 5(1):89–91, ???? 2003. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1023/A:1021926610496.

Ramsey:2003:UKE

[91] Jeffry L. Ramsey. U. Klein (ed.): Tools and modes of representation in the laboratory sciences. Foundations of Chemistry, 5(1):93–97, ???? 2003. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1023/A:1021934727335.

Anonymous:2003:IAa

[92] Anonymous. Instructions for authors. Foundations of Chemistry, 5(1): 99–106, ???? 2003. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1021912932560.

Scerri:2003:Eb

[93] Eric R. Scerri. Editorial 14. Foundations of Chemistry, 5(2):107– 111, May 2003. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1023649307266.

Paneth:2003:ESC

[94] F. A. Paneth. The epistemological status of the chemical concept of element. Foundations of Chemistry, 5(2):113–145, May 2003. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1023/A:1023600703644.

Reiher:2003:STV

[95] Markus Reiher. The systems-theoretical view of chemical concepts. Foundations of Chemistry, 5(2):147–163, May 2003. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1023/A:1023672816861.

Christie:2003:CLT

[96] John R. Christie and Maureen Christie. Chemical laws and theories: A response to Vihalemm. Foundations of Chemistry, 5(2):165–174, May 2003. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1023/A:1023631726532. See

[86].

Ramsey:2003:BRP

[97] Jeffry L. Ramsey. Book review: P. J. T. Morris and O. T. Benfey (eds.): Robert Burns Woodward: Architect and Artist in the World of Molecules (History of Modern Chemical Sciences Series). Foundations of Chemistry, 5(2):175–178, May 2003. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1023/A:1023693730053.

Hentschel:2003:IRC

[98] habil. Klaus Hentschel. The instrumental revolution in chemistry. Foundations of Chemistry, 5(2):179–183, May 2003. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1023/A:1023691917565.

Scerri:2003:Ec

[99] Eric R. Scerri. Editorial 15. Foundations of Chemistry, 5(3):185– 188, October 2003. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1025693811263.

Kaji:2003:MDP

[100] Masanori Kaji. Mendeleev’s discovery of the periodic law: The origin and the reception. Foundations of Chemistry, 5(3):189–214, October 2003. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1023/A:1025673206850.

Spector:2003:AMR

[101] Tami I. Spector. The aesthetics of molecular representation: From the empirical to the constitutive. Foundations of Chemistry, 5(3):215– 236, October 2003. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1025692518934.

Vancik:2003:PCL

[102] Hrvoj Vancik. Philosophy of chemistry and limits of complexity. Foundations of Chemistry, 5(3):237–247, October 2003. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1023/A:1025656727195.

Allin:2003:BRC

[103] Shawn B. Allin. Book review: Cathy Cobb: Magick, Mayhem, and Mavericks: The Spirited History of Physical Chemistry. Foundations of Chemistry, 5(3):249–252, October 2003. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1023/A:1025648701979.

Needham:2003:BRM

[104] Paul Needham. Book review: Maureen Christie: The Ozone Layer. A Philosophy of Science Perspective. Foundations of Chemistry, 5(3):253– 261, October 2003. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1025642128618.

Anonymous:2003:AIV

[105] Anonymous. Author index volume 5. Foundations of Chemistry, 5(3):263, October 2003. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL http://link.springer.com/accesspage/article/
10.1023/A:1025651709786.

Anonymous:2003:CV

[106] Anonymous. Contents of volume 5. Foundations of Chemistry, 5(3):265– 266, October 2003. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL http://link.springer.com/accesspage/article/
10.1023/A:1025668824808.

Anonymous:2003:IAb

[107] Anonymous. Instructions for authors. Foundations of Chemistry, 5(3): 267–273, October 2003. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
A:1025640107542.

Scerri:2004:Ea

[108] Eric R. Scerri. Editorial 16. Foundations of Chemistry, 6(1):1–2, ???? 2004. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1023/B:FOCH.0000021194.71352.22.

Bhushan:2004:I

[109] Nalini Bhushan. Introduction. Foundations of Chemistry, 6(1):3–9, ???? 2004. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1023/B:FOCH.
0000021170.34287.f7.

Hoffmann:2004:TC

[110] Roald Hoffmann. Theoretical chemistry. Foundations of Chemistry, 6(1): 11, ???? 2004. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL http://link.springer.com/accesspage/article/
10.1023/B:FOCH.0000020992.64107.f5.

Woody:2004:TSW

[111] Andrea I. Woody. Telltale signs: What common explanatory strategies in chemistry reveal about explanation itself. Foundations of Chemistry, 6 (1):13–43, ???? 2004. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
B:FOCH.0000020993.09368.b6. See response [112].

Weininger:2004:RTS

[112] Stephen J. Weininger. Response to “Telltale Signs: What Common Explanatory Strategies in Chemistry Reveal About Explanation Itself”. Foundations of Chemistry, 6(1):45–48, ???? 2004. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/article/10.1023/B:FOCH.0000020994.23928.4d. See

[111].

Douglas:2004:PED

[113] Heather Douglas. Prediction, explanation, and dioxin biochemistry: Science in public policy. Foundations of Chemistry, 6(1):49–63, ???? 2004. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1023/B:FOCH.
0000020995.75920.81.

Wiberg:2004:SSS

[114] Kenneth B. Wiberg. Strain, structure, stability and reactivity. Foundations of Chemistry, 6(1):65–80, ???? 2004. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1023/B:FOCH.0000020996.34436.71.

Ramsey:2004:SES

[115] Jeffry L. Ramsey. Straining to explain strain and synthesis. Foundations of Chemistry, 6(1):81–91, ???? 2004. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1023/B:FOCH.0000020997.72966.3b.

Scerri:2004:JHI

[116] Eric R. Scerri. Just how ab initio is ab initio quantum chemistry? Foundations of Chemistry, 6(1):93–116, ???? 2004. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1023/B:FOCH.0000020998.31689.16. See comments

[117, 149].

Friedrich:2004:LHI

[117] Bretislav Friedrich. … hasn’t it? A commentary on Eric Scerri’s paper “Has Quantum Mechanics Explained the Periodic Table?”. Foundations of Chemistry, 6(1):117–132, ???? 2004. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1023/B:FOCH.0000020999.69326.89. See [116].

Anonymous:2004:CP

[118] Anonymous. Call for papers. Foundations of Chemistry, 6(1):133, ???? 2004. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL http://link.springer.com/accesspage/article/
10.1023/B:FOCH.0000021196.65720.cc.

Scerri:2004:Eb

[119] Eric R. Scerri. Editorial 17. Foundations of Chemistry, 6(2):135–136, May 2004. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL http://link.springer.com/accesspage/article/
10.1023/B:FOCH.0000035095.09138.6b.

Eddy:2004:EPN

[120] M. D. Eddy. Elements, principles and the narrative of affinity. Foundations of Chemistry, 6(2):161–175, May 2004. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1023/B:FOCH.0000035061.02831.45.

Laszlo:2004:BRB

[121] Pierre Laszlo. Book review: Mapping the Spectrum. Techniques of Visual Representation in Research and Teaching. Foundations of Chemistry, 6 (2):177–189, May 2004. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
B:FOCH.0000034991.05140.15.

Sanchez:2004:BRB

[122] José Ramón Bertomeu Sánchez. Book review: Lavoisier in Italia. Foundations of Chemistry, 6(2):191–195, May 2004. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1023/B:FOCH.0000034990.68395.20.

Anonymous:2004:Ca

[123] Anonymous. Corrigendum. Foundations of Chemistry, 6(2):197, May 2004. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1023/B:FOCH.0000035096.25553.8e.

Earley:2004:WIC

[124] Joseph E. Earley. Would introductory chemistry courses work better with a new philosophical basis? Foundations of Chemistry, 6(3):137–160, October 2004. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1023/B:FOCH.
0000034992.42777.95.

Scerri:2004:Ec

[125] Eric R. Scerri. Editorial 18. Foundations of Chemistry, 6(3):199–201, October 2004. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1023/B:FOCH.
0000041643.10770.5d.

Deltete:2004:BRB

[126] Robert J. Deltete and Anastasios Brenner. Book review: Pierre Duhem: Mixture and Chemical Combination and Related Essays. Edited and translated, with an Introduction, by Paul Needham. Foundations of Chemistry, 6(3):203–232, October 2004. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1023/B:FOCH.0000041548.61723.77. See response [127].

Needham:2004:MCC

[127] Paul Needham. Mixture and chemical combination and related essays: A response to Robert Deltete and Anastasios Brenner. Foundations of Chemistry, 6(3):233–245, October 2004. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1023/B:FOCH.0000041549.06984.c3. See [126].

Kauffman:2004:BRF

[128] George B. Kauffman and Laurie M. Kauffman. Book review: Fred Basolo: From Coello to Inorganic Chemistry: A Lifetime of Reactions. Foundations of Chemistry, 6(3):247–250, October 2004. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1023/B:FOCH.0000041953.83552.86.

Anonymous:2004:Cb

[129] Anonymous. Corrigendum. Foundations of Chemistry, 6(3):251, October 2004. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL http://link.springer.com/accesspage/article/
10.1023/B:FOCH.0000041679.58052.9f.

Anonymous:2004:AIV

[130] Anonymous. Author index volume 6. Foundations of Chemistry, 6(3):253, October 2004. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL http://link.springer.com/accesspage/article/
10.1023/B:FOCH.0000041958.76547.c4.

Anonymous:2004:CV

[131] Anonymous. Contents of volume 6. Foundations of Chemistry, 6(3):255– 257, October 2004. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
B:FOCH.0000041957.11094.3d.

Anonymous:2004:IA

[132] Anonymous. Instructions for authors. Foundations of Chemistry, 6(3): 259–265, October 2004. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
B:FOCH.0000041956.39391.0f.

Scerri:2005:ESI

[133] Eric R. Scerri. Editorial 19 special issue on philosophical problems of chemical kinds. Foundations of Chemistry, 7(1):1–4, ???? 2005. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1023/B:FOCH.0000042937.36350.5a.

Harre:2005:I

[134] Rom Harré. Introduction. Foundations of Chemistry, 7(1):5, ???? 2005. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1023/B:FOCH.0000042884.40307.a5.

Harre:2005:CKE

[135] Rom Harré. Chemical kinds and essences revisited. Foundations of Chemistry, 7(1):7–30, ???? 2005. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1023/B:FOCH.0000042885.93975.69.

Hendry:2005:LME

[136] Robin Findlay Hendry. Lavoisier and Mendeleev on the elements. Foundations of Chemistry, 7(1):31–48, ???? 2005. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1023/B:FOCH.0000042886.65679.4e.

Simonian:2005:PCC

[137] Joseph Simonian. The paradoxes of chemical classification: Why ‘water is H_{2}O’ is not an identity statement. Foundations of Chemistry, 7(1): 49–56, ???? 2005. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
B:FOCH.0000042887.03317.11.

vanBrakel:2005:IX

[138] J. van Brakel. On the inventors of XYZ. Foundations of Chemistry, 7 (1):57–84, ???? 2005. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1023/
B:FOCH.0000042888.33192.2e.

Earley:2005:WTN

[139] Joseph E. Earley. Why there is no salt in the sea. Foundations of Chemistry, 7(1):85–102, ???? 2005. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1023/B:FOCH.0000042881.05418.15.

Needham:2005:MM

[140] Paul Needham. Mixtures and modality. Foundations of Chemistry, 7 (1):103–118, ???? 2005. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1023/B:FOCH.0000042889.36764.49.

Scerri:2005:Ea

[141] Eric R. Scerri. Editorial 20. Foundations of Chemistry, 7(2):119–123, ???? 2005. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
005-0118-5.

Lombardi:2005:OAC

[142] Olimpia Lombardi and Martín Labarca. The ontological autonomy of the chemical world. Foundations of Chemistry, 7(2):125–148, ???? 2005. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-

004-0980-6. See comment [164].

Wilholt:2005:EMT

[143] Torsten Wilholt. Explaining models: Theoretical and phenomenological models and their role for the first explanation of the hydrogen spectrum. Foundations of Chemistry, 7(2):149–169, ???? 2005. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-004-5958-x.

Vihalemm:2005:CTM

[144] Rein Vihalemm. Chemistry and a theoretical model of science: On the occasion of a recent debate with the Christies. Foundations of Chemistry, 7(2):171–182, ???? 2005. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1007/s10698-005-0959-y.

Kauffman:2005:BRF

[145] George B. Kauffman. Book review: Fathi Habashi: From Alchemy to Atomic Bombs: History of Chemistry, Metallurgy, and Civilization. Métallurgie Extractive Québec: 800 rue Alain #504, Sainte Foy, Québec, Canada G1X 4E7, 2002; distributed by Laval University Bookstore “Zone”: Cité Universitaire, Sainte Foy, Québec, Canada G1K 7P4, viii + 357 pp, Can.$70.00; U.S.$50.00; plus postage (hardbound); ISBN 2-922686-00-0. Foundations of Chemistry, 7(2):183–186, ???? 2005. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1007/s10698-004-5960-3.

Hendry:2005:BRJ

[146] Robin Findlay Hendry. Book review: Jaap van Brakel: Philosophy of Chemistry: Between the Manifest and the Scientific Image Leuven, University Press, Leuven, 2000, xiv + 246 pp., ISBN 90-5867-063-5. Foundations of Chemistry, 7(2):187–197, ???? 2005. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1007/s10698-004-0491-5.

Scerri:2005:Eb

[147] Eric R. Scerri. Editorial 21. Foundations of Chemistry, 7(3):199–202, October 2005. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
005-4510-y.

Laing:2005:RPT

[148] Michael Laing. A revised periodic table: With the lanthanides repositioned. Foundations of Chemistry, 7(3):203–233, October 2005. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-004-5959-9.

Ostrovsky:2005:RDC

[149] V. N. Ostrovsky. On recent discussion concerning quantum justification of the periodic table of the elements. Foundations of Chemistry, 7(3):235– 239, October 2005. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-005-2141-y. See [116, 117].

Novaro:2005:ACD

[150] Octavio Novaro. Activity of closed d-shells in noble metal atoms. Foundations of Chemistry, 7(3):241–268, October 2005. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1007/s10698-004-0787-5.

Kidwai:2005:GCI

[151] M. Kidwai and R. Mohan. Green chemistry: An innovative technology. Foundations of Chemistry, 7(3):269–287, October 2005. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-004-2783-1.

Balaban:2005:RAM

[152] A. T. Balaban. Reflections about mathematical chemistry. Foundations of Chemistry, 7(3):289–306, October 2005. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1007/s10698-005-0779-0.

Korobov:2005:SCR

[153] A. Korobov. Simple chemical reactions in the solid state: Towards elaborating a conception. Foundations of Chemistry, 7(3):307–314, October 2005. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
005-7326-x.

Giunta:2005:BRM

[154] Carmen Giunta. Book review: Michael D. Gordin: A Well-Ordered Thing: Dmitrii Mendeleev and the Shadow of the Periodic Table, Basic Books, New York, 2004, 364 + xx pp., ISBN 0-465-02775-X, $30. Foundations of Chemistry, 7(3):315–319, October 2005. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1007/s10698-005-7201-9.

Pence:2005:BRM

[155] Harry E. Pence. Book review: Mark Albert: Galen’s Lectures: A Novel About Chemistry, Xlibris Corporation, 2000, 493 pp. (ISBN 0-7388-4196-X). Foundations of Chemistry, 7(3):321–324, October 2005. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-005-7197-1.

Anonymous:2005:AIV

[156] Anonymous. Author index volume 7. Foundations of Chemistry, 7(3):325, October 2005. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s10698-005-4509-4;   http://link.
springer.com/article/10.1007/s10698-005-4509-4.

Anonymous:2005:CV

[157] Anonymous. Contents of volume 7. Foundations of Chemistry, 7(3):327– 329, October 2005. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-005-4515-6.

Anonymous:2005:IA

[158] Anonymous. Instructions for authors. Foundations of Chemistry, 7(3): 331–337, October 2005. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-005-4516-5.

Scerri:2006:Ea

[159] Eric Scerri. Editorial 22. Foundations of Chemistry, 8(1):1–2, February 2006. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/s10698-006-9011-0;   http://link.springer.com/article/10.
1007/s10698-006-9011-0.

Stein:2006:PTE

[160] Ross L. Stein. A process theory of enzyme catalytic power- the interplay of science and metaphysics. Foundations of Chemistry, 8(1):3–29, February 2006. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
005-7907-8.

Sacks:2006:CPH

[161] Lawrence J. Sacks. Concerning the position of hydrogen in the periodic table. Foundations of Chemistry, 8(1):31–35, February 2006. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-005-9003-5.

Shahbazian:2006:RON

[162] Shant Shahbazian and Mansour Zahedi. The role of observables and nonobservables in chemistry: A critique of chemical language. Foundations of Chemistry, 8(1):37–52, February 2006. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1007/s10698-005-8247-4.

Schummer:2006:GSM

[163] Joachim Schummer. Gestalt switch in molecular image perception: The aesthetic origin of molecular nanotechnology in supramolecular chemistry. Foundations of Chemistry, 8(1):53–72, February 2006. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-005-3358-5.

Needham:2006:ORC

[164] Paul Needham. Ontological reduction: a comment on Lombardi and Labarca. Foundations of Chemistry, 8(1):73–80, February 2006. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-005-9002-6. See [142].

Lombardi:2006:OAC

[165] Olimpia Lombardi and Martín Labarca. The ontological autonomy of the chemical world: a response to Needham. Foundations of Chemistry, 8 (1):81–92, February 2006. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1007/s10698-005-9004-4.

Scerri:2006:Eb

[166] Eric Scerri. Editorial 23. Foundations of Chemistry, 8(2):93–95, July 2006. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1007/s10698-006-9021-y.

Labinger:2006:OSN

[167] Jay A. Labinger. Organized skepticism, naïve methodism, and other-isms. Foundations of Chemistry, 8(2):97–110, July 2006. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/article/10.1007/s10698-004-3531-2.

Wink:2006:CBP

[168] Donald J. Wink. Connections between pedagogical and epistemological constructivism: Questions for teaching and research in chemistry. Foundations of Chemistry, 8(2):111–151, July 2006. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1007/s10698-006-9008-8.

deBerg:2006:SCC

[169] Kevin C. de Berg. The status of constructivism in chemical education research and its relationship to the teaching and learning of the concept of idealization in chemistry. Foundations of Chemistry, 8(2):153–176, July 2006. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
006-9010-1.

Cardellini:2006:FRC

[170] Liberato Cardellini. The foundations of radical constructivism: An interview with Ernst von Glasersfeld. Foundations of Chemistry, 8(2):177–187, July 2006. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
006-9012-z.

Taber:2006:CNC

[171] Keith S. Taber. Constructivism’s new clothes: The trivial, the contingent, and a progressive research programme into the learning of science. Foundations of Chemistry, 8(2):189–219, July 2006. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1007/s10698-005-4536-1.

Scerri:2006:Ec

[172] Eric Scerri. Editorial 24. Foundations of Chemistry, 8(3):221–223, October 2006. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
006-9028-4.

Sharlow:2006:CEP

[173] M. F. Sharlow. Chemical elements and the problem of universals. Foundations of Chemistry, 8(3):225–242, October 2006. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1007/s10698-006-9016-8.

Forster:2006:CPP

[174] Leslie S. Forster. Chromium photophysics- a prototypical case history. Foundations of Chemistry, 8(3):243–254, October 2006. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-006-9017-7.

Snooks:2006:ASP

[175] R. J. Snooks. Another scientific practice separating chemistry from physics: thought experiments. Foundations of Chemistry, 8(3):255– 270, October 2006. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-006-9019-5.

Earley:2006:SPI

[176] Joseph E. Earley, Sr. Some philosophical influences on Ilya Prigogine’s statistical mechanics. Foundations of Chemistry, 8(3):271–283, October 2006. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1007/s10698-006-9007-9.

Scerri:2006:CAK

[177] Eric R. Scerri. Commentary on Allen & Knight’s response to the Löwdin Challenge. Foundations of Chemistry, 8(3):285–292, October 2006. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1007/s10698-006-9009-7.

Bloor:2006:BR

[178] John E. Bloor and W. H. Eugen Schwarz. Book review. Foundations of Chemistry, 8(3):293–303, October 2006. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1007/s10698-006-9014-x. See response [179].

King:2006:RDH

[179] R. B. King and D. H. Rouvray. Response of D. H. Rouvray and R. B. King, Editors of the Book The Periodic Table: Into the 21st Century. Foundations of Chemistry, 8(3):305–306, October 2006. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10698-006-9013-y;
http://link.springer.com/article/10.1007/s10698-006-9013-y. See

[178].

Pagni:2006:BR

[180] R. Pagni. Book review. Foundations of Chemistry, 8(3):307–310, October 2006. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
006-9015-9.

Scerri:2007:Ea

[181] Eric R. Scerri. Editorial 25. Foundations of Chemistry, 9(1):1, April 2007. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/s10698-007-9031-4;   http://link.springer.com/article/10.
1007/s10698-007-9031-4.

Deltete:2007:WOEa

[182] R. J. Deltete. Wilhelm Ostwald’s energetics 1: Origins and motivations. Foundations of Chemistry, 9(1):3–56, April 2007. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/article/10.1007/s10698-005-6707-5.

Jacob:2007:CDC

[183] Claus Jacob. The closure of the Department of Chemistry at the University of Exeter — an insider’s view. Foundations of Chemistry, 9(1): 57–64, April 2007. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-006-9006-x.

Napinen:2007:NHU

[184] Leo Näpinen. The need for the historical understanding of nature in physics and chemistry. Foundations of Chemistry, 9(1):65–84, April 2007. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1007/s10698-006-9018-6.

Shahbazian:2007:LEC

[185] Shant Shahbazian and Mansour Zahedi. Letter to the Editor: The concept of chemical bond- some like it fuzzy but others concrete. Foundations of Chemistry, 9(1):85–95, April 2007. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1007/s10698-006-9023-9.

Giunta:2007:BR

[186] Carmen J. Giunta. Book review. Foundations of Chemistry, 9(1):97–100, April 2007. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
006-9005-y.

Stemwedel:2007:BR

[187] Janet D. Stemwedel. Book review. Foundations of Chemistry, 9(1): 101–114, April 2007. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-006-9024-8.

Scerri:2007:Eb

[188] Eric R. Scerri. Editorial 26. Foundations of Chemistry, 9(2):115–117, July 2007. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
007-9037-y.

Rayner-Canham:2007:CKM

[189] Geoff Rayner-Canham and Megan Oldford. The chemical ‘Knight’s Move’ relationship: what is its significance? Foundations of Chemistry, 9(2): 119–125, July 2007. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-007-9034-1.

Laing:2007:WPH

[190] Michael Laing. Where to put hydrogen in a periodic table? Foundations of Chemistry, 9(2):127–137, July 2007. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1007/s10698-006-9027-5.

Schwarz:2007:RQR

[191] W. H. Eugen Schwarz. Recommended questions on the road towards a scientific explanation of the periodic system of chemical elements with the help of the concepts of quantum physics. Foundations of Chemistry, 9(2):139–188, July 2007. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1007/s10698-006-9020-z.

Restrepo:2007:MAP

[192] Guillermo Restrepo and Leonardo Pachón. Mathematical aspects of the periodic law. Foundations of Chemistry, 9(2):189–214, July 2007. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-006-9026-6.

Rayner-Canham:2007:BRE

[193] G. W. Rayner-Canham. Book review: Eric R. Scerri, The Periodic Table: its Story and its Significance. Foundations of Chemistry, 9(2):215–218, July 2007. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
007-9032-3.

Scerri:2007:Ec

[194] Eric Scerri. Editorial 27. Foundations of Chemistry, 9(3):219–220, October 2007. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/s10698-007-9040-3;   http://link.springer.com/article/10.
1007/s10698-007-9040-3.

Kibler:2007:MPT

[195] Maurice R. Kibler. From the Mendeleev periodic table to particle physics and back to the periodic table. Foundations of Chemistry, 9(3):221– 234, October 2007. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-007-9039-9.

Stewart:2007:CDM

[196] Philip J. Stewart. A century on from Dmitrii Mendeleev: tables and spirals, noble gases and Nobel prizes. Foundations of Chemistry, 9(3): 235–245, October 2007. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-007-9038-x.

Erduran:2007:BLP

[197] Sibel Erduran. Breaking the law: promoting domain-specificity in chemical education in the context of arguing about the periodic law. Foundations of Chemistry, 9(3):247–263, October 2007. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1007/s10698-007-9036-z.

Deltete:2007:WOEb

[198] R. J. Deltete. Wilhelm Ostwald’s energetics 2: energetic theory and applications, part I. Foundations of Chemistry, 9(3):265–316, October 2007. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1007/s10698-006-9025-7.

Ruthenberg:2007:SCP

[199] Klaus Ruthenberg. Saving chemical phenomena outside the scientific community. Foundations of Chemistry, 9(3):317–320, October 2007. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-007-9030-5.

Scerri:2008:Ea

[200] Eric R. Scerri. Editorial 28. Foundations of Chemistry, 10(1):1–2, April 2008. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/s10698-008-9043-8;   http://link.springer.com/article/10.
1007/s10698-008-9043-8.

Novaro:2008:RPW

[201] Octavio Novaro. On the rightful place for He within the periodic table. Foundations of Chemistry, 10(1):3–12, April 2008. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/article/10.1007/s10698-007-9041-2.

Rayner-Canham:2008:NEA

[202] Geoff Rayner-Canham and Zheng Zheng. Naming elements after scientists: an account of a controversy. Foundations of Chemistry, 10(1):13–18, April 2008. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
007-9042-1.

Harre:2008:SPM

[203] Rom Harré. Some presuppositions in the metaphysics of chemical reactions. Foundations of Chemistry, 10(1):19–38, April 2008. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-005-9000-8.

Akeroyd:2008:MEV

[204] F. Michael Akeroyd. Mechanistic explanation versus deductivenomological explanation. Foundations of Chemistry, 10(1):39–48, April 2008. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
005-9001-7.

Newman:2008:CS

[205] Micah Newman. Chemical supervenience. Foundations of Chemistry, 10 (1):49–62, April 2008. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-006-9029-3.

Paparazzo:2008:WTC

[206] Ernesto Paparazzo. Why take chemistry stoically? The case of Posidonius. Foundations of Chemistry, 10(1):63–75, April 2008. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-007-9035-0.

Scerri:2008:Eb

[207] Eric Scerri. Editorial 29. Foundations of Chemistry, 10(2):77–78, July 2008. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/s10698-008-9048-3;   http://link.springer.com/article/10.
1007/s10698-008-9048-3.

Alvarez:2008:BCE

[208] Santiago Alvarez, Joaquim Sales, and Miquel Seco. On books and chemical elements. Foundations of Chemistry, 10(2):79–100, July 2008. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-008-9047-4.

Hoyningen-Huene:2008:TKC

[209] Paul Hoyningen-Huene. Thomas Kuhn and the chemical revolution. Foundations of Chemistry, 10(2):101–115, July 2008. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1007/s10698-008-9044-7.

Goodwin:2008:SFE

[210] W. M. Goodwin. Structural formulas and explanation in organic chemistry. Foundations of Chemistry, 10(2):117–127, July 2008. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-007-9033-2.

Cardellini:2008:VIE

[211] Liberato Cardellini. The views and influence of Ernst von Glasersfeld: An introduction. Foundations of Chemistry, 10(2):129–134, July 2008. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1007/s10698-006-9022-x.

Hettema:2008:NMW

[212] Hinne Hettema. A note on Michael Weisberg’s: challenges to the structural conception of chemical bonding. Foundations of Chemistry, 10(2): 135–142, July 2008. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-008-9045-6.

Scerri:2008:Ec

[213] Eric Scerri. Editorial 30. Foundations of Chemistry, 10(3):143, October 2008. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.

1007/s10698-008-9056-3;   http://link.springer.com/article/10.
1007/s10698-008-9056-3.

Spector:2008:IE

[214] Tami I. Spector. ISPC 2007 editorial. Foundations of Chemistry, 10(3):145–146, October 2008. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s10698-008-9057-2;   http://link.
springer.com/article/10.1007/s10698-008-9057-2.

Berson:2008:FTT

[215] Jerome A. Berson. Fundamental theories and their empirical patches. Foundations of Chemistry, 10(3):147–156, October 2008. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-008-9046-5.

Chalmers:2008:AAN

[216] Alan F. Chalmers. Atom and aether in nineteenth-century physical science. Foundations of Chemistry, 10(3):157–166, October 2008. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-008-9051-8.

Vemulapalli:2008:TCB

[217] G. K. Vemulapalli. Theories of the chemical bond and its true nature. Foundations of Chemistry, 10(3):167–176, October 2008. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-008-9049-2.

Goodwin:2008:IIT

[218] W. M. Goodwin. Implementation and innovation in total synthesis. Foundations of Chemistry, 10(3):177–186, October 2008. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/article/10.1007/s10698-008-9050-9.

Deltete:2008:WOE

[219] Robert J. Deltete. Wilhelm Ostwald’s energetics 3: energetic theory and applications, part II. Foundations of Chemistry, 10(3):187–221, October 2008. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
008-9053-6.

Scerri:2009:Ea

[220] Eric Scerri. Editorial 31. Foundations of Chemistry, 11(1):1–2, April 2009. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
009-9068-7;   http://link.springer.com/content/pdf/10.1007/s10698-
009-9068-7.pdf.

Spector:2009:ISE

[221] Tami I. Spector. ISPC 2007 second editorial. Foundations of Chemistry, 11(1):3–5, April 2009. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-009-9066-9;   http://link.springer.com/content/pdf/10.
1007/s10698-009-9066-9.pdf.

Sukumar:2009:CCM

[222] N. Sukumar. The chemist’s concept of molecular structure. Foundations of Chemistry, 11(1):7–20, April 2009. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1007/s10698-008-9060-7.

Salmon:2009:EE

[223] James Francis Salmon. Emergence in evolution. Foundations of Chemistry, 11(1):21–32, April 2009. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1007/s10698-008-9059-0.

Garcia:2009:DSN

[224] Pio García. Discovery by serendipity: a new context for an old riddle. Foundations of Chemistry, 11(1):33–42, April 2009. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/article/10.1007/s10698-008-9061-6.

Pagni:2009:ODA

[225] Richard Martin Pagni. The origin and development of the acidity function. Foundations of Chemistry, 11(1):43–50, April 2009. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-009-9065-x.

Tromble:2009:ACS

[226] Meredith Tromble. The advent of chemical symbolism in the art of Sonya Rapoport. Foundations of Chemistry, 11(1):51–60, April 2009. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-008-9058-1.

Scerri:2009:Eb

[227] Eric R. Scerri. Editorial 32. Foundations of Chemistry, 11(2):61–62, July 2009. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
009-9072-y;   http://link.springer.com/content/pdf/10.1007/s10698-
009-9072-y.pdf.

Spector:2009:ITE

[228] Tami I. Spector. ISPC 2007 third editorial. Foundations of Chemistry, 11 (2):63–64, July 2009. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-009-9073-x;   http://link.springer.com/content/pdf/10.
1007/s10698-009-9073-x.pdf.

Earley:2009:HCS

[229] Joseph E. Earley, Sr. How chemistry shifts horizons: element, substance, and the essential. Foundations of Chemistry, 11(2):65–77, July 2009. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1007/s10698-008-9054-5.

Ruthenberg:2009:PKP

[230] Klaus Ruthenberg. Paneth, Kant, and the philosophy of chemistry. Foundations of Chemistry, 11(2):79–91, July 2009. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1007/s10698-009-9064-y.

Harre:2009:TTO

[231] Rom Harré. Trope theory and the ontology of chemistry. Foundations of Chemistry, 11(2):93–103, July 2009. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1007/s10698-008-9052-7.

Pagni:2009:WNF

[232] Richard M. Pagni. The weak nuclear force, the chirality of atoms, and the origin of optically active molecules. Foundations of Chemistry, 11(2): 105–122, July 2009. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-009-9070-0.

Rayner-Canham:2009:ISF

[233] Geoff Rayner-Canham. Isoelectronic series: a fundamental periodic property. Foundations of Chemistry, 11(2):123–129, July 2009. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-008-9055-4.

Scerri:2009:Ec

[234] Eric R. Scerri. Editorial 33. Foundations of Chemistry, 11(3):131–133, October 2009. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
009-9080-y;   http://link.springer.com/content/pdf/10.1007/s10698-
009-9080-y.pdf.

Sima:2009:ONI

[235] Jozef Sima. Oxidation number: issues of its determination and range. Foundations of Chemistry, 11(3):135–143, October 2009. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-009-9071-z.

Hettema:2009:ETF

[236] Hinne Hettema. Explanation and theory formation in quantum chemistry. Foundations of Chemistry, 11(3):145–174, October 2009. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-009-9075-8.

DeKosky:2009:BRW

[237] Robert K. DeKosky. Book review: William H. Brock: William Crookes (1832–1919) and the Commercialization of Science. Foundations of Chemistry, 11(3):175–180, October 2009. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1007/s10698-009-9074-9.

McIntyre:2009:BRE

[238] Lee McIntyre. Book review: Eric Scerri: Collected papers on philosophy of chemistry. Foundations of Chemistry, 11(3):181–182, October 2009. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/s10698-009-9076-7;   http://link.springer.com/article/10.
1007/s10698-009-9076-7.

Labarca:2009:BRK

[239] Martín Labarca and Olimpia Lombardi. Book review: Klaus Ruthenberg and Jaap van Brakel (eds): Stuff. The nature of chemical substances. Foundations of Chemistry, 11(3):183–186, October 2009. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1007/s10698-009-9077-6.

Scerri:2010:Ea

[240] Eric R. Scerri. Editorial 34. Foundations of Chemistry, 12(1):1–3, April 2010. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
010-9087-4;   http://link.springer.com/content/pdf/10.1007/s10698-
010-9087-4.pdf.

Stewart:2010:CJU

[241] Philip J. Stewart. Charles Janet: unrecognized genius of the periodic system. Foundations of Chemistry, 12(1):5–15, April 2010. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-008-9062-5.

Griffith:2010:GVP

[242] W. P. Griffith. The group VIII platinum-group metals and the periodic table. Foundations of Chemistry, 12(1):17–25, April 2010. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-008-9063-4.

Laing:2010:QMU

[243] M. J. Laing. The question mark at uranium. Foundations of Chemistry, 12(1):27–30, April 2010. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1007/s10698-009-9067-8.

Habashi:2010:MTL

[244] Fathi Habashi. Metals: typical and less typical, transition and inner transition. Foundations of Chemistry, 12(1):31–39, April 2010. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-009-9069-6.

Tobin:2010:MMC

[245] Emma Tobin. Microstructuralism and macromolecules: the case of moonlighting proteins. Foundations of Chemistry, 12(1):41–54, April 2010. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1007/s10698-009-9078-5.

Linthorst:2010:OOD

[246] J. A. Linthorst. An overview: origins and development of green chemistry. Foundations of Chemistry, 12(1):55–68, April 2010. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-009-9079-4;   http:
//link.springer.com/content/pdf/10.1007/s10698-009-9079-4.pdf.

Scerri:2010:EPT

[247] Eric Scerri. Explaining the periodic table, and the role of chemical triads. Foundations of Chemistry, 12(1):69–83, April 2010. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-010-9082-9.

Marks:2010:NRD

[248] E. G. Marks and J. A. Marks. Newlands revisited: a display of the periodicity of the chemical elements for chemists. Foundations of Chemistry, 12 (1):85–93, April 2010. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-010-9083-8;   http://link.springer.com/content/pdf/10.
1007/s10698-010-9083-8.pdf.

Scerri:2010:Eb

[249] Eric Scerri. Editorial 35. Foundations of Chemistry, 12(2):95–96, October 2010. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
010-9095-4;   http://link.springer.com/content/pdf/10.1007/s10698-
010-9095-4.pdf.

Ruthenberg:2010:IFE

[250] K. Ruthenberg. ISPC 2008 first editorial. Foundations of Chemistry, 12 (2):97–99, October 2010. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1007/s10698-010-9096-3;   http://link.springer.com/content/pdf/
10.1007/s10698-010-9096-3.pdf.

Harre:2010:CCC

[251] Rom Harré. Causal concepts in chemical vernaculars. Foundations of Chemistry, 12(2):101–115, October 2010. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1007/s10698-010-9091-8.

Akeroyd:2010:PSM

[252] Michael Akeroyd. The philosophical significance of Mendeleev’s successful predictions of the properties of gallium and scandium. Foundations of Chemistry, 12(2):117–122, October 2010. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1007/s10698-010-9085-6.

vanBrakel:2010:CPN

[253] J. van Brakel. Chemistry and physics: no need for metaphysical glue. Foundations of Chemistry, 12(2):123–136, October 2010. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-010-9084-7.

Soentgen:2010:HPC

[254] Jens Soentgen. On the history and prehistory of CO_{2}. Foundations of Chemistry, 12(2):137–148, October 2010. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1007/s10698-009-9081-x.

Labarca:2010:WOD

[255] Martín Labarca and Olimpia Lombardi. Why orbitals do not exist? Foundations of Chemistry, 12(2):149–157, October 2010. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/article/10.1007/s10698-010-9086-5. See critical note

[335].

Lombardi:2010:MCP

[256] Olimpia Lombardi and Mario Castagnino. Matters are not so clear on the physical side. Foundations of Chemistry, 12(2):159–166, October 2010. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1007/s10698-010-9090-9.

Scerri:2010:Ec

[257] Eric Scerri. Editorial 36. Foundations of Chemistry, 12(3):167–169, November 2010. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-010-9099-0;   http://link.springer.com/content/pdf/10.
1007/s10698-010-9099-0.pdf.

Woods:2010:MMH

[258] Gordon T. Woods. Mendeleev, the man and his matrix: Dmitri Mendeleev, aspects of his life and work: was he a somewhat fortunate man? Foundations of Chemistry, 12(3):171–186, November 2010. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1007/s10698-010-9088-3.

Boyce:2010:BBL

[259] Conal Boyce. On the boundary between laboratory ‘givens’ and laboratory ‘tangibles’. Foundations of Chemistry, 12(3):187–202, November 2010. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
010-9093-6;   http://link.springer.com/content/pdf/10.1007/s10698-
010-9093-6.pdf.

Llored:2010:MQC

[260] Jean-Pierre Llored. Mereology and quantum chemistry: the approximation of molecular orbital. Foundations of Chemistry, 12(3):203–221, November 2010. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-010-9092-7.

Patterson:2010:ALB

[261] Gary D. Patterson. Les atomes: a landmark book in chemistry. Foundations of Chemistry, 12(3):223–233, November 2010. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/article/10.1007/s10698-010-9098-1.

Thyssen:2010:BRE

[262] Pieter Thyssen. Book review: Eric R. Scerri: Selected papers on the periodic table. Foundations of Chemistry, 12(3):235–238, November 2010. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1007/s10698-010-9089-2.

Psarros:2010:BRK

[263] Nikos Psarros. Book review: Klaus Ruthenberg (ed): Frantisek Wald: Essays 1891–1929, Wald Press, Prague, 2009, 231 p. Foundations of Chemistry, 12(3):239–241, November 2010. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1007/s10698-010-9097-2.

Scerri:2011:Ea

[264] Eric R. Scerri. Editorial 37. Foundations of Chemistry, 13(1):1–7, April 2011. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
011-9107-z;   http://link.springer.com/content/pdf/10.1007/s10698-

011-9107-z.pdf. See comment [274].

Anonymous:2011:ISP

[265] Anonymous. International Society for the Philosophy of Chemistry: summer symposium 2011 in Latin America. Foundations of Chemistry, 13 (1):9, April 2011. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/

s10698-011-9109-x;   http://link.springer.com/content/pdf/10.
1007/s10698-011-9109-x.pdf.

Bader:2011:NEP

[266] Richard F. W. Bader. On the non-existence of parallel universes in chemistry. Foundations of Chemistry, 13(1):11–37, April 2011. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-011-9106-0.

Healy:2011:HCL

[267] Eamonn F. Healy. Heisenberg’s chemical legacy: resonance and the chemical bond. Foundations of Chemistry, 13(1):39–49, April 2011. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-011-9104-2.

Pallo:2011:EIQ

[268] Gabor Pallo. Early impact of quantum physics on chemistry: George Hevesy’s work on rare earth elements and Michael Polanyi’s absorption theory. Foundations of Chemistry, 13(1):51–61, April 2011. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-011-9105-1.

Harre:2011:MGC

[269] Rom Harré and Jean-Pierre Llored. Mereologies as the grammars of chemical discourses. Foundations of Chemistry, 13(1):63–76, April 2011. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1007/s10698-011-9103-3.

Laing:2011:BRS

[270] Michael Laing. Book review: Sam Kean: The disappearing spoon: and other true tales of madness, love, and the history of the world from the periodic table of the elements. Foundations of Chemistry, 13(1):77, April 2011. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
010-9101-x;   http://link.springer.com/content/pdf/10.1007/s10698-
010-9101-x.pdf.

Earley:2011:BRA

[271] Joseph E. Earley, Sr. Book review: Alan Chalmers: The scientist’s atom and the philosopher’s stone: how science succeeded and philosophy failed to gain knowledge of atoms. Foundations of Chemistry, 13(1):79–83, April 2011. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
010-9102-9.

Scerri:2011:Eb

[272] Eric R. Scerri. Editorial 38. Foundations of Chemistry, 13(2):85–86, July 2011. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
011-9115-z;   http://link.springer.com/content/pdf/10.1007/s10698-
011-9115-z.pdf.

Ogilvie:2011:MOM

[273] J. F. Ogilvie. Is a molecular orbital measurable by means of tomographic imaging? Foundations of Chemistry, 13(2):87–91, July 2011. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-011-9113-1.

Sutcliffe:2011:CE

[274] Brian T. Sutcliffe and R. Guy Woolley. A comment on “Editorial 37”. Foundations of Chemistry, 13(2):93–95, July 2011. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/article/10.1007/s10698-011-9110-4. See [264].

Vihalemm:2011:ACO

[275] Rein Vihalemm. The autonomy of chemistry: old and new problems. Foundations of Chemistry, 13(2):97–107, July 2011. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/article/10.1007/s10698-010-9094-5.

Kidd:2011:ETK

[276] R. Garth Kidd. Elements of the third kind and the spin-dependent chemical force. Foundations of Chemistry, 13(2):109–119, July 2011. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-010-9100-y.

Rayner-Canham:2011:IPT

[277] Geoff Rayner-Canham. Isodiagonality in the periodic table. Foundations of Chemistry, 13(2):121–129, July 2011. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1007/s10698-011-9108-y.

Pagni:2011:DSR

[278] Richard M. Pagni. Do the solvolysis reactions of secondary substrates occur by the S_{N}1 or S_{N}2 mechanism: or something else? Foundations of Chemistry, 13(2):131–143, July 2011. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1007/s10698-011-9111-3.

Fergusson:2011:HDN

[279] Jack E. Fergusson. The history of the discovery of nuclear fission. Foundations of Chemistry, 13(2):145–166, July 2011. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1007/s10698-011-9112-2.

Hojniak:2011:BRD

[280] Sandra D. Hojniak. Book review: David E. Fisher: Much Ado about (Practically) Nothing. A History of the Noble Gases. Foundations of Chemistry, 13(2):167–169, July 2011. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1007/s10698-011-9114-0.

Scerri:2011:Ec

[281] Eric Scerri. Editorial 39. Foundations of Chemistry, 13(3):171–172, October 2011. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
011-9138-5;   http://link.springer.com/content/pdf/10.1007/s10698-
011-9138-5.pdf.

Banchetti-Robino:2011:OTS

[282] Marina Paola Banchetti-Robino. Ontological tensions in sixteenth and seventeenth century chemistry: between mechanism and vitalism. Foundations of Chemistry, 13(3):173–186, October 2011. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/article/10.1007/s10698-011-9126-9.

Harre:2011:DEF

[283] Rom Harré. Do explanation formats in elementary chemistry depend on agent causality? Foundations of Chemistry, 13(3):187–200, October 2011. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1007/s10698-011-9120-2.

Kim:2011:PCC

[284] Mi Gyung Kim. From phlogiston to caloric: chemical ontologies. Foundations of Chemistry, 13(3):201–222, October 2011. CODEN FOCHFL.

ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/article/10.1007/s10698-011-9116-y.

Rayner-Canham:2011:RAT

[285] Geoff Rayner-Canham. Relationships among the transition elements. Foundations of Chemistry, 13(3):223–232, October 2011. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-011-9117-x.

Restrepo:2011:CLP

[286] Guillermo Restrepo and José L. Villaveces. Chemistry, a lingua philosophica. Foundations of Chemistry, 13(3):233–249, October 2011. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1007/s10698-011-9123-z.

Rayner-Canham:2011:BRA

[287] Marelene Rayner-Canham and Geoff Rayner-Canham. Book review: Anne-Marie Weidler Kubanek: Nothing less than an adventure: Ellen Gleditsch and her life in science. Foundations of Chemistry, 13 (3):251–252, October 2011. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s10698-011-9119-8;   http://link.
springer.com/article/10.1007/s10698-011-9119-8.

Scerri:2012:Ea

[288] Eric R. Scerri. Editorial 40. Foundations of Chemistry, 14(1):1–2, April 2012. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
012-9148-y;   http://link.springer.com/content/pdf/10.1007/s10698-
012-9148-y.pdf.

Chalmers:2012:GE

[289] Alan Chalmers. Guest editorial. Foundations of Chemistry, 14(1):3–6, April 2012. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
012-9147-z;   http://link.springer.com/content/pdf/10.1007/s10698-
012-9147-z.pdf.

Klein:2012:OIC

[290] Ursula Klein. Objects of inquiry in classical chemistry: material substances. Foundations of Chemistry, 14(1):7–23, April 2012. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-011-9122-0.

Lefevre:2012:VCT

[291] Wolfgang Lefèvre. Viewing chemistry through its ways of classifying. Foundations of Chemistry, 14(1):25–36, April 2012. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/article/10.1007/s10698-011-9125-x.

Chalmers:2012:KOC

[292] Alan Chalmers. Klein on the origin of the concept of chemical compound. Foundations of Chemistry, 14(1):37–53, April 2012. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-011-9121-1.

Hendry:2012:CSL

[293] Robin Findlay Hendry. Chemical substances and the limits of pluralism. Foundations of Chemistry, 14(1):55–68, April 2012. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/article/10.1007/s10698-011-9145-6.

Scerri:2012:WEW

[294] Eric R. Scerri. What is an element? What is the periodic table? And what does quantum mechanics contribute to the question? Foundations of Chemistry, 14(1):69–81, April 2012. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1007/s10698-011-9124-y.

Simon:2012:PPP

[295] Jonathan Simon. The production of purity as the production of knowledge. Foundations of Chemistry, 14(1):83–96, April 2012. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-011-9144-7.

Kauffman:2012:BRBa

[296] George B. Kauffman. Book review: Bernadette Bensaude-Vincent and Jonathan Simon: Chemistry, the impure science. Foundations of Chemistry, 14(1):97–98, April 2012. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s10698-011-9132-y;   http://link.
springer.com/article/10.1007/s10698-011-9132-y.

Kauffman:2012:BRI

[297] George B. Kauffman. Book review: István Hargittai: Judging Edward Teller: A closer look at one of the most influential scientists of the twentieth century. Foundations of Chemistry, 14(1):99–101, April 2012. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1007/s10698-011-9133-x.

Anonymous:2012:ISP

[298] Anonymous. International Society for the Philosophy of Chemistry (ISPC) — Summer Symposium 2012. Foundations of Chemistry, 14(1):103–104, April 2012. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s10698-012-9150-4;   http://link.
springer.com/article/10.1007/s10698-012-9150-4.

Anonymous:2012:TIC

[299] Anonymous. Third international conference on the periodic table, Cusco, Peru, 2012. Foundations of Chemistry, 14(1):105–106, April 2012. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/s10698-012-9149-x;   http://link.springer.com/article/10.
1007/s10698-012-9149-x.

Scerri:2012:Eb

[300] Eric Scerri. Editorial 41. Foundations of Chemistry, 14(2):107, July 2012. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
012-9157-x;   http://link.springer.com/content/pdf/10.1007/s10698-

012-9157-x.pdf. See erratum [362].

Thomas:2012:EC

[301] Gerald F. Thomas. The emancipation of chemistry. Foundations of Chemistry, 14(2):109–155, July 2012. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1007/s10698-011-9118-9.

Martins:2012:RMR

[302] Roberto de Andrade Martins. The rise of magnetochemistry from Ritter to Hurmuzescu. Foundations of Chemistry, 14(2):157–182, July 2012. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1007/s10698-011-9127-8.

Kauffman:2012:BRK

[303] George B. Kauffman. Book review: Kenneth J. Klabunde and Ryan M. Richards (Eds): Nanoscale materials in chemistry, 2nd edn. Foundations of Chemistry, 14(2):183–184, July 2012. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s10698-011-9131-z;   http://link.
springer.com/article/10.1007/s10698-011-9131-z.

Kauffman:2012:BRA

[304] George B. Kauffman. Book review: Alexander Y. Grosberg and Alexei R. Khokhlov: Giant molecules: here, there, and everywhere, 2nd edn. Foundations of Chemistry, 14(2):185–186, July 2012. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10698-011-9134-9;
http://link.springer.com/article/10.1007/s10698-011-9134-9.

Kauffman:2012:BRBb

[305] George B. Kauffman. Book review: Bob B. He: Two-dimensional X-ray diffraction. Foundations of Chemistry, 14(2):187–188, July 2012. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/s10698-011-9135-8;   http://link.springer.com/article/10.
1007/s10698-011-9135-8.

Scerri:2012:Ec

[306] Eric Scerri. Editorial 42. Foundations of Chemistry, 14(3):189–190, October 2012. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/s10698-012-9167-8;   http://link.springer.com/article/10.
1007/s10698-012-9167-8.

Harre:2012:I

[307] Rom Harré. Introduction. Foundations of Chemistry, 14(3):191, October 2012. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
012-9162-0;   http://link.springer.com/content/pdf/10.1007/s10698-
012-9162-0.pdf.

Ruthenberg:2012:PCI

[308] Klaus Ruthenberg and Rom Harré. Philosophy of chemistry as intercultural philosophy: Jaap van Brakel. Foundations of Chemistry, 14(3): 193–203, October 2012. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-012-9161-1.

Lombardi:2012:PMV

[309] Olimpia Lombardi. Prigogine and the many voices of nature. Foundations of Chemistry, 14(3):205–219, October 2012. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1007/s10698-011-9140-y.

Banchetti-Robino:2012:OFF

[310] Marina Paola Banchetti-Robino. The ontological function of first-order and second-order corpuscles in the chemical philosophy of Robert Boyle: the redintegration of potassium nitrate. Foundations of Chemistry, 14 (3):221–234, October 2012. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1007/s10698-012-9159-8.

Earley:2012:NAP

[311] Joseph E. Earley, Sr. A neglected aspect of the puzzle of chemical structure: how history helps. Foundations of Chemistry, 14(3):235–243, October 2012. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
012-9146-0.

Llored:2012:EQC

[312] Jean-Pierre Llored. Emergence and quantum chemistry. Foundations of Chemistry, 14(3):245–274, October 2012. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1007/s10698-012-9163-z.

Scerri:2012:CWV

[313] Eric R. Scerri. A critique of Weisberg’s view on the periodic table and some speculations on the nature of classifications. Foundations of Chemistry, 14 (3):275–284, October 2012. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1007/s10698-012-9164-y.

Scerri:2013:Ea

[314] Eric Scerri. Editorial 43. Foundations of Chemistry, 15(1):1–2, April 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
013-9188-y;   http://link.springer.com/content/pdf/10.1007/s10698-
013-9188-y.pdf.

Nelson:2013:WM

[315] Peter G. Nelson. What is the mole? Foundations of Chemistry, 15(1): 3–11, April 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-013-9185-1.

Leach:2013:CEB

[316] Mark R. Leach. Concerning electronegativity as a basic elemental property and why the periodic table is usually represented in its medium form. Foundations of Chemistry, 15(1):13–29, April 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-012-9151-3.

Marchese:2013:PVD

[317] Francis T. Marchese. Periodicity, visualization, and design. Foundations of Chemistry, 15(1):31–55, April 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1007/s10698-012-9154-0.

Sima:2013:RRI

[318] Jozef Sima. Redox reactions: inconsistencies in their description. Foundations of Chemistry, 15(1):57–64, April 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1007/s10698-011-9143-8.

Lewowicz:2013:SVI

[319] Lucía Lewowicz and Olimpia Lombardi. Stuff versus individuals. Foundations of Chemistry, 15(1):65–77, April 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1007/s10698-012-9152-2.

Sarikaya:2013:VAS

[320] Mustafa Sarikaya. A view about the short histories of the mole and Avogadro’s number. Foundations of Chemistry, 15(1):79–91, April 2013. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1007/s10698-011-9128-7.

Blumenthal:2013:KCR

[321] Geoffrey Blumenthal. Kuhn and the chemical revolution: a re-assessment. Foundations of Chemistry, 15(1):93–101, April 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/article/10.1007/s10698-011-9129-6.

Gomez:2013:SCE

[322] Pedro J. Sánchez Gómez. The semantics of chemical education: constructivism, externalism and the language of chemistry. Foundations of Chemistry, 15(1):103–116, April 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1007/s10698-011-9130-0.

Kauffman:2013:BRR

[323] George B. Kauffman. Book review: Richard P. Pohanish and Stanley A. Greene: Wiley guide to chemical incompatibilities, 3rd edn. Foundations of Chemistry, 15(1):117, April 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-011-9136-7;   http:
//link.springer.com/content/pdf/10.1007/s10698-011-9136-7.pdf.

Kauffman:2013:BRB

[324] George B. Kauffman. Book review: Bassam Z. Shakhashiri: Chemical demonstrations: a handbook for teachers of chemistry, Volume 5. Foundations of Chemistry, 15(1):119–120, April 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10698-011-9137-6;
http://link.springer.com/article/10.1007/s10698-011-9137-6.

Kauffman:2013:BRT

[325] George B. Kauffman. Book review: Thomas B. Rauchfuss, Editorin-Chief: Inorganic Syntheses, Volume 35. Foundations of Chemistry, 15(1):121–122, April 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s10698-011-9139-4;   http://link.
springer.com/article/10.1007/s10698-011-9139-4.

Scerri:2013:Eb

[326] Eric Scerri. Editorial 44. Foundations of Chemistry, 15(2):123, July 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
013-9191-3;   http://link.springer.com/content/pdf/10.1007/s10698-
013-9191-3.pdf.

Restrepo:2013:E

[327] Guillermo Restrepo. Editorial. Foundations of Chemistry, 15(2):125–126, July 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-

013-9180-6;   http://link.springer.com/content/pdf/10.1007/s10698-
013-9180-6.pdf.

Harre:2013:MM

[328] Rom Harré and Jean-Pierre Llored. Molecules and mereology. Foundations of Chemistry, 15(2):127–144, July 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1007/s10698-013-9181-5.

Rocke:2013:WDT

[329] Alan Rocke. What did “theory” mean to nineteenth-century chemists? Foundations of Chemistry, 15(2):145–156, July 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/article/10.1007/s10698-013-9184-2. See comments

[359] and response [360].

Chamizo:2013:TOC

[330] José Antonio Chamizo. Technochemistry: One of the chemists’ ways of knowing. Foundations of Chemistry, 15(2):157–170, July 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-013-9179-z.

Mahootian:2013:PEC

[331] Farzad Mahootian. Paneth’s epistemology of chemical elements in light of Kant’s Opus postumum. Foundations of Chemistry, 15(2):171–184, July 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
013-9182-4.

Restrepo:2013:MMC

[332] Guillermo Restrepo. To mathematize, or not to mathematize chemistry. Foundations of Chemistry, 15(2):185–197, July 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-013-9183-3.

Nelson:2013:PFC

[333] Peter G. Nelson. Periodicity in the formulae of carbonyls and the electronic basis of the periodic table. Foundations of Chemistry, 15(2):199–208, July 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
012-9155-z.

Gilead:2013:STQ

[334] Amihud Gilead. Shechtman’s three question marks: possibility, impossibility, and quasicrystals. Foundations of Chemistry, 15(2):209–224, July 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
012-9156-y.

Manafu:2013:IRP

[335] Alexandru Manafu. Internal realism and the problem of ontological autonomy: a critical note on Lombardi and Labarca. Foundations of Chemistry, 15(2):225–228, July 2013. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1007/s10698-012-9165-x. See [255].

Rayner-Canham:2013:PPG

[336] Geoff Rayner-Canham. Periodic patterns: the group (n) and group (n+10) linkage. Foundations of Chemistry, 15(2):229–237, July 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-012-9169-6.

Kauffman:2013:BRG

[337] George B. Kauffman. Book review: George A. Olah, Alain Goeppert and G. K. Surya Prakash (eds): Beyond oil and gas: the methanol economy, 2nd updated and enlarged edition. Foundations of Chemistry, 15(2):239–240, July 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s10698-011-9141-x;   http://link.
springer.com/article/10.1007/s10698-011-9141-x.

Kauffman:2013:BRH

[338] George B. Kauffman. Book review: Herbert W. Roesky and Dietmar K. Kennepohl (eds): Experiments in green and sustainable chemistry. Foundations of Chemistry, 15(2):241–242, July 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10698-011-9142-9;
http://link.springer.com/article/10.1007/s10698-011-9142-9.

Scerri:2013:Ec

[339] Eric Scerri. Editorial 45. Foundations of Chemistry, 15(3):243, October 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
013-9195-z;   http://link.springer.com/content/pdf/10.1007/s10698-
013-9195-z.pdf.

Matta:2013:SIP

[340] Chérif F. Matta. Special issue: Philosophical aspects and implications of the quantum theory of atoms in molecules (QTAIM). Foundations of Chemistry, 15(3):245–251, October 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-013-9194-0;   http:
//link.springer.com/content/pdf/10.1007/s10698-013-9194-0.pdf.

Bader:2013:AMN

[341] Richard F. W. Bader and Chérif F. Matta. Atoms in molecules as nonoverlapping, bounded, space-filling open quantum systems. Foundations of Chemistry, 15(3):253–276, October 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1007/s10698-012-9153-1.

Eberhart:2013:TFC

[342] Mark E. Eberhart and Travis E. Jones. The two faces of chemistry: can they be reconciled? Foundations of Chemistry, 15(3):277–285, October 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
012-9172-y.

Shahbazian:2013:BOQ

[343] Shant Shahbazian. Beyond the orthodox QTAIM: motivations, current status, prospects and challenges. Foundations of Chemistry, 15(3):287– 302, October 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-012-9170-0.

Sukumar:2013:AMM

[344] N. Sukumar. The atom in a molecule as a mereological construct in chemistry. Foundations of Chemistry, 15(3):303–309, October 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-012-9171-z.

Hettema:2013:AQM

[345] Hinne Hettema. Austere quantum mechanics as a reductive basis for chemistry. Foundations of Chemistry, 15(3):311–326, October 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http:/

/link.springer.com/article/10.1007/s10698-012-9173-x. See com-

ment [346].

Shahbazian:2013:CAQ

[346] Shant Shahbazian. Comment on “Austere quantum mechanics as a reductive basis for chemistry”. Foundations of Chemistry, 15(3):327–334, October 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-

012-9177-6. See [345] and reply [347].

Hettema:2013:QRP

[347] Hinne Hettema. QTAIM as a research programme: a reply to Shahbazian. Foundations of Chemistry, 15(3):335–341, October 2013. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-013-9193-1. See [346].

Scerri:2014:Ea

[348] Eric R. Scerri. Editorial 46. Foundations of Chemistry, 16(1):1–2, April 2014. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
014-9202-z;   http://link.springer.com/content/pdf/10.1007/s10698-
014-9202-z.pdf.

Causa:2014:ABM

[349] Mauro Causá, Andreas Savin, and Bernard Silvi. Atoms and bonds in molecules and chemical explanations. Foundations of Chemistry, 16(1): 3–26, April 2014. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-013-9192-2.

Nalewajski:2014:ECE

[350] Roman F. Nalewajski. Entropic concepts in electronic structure theory. Foundations of Chemistry, 16(1):27–62, April 2014. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-012-9168-7;   http:
//link.springer.com/content/pdf/10.1007/s10698-012-9168-7.pdf.

Joypazadeh:2014:WDS

[351] Hamidreza Joypazadeh and Shant Shahbazian. What does shape a topological atom? Foundations of Chemistry, 16(1):63–75, April 2014. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1007/s10698-013-9186-0.

Shahbazian:2014:LET

[352] Shant Shahbazian. Letter to the editor: Are there “really” atoms in molecules? Foundations of Chemistry, 16(1):77–84, April 2014. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-013-9187-z.

Scerri:2014:Eb

[353] Eric Scerri. Editorial 47. Foundations of Chemistry, 16(2):85–86, July 2014. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
014-9210-z;   http://link.springer.com/content/pdf/10.1007/s10698-
014-9210-z.pdf.

Ramirez-Solis:2014:FMM

[354] Alejandro Ramírez-Solís and Octavio Novaro. The first metals in Mendeleiev’s table: further arguments to place He above Ne and not above be. Foundations of Chemistry, 16(2):87–91, July 2014. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-012-9174-9.

Boyce:2014:ULD

[355] Conal Boyce. Using logic to define the Aufbau–Hund–Pauli relation: a guide to teaching orbitals as a single, natural, unfragmented ruleset. Foundations of Chemistry, 16(2):93–106, July 2014. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-012-9176-7;   http:
//link.springer.com/content/pdf/10.1007/s10698-012-9176-7.pdf.

Villani:2014:SSC

[356] Giovanni Villani. Structured system in chemistry: comparison with mechanics and biology. Foundations of Chemistry, 16(2):107–123, July 2014. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1007/s10698-013-9178-0.

Marques:2014:ESI

[357] Carlos Alberto Marques and Adélio A. S. C. Machado. Environmental sustainability: implications and limitations to green chemistry. Foundations of Chemistry, 16(2):125–147, July 2014. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1007/s10698-013-9189-x.

Gilead:2014:PPS

[358] Amihud Gilead. Pure possibilities and some striking scientific discoveries. Foundations of Chemistry, 16(2):149–163, July 2014. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-013-9190-4.

Needham:2014:NCC

[359] Paul Needham. Nineteenth-century chemical theory: correction of a misunderstanding. Foundations of Chemistry, 16(2):165–167, July 2014. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-

013-9198-9. See [329] and response [360].

Rocke:2014:RLE

[360] Alan Rocke. Response to a letter to the editor. Foundations of Chemistry, 16(2):169–170, July 2014. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s10698-013-9199-8;   http://link.
springer.com/article/10.1007/s10698-013-9199-8. See [359, 329].

Kauffman:2014:BREa

[361] George B. Kauffman. Book review: Eric R. Scerri: The periodic table: a very short introduction. Foundations of Chemistry, 16(2):171–172, July 2014. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s10698-014-9206-8;   http://link.
springer.com/article/10.1007/s10698-014-9206-8.

Scerri:2014:EE

[362] Eric Scerri. Erratum to: Editorial 41. Foundations of Chemistry, 16 (2):173, July 2014. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-012-9166-9;   http://link.springer.com/content/pdf/10.
1007/s10698-012-9166-9.pdf. See [300].

Scerri:2014:Ec

[363] Eric R. Scerri. Editorial 48. Foundations of Chemistry, 16(3):175, October 2014. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
014-9213-9;   http://link.springer.com/content/pdf/10.1007/s10698-
014-9213-9.pdf.

Hernandez:2014:FMM

[364] Raymundo Hernández and Octavio Novaro. The first metals in Mendeleiev’s table: Part II. A new argument against the placement of hydrogen atop the alkali metal column. Foundations of Chemistry, 16 (3):177–180, October 2014. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1007/s10698-013-9196-y.

Lombardi:2014:LCP

[365] Olimpia Lombardi. Linking chemistry with physics: arguments and counterarguments. Foundations of Chemistry, 16(3):181–192, October 2014. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-

013-9197-x. See reply [366].

Hettema:2014:LCP

[366] Hinne Hettema. Linking chemistry with physics: a reply to Lombardi. Foundations of Chemistry, 16(3):193–200, October 2014. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-014-9200-1. See [365].

Gilead:2014:CRI

[367] Amihud Gilead. Chain reactions, “impossible” reactions, and panenmentalist possibilities. Foundations of Chemistry, 16(3):201–214, October 2014. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
014-9201-0.

Bengoetxea:2014:SRC

[368] Juan Bautista Bengoetxea, Oliver Todt, and José Luis Luján. Similarity and representation in chemical knowledge practices. Foundations of Chemistry, 16(3):215–233, October 2014. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1007/s10698-014-9203-y.

Imyanitov:2014:ANF

[369] Naum S. Imyanitov. Adequacy of the new formulation of the periodic law when fundamental variations occur in blocks and periods. Foundations of Chemistry, 16(3):235–247, October 2014. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1007/s10698-014-9204-x.

Kauffman:2014:BRU

[370] George B. Kauffman. Book review: Ulf Lagerkvist: Erling Norrby (ed.): The periodic table and a missed Nobel Prize. Foundations of Chemistry, 16(3):249–251, October 2014. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1007/s10698-014-9205-9.

Kauffman:2014:BREb

[371] George B. Kauffman. Book review: Eric Scerri: A tale of 7 elements. Foundations of Chemistry, 16(3):253–256, October 2014. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-014-9207-7.

Kauffman:2014:BREc

[372] George B. Kauffman. Book review: Eric Scerri (ed): 30-Second elements: the 50 most significant elements, each explained in half a minute. Foundations of Chemistry, 16(3):257–258, October 2014. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10698-014-9208-6;
http://link.springer.com/article/10.1007/s10698-014-9208-6.

Scerri:2015:Ea

[373] Eric Scerri. Editorial 49. Foundations of Chemistry, 17(1):1–3, April 2015. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
015-9230-3;   http://link.springer.com/content/pdf/10.1007/s10698-
015-9230-3.pdf.

Maar:2015:BCA

[374] Juergen Heinrich Maar and Eder João Lenardão. The Brazilian contribution of Alcindo Flores Cabral to the periodic classification. Foundations of Chemistry, 17(1):5–22, April 2015. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1007/s10698-015-9223-2.

Jensen:2015:PLA

[375] William B. Jensen. The positions of lanthanum (actinium) and lutetium (lawrencium) in the periodic table: an update. Foundations of Chemistry, 17(1):23–31, April 2015. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1007/s10698-015-9216-1.

deBerg:2015:FCE

[376] Kevin Charles de Berg. Foundations of and challenges to electrolyte chemistry. Foundations of Chemistry, 17(1):33–48, April 2015. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-015-9219-y.

Garcia:2015:CSE

[377] Pio Garcia. Computer simulations and experiments: in vivo-in vitro conditions in biochemistry. Foundations of Chemistry, 17(1):49–65, April 2015. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1007/s10698-015-9215-2.

DiGiacomo:2015:SCM

[378] Francesco Di Giacomo. On some characteristics of Marcus’ work in the light of the history of science. Foundations of Chemistry, 17(1):67–78, April 2015. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
015-9222-3.

Rulev:2015:PNB

[379] Alexander Yu. Rulev. Pegniochemistry as a new branch of the chemical science. Foundations of Chemistry, 17(1):79–86, April 2015. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-015-9214-3.

Pallo:2015:BRJ

[380] Gábor Palló. Book review: Jean-Pierre Llored (ed.): The philosophy of chemistry: practices, methodologies, and concepts. Foundations of Chemistry, 17(1):87–89, April 2015. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1007/s10698-014-9209-5.

Scerri:2015:Eb

[381] Eric Scerri. Editorial 50. Foundations of Chemistry, 17(2):91–92, July 2015. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
015-9232-1;   http://link.springer.com/content/pdf/10.1007/s10698-
015-9232-1.pdf.

Barragan:2015:EKT

[382] Daniel Barragán. Essentials of kinetics and thermodynamics for understanding chemical oscillations. Foundations of Chemistry, 17(2):93–106, July 2015. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
015-9221-4.

Moretti:2015:ERP

[383] Giuliano Moretti. The ”extent of reaction”: a powerful concept to study chemical transformations at the first-year general chemistry courses. Foundations of Chemistry, 17(2):107–115, July 2015. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1007/s10698-014-9212-x.

Gambini:2015:QMS

[384] Rodolfo Gambini, Lucía Lewowicz, and Jorge Pullin. Quantum mechanics, strong emergence and ontological non-reducibility. Foundations of Chemistry, 17(2):117–127, July 2015. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1007/s10698-015-9224-1.

Khalilov:2015:SIC

[385] Leonard M. Khalilov. Symmetry, inertness and chirality in theory of chiral systems. Foundations of Chemistry, 17(2):129–135, July 2015. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-015-9231-2.

Best:2015:LRP

[386] Nicholas W. Best. Lavoisier’s ”reflections on phlogiston” I: against phlogiston theory. Foundations of Chemistry, 17(2):137–151, July 2015. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1007/s10698-015-9220-5.

Nelson:2015:MVL

[387] Peter G. Nelson. A modern version of Lewis’s theory of valency. Foundations of Chemistry, 17(2):153–162, July 2015. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1007/s10698-015-9217-0. See erratum [404].

Tangde:2015:AHC

[388] Vijay M. Tangde. Advances in hadronic chemistry and its applications. Foundations of Chemistry, 17(2):163–179, July 2015. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-015-9218-z.

Scerri:2015:Ec

[389] Eric Scerri. Editorial 51. Foundations of Chemistry, 17(3):181, October 2015. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
015-9239-7;   http://link.springer.com/content/pdf/10.1007/s10698-
015-9239-7.pdf.

Llored:2015:GE

[390] Jean-Pierre Llored. Guest editorial. Foundations of Chemistry, 17(3): 183–185, October 2015. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-015-9240-1;   http://link.springer.com/content/pdf/10.
1007/s10698-015-9240-1.pdf.

Schaming:2015:NAT

[391] Delphine Schaming and Hynd Remita. Nanotechnology: from the ancient time to nowadays. Foundations of Chemistry, 17(3):187–205, October 2015. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
015-9235-y.

Laurent:2015:OPE

[392] Brice Laurent. Operations for a problem of existence: dealing with the ontological uncertainty of nano substances. Foundations of Chemistry, 17 (3):207–224, October 2015. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1007/s10698-015-9233-0.

Lacour:2015:DRC

[393] Stéphanie Lacour, Sacha Loeve, Brice Laurent, and Virginie Albe. Deliberating responsibility: a collective contribution by the C’Nano IdF Nanoscience & Society Office. Foundations of Chemistry, 17(3):225– 245, October 2015. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-015-9234-z.

Boeyens:2015:WMM

[394] Jan C. A. Boeyens. Wave-mechanical model for chemistry. Foundations of Chemistry, 17(3):247–262, October 2015. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1007/s10698-015-9225-0.

Hartman:2015:DWN

[395] JudithAnn R. Hartman and Eric A. Nelson. “Do we need to memorize that?” or cognitive science for chemists. Foundations of Chemistry, 17 (3):263–274, October 2015. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL http://link.springer.com/article/10.
1007/s10698-015-9226-z.

Downward:2015:SRT

[396] Michael Downward. Symmetry and representation in a three dimensional space. Foundations of Chemistry, 17(3):275–287, October 2015. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-015-9227-y.

Scerri:2016:Ea

[397] Eric Scerri. Editorial 52. Foundations of Chemistry, 18(1):1–2, April 2016. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
016-9254-3;   http://link.springer.com/content/pdf/10.1007/s10698-
016-9254-3.pdf.

Best:2016:LRP

[398] Nicholas W. Best. Lavoisier’s ”reflections on phlogiston” II: on the nature of heat. Foundations of Chemistry, 18(1):3–13, April 2016. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-015-9236-x.

Nelson:2016:ITR

[399] Peter G. Nelson. Introducing the theory of relativity. Foundations of Chemistry, 18(1):15–19, April 2016. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1007/s10698-015-9247-7;   http://link.springer.com/
content/pdf/10.1007/s10698-015-9247-7.pdf.

Imyanitov:2016:DSC

[400] Naum S. Imyanitov. Dialectics and synergetics in chemistry. Periodic Table and oscillating reactions. Foundations of Chemistry, 18(1):21–56, April 2016. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
015-9248-6.

Salehzadeh:2016:SWO

[401] Sadegh Salehzadeh and Farahnaz Maleki. The 4s and 3d subshells: Which one fills first in progressing through the periodic table and which one fills first in any particular atom? Foundations of Chemistry, 18(1):57–65, April 2016. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
016-9249-0.

Sima:2016:SRM

[402] Jozef Sima. Structure-related melting and boiling points of inorganic compounds. Foundations of Chemistry, 18(1):67–79, April 2016. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-015-9228-x.

Llored:2016:KGA

[403] Jean-Pierre Llored. Book review: Kostas Gavroglu and Ana Simões: Neither physics nor chemistry. A history of quantum chemistry. Foundations of Chemistry, 18(1):81–84, April 2016. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1007/s10698-014-9211-y.

Nelson:2016:EMV

[404] Peter G. Nelson. Erratum to: A modern version of Lewis’s theory of valency. Foundations of Chemistry, 18(1):85, April 2016. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-015-9238-8;   http:
//link.springer.com/content/pdf/10.1007/s10698-015-9238-8.pdf.

See [387].

Scerri:2016:Eb

[405] Eric R. Scerri. Editorial 53. Foundations of Chemistry, 18(2):87, July 2016. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
016-9256-1;   http://link.springer.com/content/pdf/10.1007/s10698-
016-9256-1.pdf.

Harre:2016:RTP

[406] Rom Harré. Realism and the turn to practice. Foundations of Chemistry, 18(2):89, July 2016. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-015-9245-9;   http://link.springer.com/content/pdf/10.
1007/s10698-015-9245-9.pdf.

Vihalemm:2016:CPP

[407] Rein Vihalemm. Chemistry and the problem of pluralism in science: an analysis concerning philosophical and scientific disagreements. Foundations of Chemistry, 18(2):91–102, July 2016. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.

com/article/10.1007/s10698-015-9241-0.

Banchetti-Robino:2016:VHH

[408] Marina Paola Banchetti-Robino. Van Helmont’s hybrid ontology and its influence on the chemical interpretation of spirit and ferment. Foundations of Chemistry, 18(2):103–112, July 2016. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1007/s10698-015-9243-y.

Muursepp:2016:CPSa

[409] Peeter Müürsepp. Chemistry as a practical science: Edward Caldin revisited. Foundations of Chemistry, 18(2):113–123, July 2016. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/article/10.1007/s10698-015-9244-x.

Llored:2016:CPC

[410] Jean-Pierre Llored and Stéphane Sarrade. Connecting the philosophy of chemistry, green chemistry, and moral philosophy. Foundations of Chemistry, 18(2):125–152, July 2016. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

article/10.1007/s10698-015-9242-z.

Imyanitov:2016:SFG

[411] Naum S. Imyanitov. Spiral as the fundamental graphic representation of the Periodic Law. Blocks of elements as the autonomic parts of the Periodic System. Foundations of Chemistry, 18(2):153–173, July 2016. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

http://link.springer.com/article/10.1007/s10698-015-9246-8.

Kauffman:2016:AHZ

[412] George B. Kauffman and Laurie M. Kauffman. Ahmed H. Zewail: 4D visualization of matter: recent collected works. Foundations of Chemistry, 18(2):175–176, July 2016. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s10698-015-9237-9;   http://link.
springer.com/article/10.1007/s10698-015-9237-9.

Scerri:2016:Ec

[413] Eric Scerri. Editorial 54. Foundations of Chemistry, 18(3):177–178, October 2016. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
016-9261-4;   http://link.springer.com/content/pdf/10.1007/s10698-
016-9261-4.pdf.

Siderer:2016:MPM

[414] Yona Siderer. In memory of Professor Masanori Kaji (8 February 1956–18 July 2016). Foundations of Chemistry, 18(3):179–181, October 2016. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/s10698-016-9262-3;   http://link.springer.com/article/10.
1007/s10698-016-9262-3.

Gilead:2016:EEC

[415] Amihud Gilead. Eka-elements as chemical pure possibilities. Foundations of Chemistry, 18(3):183–194, October 2016. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10698-016-9250-7;
http://link.springer.com/article/10.1007/s10698-016-9250-7.

Weinstein:2016:PTM

[416] Mark Weinstein. The periodic table and the model of emerging truth. Foundations of Chemistry, 18(3):195–212, October 2016. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s10698-016-9252-
5;   http://link.springer.com/article/10.1007/s10698-016-9252-
5.

Muursepp:2016:CPSb

[417] Peeter Müürsepp. Chemistry as a practical science (Edward Caldin revisited). Foundations of Chemistry, 18(3):213–223, October 2016. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s10698-016-9257-
0;   http://link.springer.com/article/10.1007/s10698-016-9257-
0.

Fortin:2016:ID

[418] Sebastian Fortin, Olimpia Lombardi, and Juan Camilo Martínez González. Isomerism and decoherence. Foundations of Chemistry, 18(3):225–240, October 2016. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s10698-016-9251-6;   http://link.
springer.com/article/10.1007/s10698-016-9251-6.

Fisher:2016:DCO

[419] Grant Fisher. Diagnostics in computational organic chemistry. Foundations of Chemistry, 18(3):241–262, October 2016. CODEN FOCHFL.

ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10698-016-9253-4;
http://link.springer.com/article/10.1007/s10698-016-9253-4.

Scerri:2017:Ea

[420] Eric Scerri. Editorial 55. Foundations of Chemistry, 19(1):1, April 2017. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10698-
017-9279-2;   http://link.springer.com/content/pdf/10.1007/s10698-
017-9279-2.pdf.

Banchetti-Robino:2017:GEa

[421] Marina P. Banchetti-Robino. Guest editorial. Foundations of Chemistry, 19(1):3–4, April 2017. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL http://link.springer.com/article/10.1007/
s10698-017-9276-5;   http://link.springer.com/content/pdf/10.
1007/s10698-017-9276-5.pdf.

DeLanda:2017:RHC

[422] Manuel DeLanda. Realism and the history of chemistry. Foundations of Chemistry, 19(1):5–15, April 2017. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s10698-017-9274-7;   http://link.
springer.com/article/10.1007/s10698-017-9274-7.

Fisher:2017:CDR

[423] Grant Fisher. Content, design, and representation in chemistry. Foundations of Chemistry, 19(1):17–28, April 2017. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10698-017-9275-6;
http://link.springer.com/article/10.1007/s10698-017-9275-6.

Prentner:2017:CCO

[424] Robert Prentner. Chemistry, context and the objects of thought. Foundations of Chemistry, 19(1):29–41, April 2017. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10698-017-9273-8;
http://link.springer.com/article/10.1007/s10698-017-9273-8.

Fortin:2017:RBC

[425] Sebastian Fortin, Olimpia Lombardi, and Juan Camilo Martínez González. The relationship between chemistry and physics from the perspective of Bohmian mechanics. Foundations of Chemistry, 19(1):43–59, April 2017. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s10698-017-9277-4;   http://link.
springer.com/article/10.1007/s10698-017-9277-4.

Ruthenberg:2017:EMF

[426] Klaus Ruthenberg and Juan Camilo Martínez González. Electronegativity and its multiple faces: persistence and measurement. Foundations of Chemistry, 19(1):61–75, April 2017. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s10698-017-9278-3;   http://link.
springer.com/article/10.1007/s10698-017-9278-3.

Villani:2017:CPS

[427] Giovanni Villani. Chemical perspective in the study of living beings: a systemic complexity approach. Foundations of Chemistry, 19(1):77–91, April 2017. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s10698-016-9258-z;   http://link.
springer.com/article/10.1007/s10698-016-9258-z.

Black:2017:BRP

[428] Roderick S. Black. Book review: Peter J. T. Morris: The Matter Factory: A History of the Chemistry Laboratory. Foundations of Chemistry, 19(1):93–94, April 2017. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s10698-016-9255-2;   http://link.
springer.com/article/10.1007/s10698-016-9255-2.

Scerri:2017:Eb

[429] Eric Scerri. Editorial 56. Foundations of Chemistry, 19(2):95–96, July 2017. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1007/
s10698-017-9288-1.pdf.

Ghibaudi:2017:CSM

[430] Elena Ghibaudi and Luigi Cerruti. Chemical substance, material, product, goods, waste: a changing ontology. Foundations of Chemistry, 19(2):97– 123, July 2017. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic).

Francisco:2017:CBB

[431] Nuno Francisco, Carla Morais, João C. Paiva, and Paula Gameiro. A colourful bond between art and chemistry. Foundations of Chemistry, 19(2):125–138, July 2017. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic).

Vitas:2017:QRT

[432] Marko Vitas and Andrej Dobovisek. On a quest of reverse translation. Foundations of Chemistry, 19(2):139–155, July 2017. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic).

Chamizo:2017:FCR

[433] José A. Chamizo. The fifth chemical revolution: 1973–1999. Foundations of Chemistry, 19(2):157–179, July 2017. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic).

Scerri:2017:Ec

[434] Eric Scerri. Editorial 57. Foundations of Chemistry, 19(3):181, October 2017. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL https://link.springer.com/content/pdf/10.
1007/s10698-017-9294-3.pdf.

Banchetti-Robino:2017:GEb

[435] Marina P. Banchetti-Robino. Guest editorial. Foundations of Chemistry, 19(3):183–184, October 2017. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

content/pdf/10.1007/s10698-017-9286-3.pdf.

Cordoba:2017:HHN

[436] Mariana Córdoba and Alfio Zambon. How to handle nanomaterials? The re-entry of individuals into the philosophy of chemistry. Foundations of Chemistry, 19(3):185–196, October 2017. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic).

Ochiai:2017:DMS

[437] Hirofumi Ochiai. Does a molecule have structure? Foundations of Chemistry, 19(3):197–207, October 2017. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic).

Seifert:2017:AAU

[438] Vanessa A. Seifert. An alternative approach to unifying chemistry with quantum mechanics. Foundations of Chemistry, 19(3):209–222, October 2017. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-

8463 (electronic). URL https://link.springer.com/content/pdf/10.
1007/s10698-017-9285-4.pdf.

Siderer:2017:UYT

[439] Yona Siderer. Udagawa Youan’s (1798–1846) translation of light and heat reactions in his book Kouso Seimika. Foundations of Chemistry, 19(3): 223–240, October 2017. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic).

Blumenthal:2017:DPW

[440] Geoffrey Blumenthal and James Ladyman. The development of problems within the phlogiston theories, 1766–1791. Foundations of Chemistry, 19(3):241–280, October 2017. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

content/pdf/10.1007/s10698-017-9289-0.pdf.

Wray:2017:NPS

[441] K. Brad Wray. A new philosophy of science from the history of arcane natural science. Foundations of Chemistry, 19(3):281–285, October 2017. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic).

Scerri:2018:Ea

[442] Eric Scerri. Editorial 58. Foundations of Chemistry, 20(1):1–2, April 2018. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
018-9309-8;   https://link.springer.com/content/pdf/10.1007/
s10698-018-9309-8.pdf.

Dolino:2018:CCS

[443] Le Grande O. Dolino. Chemistry as a creative science. Foundations of Chemistry, 20(1):3–13, April 2018. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-017-9272-9.

Johnson:2018:VL

[444] David A. Johnson and Peter G. Nelson. Valencies of the lanthanides. Foundations of Chemistry, 20(1):15–27, April 2018. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

https://link.springer.com/article/10.1007/s10698-017-9291-6;
 https://link.springer.com/content/pdf/10.1007/s10698-017-9291-

6.pdf. See erratum [445] and correction [467].

Johnson:2018:EVL

[445] David A. Johnson and Peter G. Nelson. Erratum to: Valencies of the lanthanides. Foundations of Chemistry, 20(1):29–30, April 2018. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
017-9296-1;   https://link.springer.com/content/pdf/10.1007/
s10698-017-9296-1.pdf. See [444].

Iga:2018:BPS

[446] Dumitru Petru Iga. Basic principles of the strategy concerning the elucidation of configuration of chiral centers of linear isomeric aldohexoses. Foundations of Chemistry, 20(1):31–41, April 2018. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.

com/article/10.1007/s10698-017-9292-5.

Hefferlin:2018:AMP

[447] Ray Hefferlin. An allegory on molecular periodicity. Foundations of Chemistry, 20(1):43–49, April 2018. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-017-9295-2.

Zambon:2018:RPS

[448] Alfio Zambon. A representation of the periodic system based on atomicnumber triads. Foundations of Chemistry, 20(1):51–74, April 2018. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

https://link.springer.com/article/10.1007/s10698-017-9297-0.

Seeman:2018:MBI

[449] Jeffrey I. Seeman. Moving beyond insularity in the history, philosophy, and sociology of chemistry. Foundations of Chemistry, 20(1):75–86, April 2018. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
017-9290-7;   https://link.springer.com/content/pdf/10.1007/
s10698-017-9290-7.pdf. See response [461].

Scerri:2018:Eb

[450] Eric Scerri. Editorial 59. Foundations of Chemistry, 20(2):87–88, July 2018. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
018-9317-8;   https://link.springer.com/content/pdf/10.1007/
s10698-018-9317-8.pdf.

Villani:2018:OPC

[451] Giovanni Villani, Elena Ghibaudi, and Luigi Cerruti. The orbital: a pivotal concept in the relationship between chemistry and physics? A comment to the work by Fortin and coauthors. Foundations of Chemistry, 20(2):89–97, July 2018. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-017-9293-4.

Hodder:2018:PTR

[452] Peter Hodder. The periodic table: revelation by quest rather than by revolution. Foundations of Chemistry, 20(2):99–110, July 2018. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https:/

/link.springer.com/article/10.1007/s10698-017-9298-z.

Stewart:2018:TSR

[453] Philip J. Stewart. Tetrahedral and spherical representations of the periodic system. Foundations of Chemistry, 20(2):111–120, July 2018. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
017-9299-y;   https://link.springer.com/content/pdf/10.1007/
s10698-017-9299-y.pdf.

Murphy:2018:EIR

[454] Mark A. Murphy. Early industrial roots of green chemistry and the history of the BHC ibuprofen process invention and its quality connection. Foundations of Chemistry, 20(2):121–165, July 2018. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
017-9300-9;   https://link.springer.com/content/pdf/10.1007/
s10698-017-9300-9.pdf.

Scerri:2018:Ec

[455] E. R. Scerri. Editorial 60. Foundations of Chemistry, 20(3):167–168, October 2018. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
018-9322-y;   https://link.springer.com/content/pdf/10.1007/
s10698-018-9322-y.pdf.

Blumenthal:2018:TCC

[456] Geoffrey Blumenthal and James Ladyman. Theory comparison and choice in chemistry, 1766–1791. Foundations of Chemistry, 20(3):169–189, October 2018. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
017-9301-8;   https://link.springer.com/content/pdf/10.1007/
s10698-017-9301-8.pdf.

Grochala:2018:PHN

[457] Wojciech Grochala. On the position of helium and neon in the Periodic Table of Elements. Foundations of Chemistry, 20(3):191–207, October 2018. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
017-9302-7;   https://link.springer.com/content/pdf/10.1007/
s10698-017-9302-7.pdf.

Wray:2018:ANR

[458] K. Brad Wray. The atomic number revolution in chemistry: a Kuhnian analysis. Foundations of Chemistry, 20(3):209–217, October 2018. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

https://link.springer.com/article/10.1007/s10698-017-9303-6.

Seeman:2018:MSI

[459] Jeffrey I. Seeman. From “multiple simultaneous independent discoveries” to the theory of “multiple simultaneous independent errors”: a conduit in science. Foundations of Chemistry, 20(3):219–249, October 2018. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https:/

/link.springer.com/article/10.1007/s10698-018-9304-0.

Petrusevski:2018:TPH

[460] Vladimir M. Petrusevski and Julijana Cvetković. On the ‘true position’ of hydrogen in the Periodic Table. Foundations of Chemistry, 20 (3):251–260, October 2018. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-018-9306-y.

Fortin:2018:LUB

[461] Sebastian Fortin, Olimpia Lombardi, and Juan Camilo Martínez González. Let us build better boats: an answer to Jeffrey Seeman’s “Moving beyond insularity in the history, philosophy, and sociology of chemistry”. Foundations of Chemistry, 20(3):261–264, October 2018. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
018-9307-x;   https://link.springer.com/content/pdf/10.1007/
s10698-018-9307-x.pdf. See [449].

Malinen:2018:CCB

[462] KelleyAnne Malinen and Chérif F. Matta. “Climate change” and the “butterfly effect” in an eighteenth century monograph. Foundations of Chemistry, 20(3):265–268, October 2018. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-018-9305-z.

Scerri:2019:Ea

[463] Eric Scerri. Editorial 61. Foundations of Chemistry, 21(1):1–2, April 2019. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL https://link.springer.com/article/10.1007/
s10698-019-09335-6;   https://link.springer.com/content/pdf/
10.1007/s10698-019-09335-6.pdf.

Stewart:2019:MPS

[464] Philip J. Stewart. Mendeleev’s predictions: success and failure. Foundations of Chemistry, 21(1):3–9, April 2019. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-018-9312-0;   https://link.springer.com/
content/pdf/10.1007/s10698-018-9312-0.pdf.

Chamizo:2019:ACR

[465] José A. Chamizo. About continuity and rupture in the history of chemistry: the fourth chemical revolution (1945–1966). Foundations of Chemistry, 21(1):11–29, April 2019. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-018-9308-9.

Gonzalez:2019:WMS

[466] Juan Camilo Martínez González, Sebastian Fortin, and Olimpia Lombardi. Why molecular structure cannot be strictly reduced to quantum mechanics. Foundations of Chemistry, 21(1):31–45, April 2019. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https:/

/link.springer.com/article/10.1007/s10698-018-9310-2.

Johnson:2019:CVL

[467] David A. Johnson and Peter G. Nelson. Correction to: Valencies of the lanthanides. Foundations of Chemistry, 21(1):47–48, April 2019. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
018-9311-1;   https://link.springer.com/content/pdf/10.1007/
s10698-018-9311-1.pdf. See [444].

Fortin:2019:GEI

[468] Sebastian Fortin, Camilo Martinez Gonzalez, Alfio Zambon, and Waldmir Araujo Neto. Guest editorial: ISPC 2015 special issue. Foundations of Chemistry, 21(1):49–50, April 2019. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-019-09336-5;   https://link.springer.
com/content/pdf/10.1007/s10698-019-09336-5.pdf.

Banchetti-Robino:2019:FMB

[469] Marina Paola Banchetti-Robino. The function of microstructure in Boyle’s chemical philosophy: ‘chymical atoms’ and structural explanation. Foundations of Chemistry, 21(1):51–59, April 2019. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.

com/article/10.1007/s10698-018-09326-z.

Scerri:2019:FIC

[470] Eric R. Scerri. Five ideas in chemical education that must die. Foundations of Chemistry, 21(1):61–69, April 2019. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-018-09327-y.

Accorinti:2019:IMC

[471] Hernán Lucas Accorinti. Incompatible models in chemistry: the case of electronegativity. Foundations of Chemistry, 21(1):71–81, April 2019. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

https://link.springer.com/article/10.1007/s10698-018-09328-x.

Polzella:2019:RES

[472] María Silvia Polzella and Penélope Lodeyro. Re-evaluating semi-empirical computer simulations in quantum chemistry. Foundations of Chemistry, 21(1):83–95, April 2019. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-018-09329-w.

Gonzalez:2019:POI

[473] Juan Camilo Martínez González. The problem of optical isomerism and the interpretation of quantum mechanics. Foundations of Chemistry, 21 (1):97–107, April 2019. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-018-09330-3.

Martins:2019:EMM

[474] Roberto de Andrade Martins. Émile Meyerson and mass conservation in chemical reactions: a priori expectations versus experimental tests. Foundations of Chemistry, 21(1):109–124, April 2019. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.

com/article/10.1007/s10698-018-09331-2.

Arriaga:2019:NCP

[475] Jesus Alberto Jaimes Arriaga, Sebastian Fortin, and Olimpia Lombardi. A new chapter in the problem of the reduction of chemistry to physics: the Quantum Theory of Atoms in Molecules. Foundations of Chemistry, 21(1):125–136, April 2019. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-018-09332-1.

Scerri:2019:Eb

[476] Eric Scerri. Editorial 62. Foundations of Chemistry, 21(2):137–138, July 2019. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL https://link.springer.com/article/10.1007/
s10698-019-09339-2;   https://link.springer.com/content/pdf/
10.1007/s10698-019-09339-2.pdf.

Wray:2019:WMM

[477] K. Brad Wray. What to make of Mendeleev’s predictions? Foundations of Chemistry, 21(2):139–143, July 2019. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-018-9313-z.

Stewart:2019:GIR

[478] Philip J. Stewart. Getting it right is not equivalent to getting it wrong. Foundations of Chemistry, 21(2):145–146, July 2019. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

https://link.springer.com/article/10.1007/s10698-018-9316-9;
 https://link.springer.com/content/pdf/10.1007/s10698-018-9316-
9.pdf.

Blumenthal:2019:PVC

[479] Geoffrey Blumenthal. Priestley’s views on the composition of water and related airs. Foundations of Chemistry, 21(2):147–178, July 2019. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
018-9314-y;   https://link.springer.com/content/pdf/10.1007/
s10698-018-9314-y.pdf.

Nelson:2019:WCM

[480] Peter G. Nelson. What is chemistry that I may teach it? Foundations of Chemistry, 21(2):179–191, July 2019. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-018-9315-x;   https://link.springer.com/
content/pdf/10.1007/s10698-018-9315-x.pdf.

Liu:2019:RCE

[481] Runzhan Liu, Guoyong Mao, and Ning Zhang. Research of chemical elements and chemical bonds from the view of complex network. Foundations of Chemistry, 21(2):193–206, July 2019. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-018-9318-7.

Suarez:2019:NTN

[482] Rosana del P. Suárez. The network theory: a new language for speaking about chemical elements relations through stoichiometric binary compounds. Foundations of Chemistry, 21(2):207–220, July 2019. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https:/

/link.springer.com/article/10.1007/s10698-018-9319-6.

Quilez:2019:HEA

[483] Juan Quílez. A historical/epistemological account of the foundation of the key ideas supporting chemical equilibrium theory. Foundations of Chemistry, 21(2):221–252, July 2019. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-018-9320-0.

Scerri:2019:E

[484] Eric Scerri. Editorial 63. Foundations of Chemistry, 21(3):253–254, October 2019. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL https://link.springer.com/article/10.1007/
s10698-019-09341-8;   https://link.springer.com/content/pdf/
10.1007/s10698-019-09341-8.pdf.

Imyanitov:2019:DPT

[485] Naum S. Imyanitov. Does the period table appear doubled? Two variants of division of elements into two subsets. Internal and secondary periodicity. Foundations of Chemistry, 21(3):255–284, October 2019. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https:/

/link.springer.com/article/10.1007/s10698-018-9321-z.

Gilead:2019:FLP

[486] Amihud Gilead. Further light on the philosophical significance of Mackay’s theoretical discovery of crystalline pure possibilities. Foundations of Chemistry, 21(3):285–296, October 2019. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-018-9323-x.

Pisano:2019:TFP

[487] Raffaele Pisano, Abdelkader Anakkar, Emilio Marco Pellegrino, and Maxime Nagels. Thermodynamic foundations of physical chemistry: reversible processes and thermal equilibrium into the history. Foundations of Chemistry, 21(3):297–323, October 2019. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-018-09324-1.

Nelson:2019:SMI

[488] Peter G. Nelson. Statistical mechanical interpretation of temperature. Foundations of Chemistry, 21(3):325–331, October 2019. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://

link.springer.com/article/10.1007/s10698-019-09337-4;   https:
//link.springer.com/content/pdf/10.1007/s10698-019-09337-4.pdf.

Santana:2019:MMW

[489] Carlos Santana. Mineral misbehavior: why mineralogists don’t deal in natural kinds. Foundations of Chemistry, 21(3):333–343, October 2019. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

https://link.springer.com/article/10.1007/s10698-019-09338-3.

StGermain:2019:GVC

[490] Elijah St. Germain. Giovanni Villani: Chemistry: A systemic complexity science. Foundations of Chemistry, 21(3):345–348, October 2019. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https:/

/link.springer.com/article/10.1007/s10698-019-09340-9.

Scerri:2020:Ea

[491] Eric Scerri. Editorial 64. Foundations of Chemistry, 22(1):1–2, April 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL https://link.springer.com/article/10.1007/
s10698-020-09360-w;   https://link.springer.com/content/pdf/
10.1007/s10698-020-09360-w.pdf.

Stewart:2020:THT

[492] Philip J. Stewart. From telluric helix to telluric remix. Foundations of Chemistry, 22(1):3–14, April 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-019-09334-7;   https://link.springer.
com/content/pdf/10.1007/s10698-019-09334-7.pdf.

Seifert:2020:RID

[493] Vanessa A. Seifert. The role of idealisations in describing an isolated molecule. Foundations of Chemistry, 22(1):15–29, April 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL https://link.springer.com/article/10.1007/
s10698-019-09342-7;   https://link.springer.com/content/pdf/
10.1007/s10698-019-09342-7.pdf.

Friend:2020:BFL

[494] Michèle Friend. The beginnings of a formal language for conceptual analysis of processes in macro-chemistry. Foundations of Chemistry, 22(1):31–42, April 2020. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-019-09343-6.

Price:2020:MTC

[495] Justin Price. Model transfer and conceptual progress: tales from chemistry and biology. Foundations of Chemistry, 22(1):43–57, April 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https:/

/link.springer.com/article/10.1007/s10698-019-09344-5.

Buyse:2020:BSG

[496] Filip A. A. Buyse. Boyle, Spinoza and Glauber: on the philosophical redintegration of saltpeter —a reply to Antonio Clericuzio. Foundations of Chemistry, 22(1):59–76, April 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-019-09345-4.

Ochiai:2020:OSA

[497] Hirofumi Ochiai. Overcoming skepticism about molecular structure by developing the concept of affordance. Foundations of Chemistry, 22 (1):77–86, April 2020. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-019-09349-0.

Eaton:2020:WSW

[498] Gareth R. Eaton. Whom should we credit for the discovery of isotopes? Foundations of Chemistry, 22(1):87–98, April 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.

springer.com/article/10.1007/s10698-019-09347-2.

Dias:2020:TCB

[499] Jerry Ray Dias. Teaching of chemistry before and after the periodic table. Foundations of Chemistry, 22(1):99–106, April 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.

springer.com/article/10.1007/s10698-020-09354-8.

Matta:2020:TSE

[500] Chérif F. Matta, Olimpia Lombardi, and Jesús Jaimes Arriaga. Two-step emergence: the quantum theory of atoms in molecules as a bridge between quantum mechanics and molecular chemistry. Foundations of Chemistry, 22(1):107–129, April 2020. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-020-09352-w.

Levere:2020:PGI

[501] Trevor H. Levere. Pere Grapí: Inspiring air: a history of air-related science. Foundations of Chemistry, 22(1):131–133, April 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https:/

/link.springer.com/article/10.1007/s10698-019-09351-6.

Scerri:2020:Eb

[502] Eric Scerri. Editorial 65. Foundations of Chemistry, 22(2):135–136, July 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL https://link.springer.com/article/10.1007/
s10698-020-09374-4;   https://link.springer.com/content/pdf/
10.1007/s10698-020-09374-4.pdf.

Wray:2020:RDN

[503] K. Brad Wray and Line Edslev Andersen. Reporting the discovery of new chemical elements: working in different worlds, only 25 years apart. Foundations of Chemistry, 22(2):137–146, July 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.

com/article/10.1007/s10698-019-09348-1.

Lalvani:2020:CLC

[504] Haresh Lalvani. 4D-cubic lattice of chemical elements. Foundations of Chemistry, 22(2):147–194, July 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-019-09350-7. See correction [505].

Lalvani:2020:CCL

[505] Haresh Lalvani. Correction to: 4D-cubic lattice of chemical elements. Foundations of Chemistry, 22(2):195, July 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.

com/article/10.1007/s10698-020-09353-9;   https://link.springer.
com/content/pdf/10.1007/s10698-020-09353-9.pdf. See [504].

Campbell:2020:HMI

[506] Chris Campbell and Karoliina Pulkkinen. How Mendeleev issued his predictions: comment on Andrea Woody. Foundations of Chemistry, 22(2):197–215, July 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-020-09355-7;   https://link.springer.
com/content/pdf/10.1007/s10698-020-09355-7.pdf.

Vernon:2020:OMN

[507] René E. Vernon. Organising the metals and nonmetals. Foundations of Chemistry, 22(2):217–233, July 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-020-09356-6;   https://link.springer.
com/content/pdf/10.1007/s10698-020-09356-6.pdf.

Samanez:2020:BPS

[508] Julio Antonio Gutiérrez Samanez. Binódic periodic system: a mathematical approach. Foundations of Chemistry, 22(2):235–266, July 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

https://link.springer.com/article/10.1007/s10698-020-09359-3.

Hagino:2020:NPT

[509] K. Hagino and Y. Maeno. A nuclear periodic table. Foundations of Chemistry, 22(2):267–273, July 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-020-09365-5;   https://link.springer.
com/content/pdf/10.1007/s10698-020-09365-5.pdf.

Emsley:2020:EC

[510] John Emsley. Elements of chemistry. Foundations of Chemistry, 22 (2):275–277, July 2020. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-020-09366-4.

Ghibaudi:2020:SST

[511] Elena Ghibaudi, Luigi Cerruti, and Giovanni Villani. Structure, shape, topology: entangled concepts in molecular chemistry. Foundations of Chemistry, 22(2):279–307, July 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-019-09333-8.

Taber:2020:CCC

[512] Keith S. Taber. Conceptual confusion in the chemistry curriculum: exemplifying the problematic nature of representing chemical concepts as target knowledge. Foundations of Chemistry, 22(2):309–334, July 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL https://link.springer.com/article/10.1007/
s10698-019-09346-3;   https://link.springer.com/content/pdf/
10.1007/s10698-019-09346-3.pdf.

Tandon:2020:SAE

[513] Hiteshi Tandon, Tanmoy Chakraborty, and Vandana Suhag. A scale of atomic electronegativity in terms of atomic nucleophilicity index. Foundations of Chemistry, 22(2):335–346, July 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.

com/article/10.1007/s10698-020-09358-4.

Scerri:2020:Ec

[514] Eric Scerri. Editorial 66. Foundations of Chemistry, 22(3):347, October 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL https://link.springer.com/article/10.1007/
s10698-020-09386-0;   https://link.springer.com/content/pdf/
10.1007/s10698-020-09386-0.pdf.

Ghibaudi:2020:GE

[515] Elena Ghibaudi and Luigi Cerruti. Guest editorial. Foundations of Chemistry, 22(3):349–351, October 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-020-09385-1;   https://link.springer.
com/content/pdf/10.1007/s10698-020-09385-1.pdf.

Sukumar:2020:OST

[516] N. Sukumar. Ontological status of time in chemistry. Foundations of Chemistry, 22(3):353–361, October 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-020-09370-8.

Santos:2020:RCA

[517] Gil Santos, Gabriel Vallejos, and Davide Vecchi. A relationalconstructionist account of protein macrostructure and function. Foundations of Chemistry, 22(3):363–382, October 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.

springer.com/article/10.1007/s10698-020-09373-5.

Gini:2020:QSP

[518] Giuseppina Gini. The QSAR similarity principle in the deep learning era: Confirmation or revision? Foundations of Chemistry, 22(3):383–402, October 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
020-09380-6.

Ruthenberg:2020:CP

[519] Klaus Ruthenberg and Ave Mets. Chemistry is pluralistic. Foundations of Chemistry, 22(3):403–419, October 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-020-09378-0;   https://link.springer.
com/content/pdf/10.1007/s10698-020-09378-0.pdf.

Drago:2020:NDR

[520] Antonino Drago. A new definition of reduction between two scientific theories: no reduction of chemistry to quantum mechanics. Foundations of Chemistry, 22(3):421–445, October 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.

com/article/10.1007/s10698-020-09377-1.

Hijmans:2020:HIU

[521] Sarah Hijmans and Jean-Pierre Llored. How to investigate the underpinnings of sciences? The case of the element chlorine. Foundations of Chemistry, 22(3):447–456, October 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-020-09372-6.

Ochiai:2020:UMS

[522] Hirofumi Ochiai. Understanding molecular structure requires constructive realism. Foundations of Chemistry, 22(3):457–465, October 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https:/

/link.springer.com/article/10.1007/s10698-020-09362-8.

Zerecero:2020:MMS

[523] Gabriela García Zerecero. Molecular models and scientific realism. Foundations of Chemistry, 22(3):467–476, October 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.

springer.com/article/10.1007/s10698-020-09363-7.

Sereno:2020:PAP

[524] Sergio Gabriele Maria Sereno. Prediction, accommodation and the periodic table: a reappraisal. Foundations of Chemistry, 22(3):477–488, October 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
020-09371-7.

Headley:2020:PNK

[525] Clevis Headley. Philosophy, natural kinds, microstructuralism, and the (mis)use of chemical examples: intimacy versus integrity as orientations towards chemical practice. Foundations of Chemistry, 22(3):489–500, October 2020. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
020-09368-2.

Scerri:2021:Ea

[526] Eric Scerri. Editorial 67. Foundations of Chemistry, 23(1):1–2, April 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL https://link.springer.com/article/10.1007/
s10698-021-09397-5;   https://link.springer.com/content/pdf/
10.1007/s10698-021-09397-5.pdf.

Escribano-Cabeza:2021:CDTb

[527] Miguel Escribano-Cabeza. Chemistry and dynamics in the thought of G. W. Leibniz II. Foundations of Chemistry, 23(1):3–16, April 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https:/

/link.springer.com/article/10.1007/s10698-020-09375-3.

DiGiacomo:2021:ETC

[528] Francesco Di Giacomo. Early theoretical chemistry: Plato’s chemistry in Timaeus. Foundations of Chemistry, 23(1):17–30, April 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https:/

/link.springer.com/article/10.1007/s10698-020-09364-6.

Gilead:2021:CAM

[529] Amihud Gilead. A comment about the meaning and significance of life in the light of generalized crystallography. Foundations of Chemistry, 23(1):31–40, April 2021. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-020-09367-3.

Grochala:2021:WCA

[530] Wojciech Grochala. Watch the colors: or about qualitative thinking in chemistry. Foundations of Chemistry, 23(1):41–58, April 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL https://link.springer.com/article/10.1007/
s10698-020-09369-1;   https://link.springer.com/content/pdf/
10.1007/s10698-020-09369-1.pdf.

Healy:2021:OCR

[531] Eamonn F. Healy. Organic chemistry as representation. Foundations of Chemistry, 23(1):59–68, April 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-020-09379-z.

Muursepp:2021:CBS

[532] Peeter Müürsepp, Gulzhikhan Nurysheva, Aliya Ramazanova, and Zhamilya Amirkulova. Chemistry as the basic science. Foundations of Chemistry, 23(1):69–83, April 2021. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-020-09357-5.

Quilez:2021:ETK

[533] Juan Quílez. On the early thermodynamic and kinetic deductions of the equilibrium constant equation. Foundations of Chemistry, 23(1):85–103, April 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
020-09376-2.

Syropoulos:2021:VC

[534] Apostolos Syropoulos. On vague chemistry. Foundations of Chemistry, 23(1):105–113, April 2021. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-020-09383-3.

Zwart:2021:RPL

[535] Hub Zwart. Revolutionary poetry and liquid crystal chemistry: Herman Gorter, Ada Prins and the interface between literature and science. Foundations of Chemistry, 23(1):115–132, April 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.

com/article/10.1007/s10698-020-09381-5;   https://link.springer.
com/content/pdf/10.1007/s10698-020-09381-5.pdf.

Scerri:2021:BRC

[536] Eric R. Scerri. Book review: Charles S. McCaw: Orbitals with applications in atomic spectra, 2nd edition. Foundations of Chemistry, 23(1):133–134, April 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
020-09388-y.

Scerri:2021:Eb

[537] Eric Scerri. Editorial 68. Foundations of Chemistry, 23(2):135–136, July 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
021-09407-6.

Escribano-Cabeza:2021:CDTa

[538] Miguel Escribano-Cabeza. Chemistry and dynamics in the thought of G. W. Leibniz I. Foundations of Chemistry, 23(2):137–153, July 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

https://link.springer.com/article/10.1007/s10698-020-09382-4.

Vernon:2021:LCG

[539] René E. Vernon. The location and composition of Group 3 of the periodic table. Foundations of Chemistry, 23(2):155–197, July 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https:/

/link.springer.com/article/10.1007/s10698-020-09384-2. See cor-

rection [540].

Vernon:2021:CLC

[540] René E. Vernon. Correction to: The location and composition of Group 3 of the periodic table. Foundations of Chemistry, 23(2):199, July 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
020-09390-4. See [539].

Maeno:2021:TRT

[541] Yoshiteru Maeno, Kouichi Hagino, and Takehiko Ishiguro. Three related topics on the periodic tables of elements. Foundations of Chemistry, 23 (2):201–214, July 2021. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-020-09387-z.

Marks:2021:MR

[542] E. G. Marks and J. A. Marks. Mendeleyev revisited. Foundations of Chemistry, 23(2):215–223, July 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-021-09398-4. See comment [557].

Lightfield:2021:LAC

[543] Ceth Lightfield. Logics for algorithmic chemistries. Foundations of Chemistry, 23(2):225–237, July 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-020-09391-3.

Neuss:2021:PEE

[544] Geoffrey R. H. Neuss. A problem with explaining the electron configuration of scandium. Foundations of Chemistry, 23(2):239–245, July 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

https://link.springer.com/article/10.1007/s10698-020-09392-2.

Scerri:2021:RGN

[545] Eric Scerri. Response to Geoffrey Neuss on how to teach the 4s and 3d orbital conundrum. Foundations of Chemistry, 23(2):247–251, July 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

https://link.springer.com/article/10.1007/s10698-021-09395-7.

deBerg:2021:ADA

[546] Kevin C. de Berg. An analysis of the difficulties associated with determining that a reaction in chemical equilibrium is incomplete. Foundations of Chemistry, 23(2):253–275, July 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-020-09393-1.

Dunmur:2021:LCC

[547] David Dunmur. Liquid crystal chemistry and poetry. Foundations of Chemistry, 23(2):277–287, July 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-020-09394-0.

Chalmers:2021:RMP

[548] Alan Chalmers. Review of Marina Paola Banchetti-Robino, The chemical philosophy of Robert Boyle: mechanicism, chymical atoms and emergence. Pp. x + 196, Oxford: Oxford University Press, 2020. Foundations of Chemistry, 23(2):289–291, July 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-021-09399-3.

Scerri:2021:BRG

[549] Eric Scerri. Book review: Geoff Rayner-Canham: The periodic table: past present, and future. Foundations of Chemistry, 23(2):293–295, July 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

https://link.springer.com/article/10.1007/s10698-020-09389-x.

Scerri:2021:Ec

[550] Eric Scerri. Editorial 69. Foundations of Chemistry, 23(3):297–298, October 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
021-09413-8.

Marques:2021:IVG

[551] Carlos Alberto Marques and Adelio A. S. C. Machado. An integrated vision of the Green Chemistry evolution along 25 years. Foundations of Chemistry, 23(3):299–328, October 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-021-09396-6.

Silverberg:2021:WOS

[552] Lee J. Silverberg. What is an organic substance? Foundations of Chemistry, 23(3):329–336, October 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-021-09400-z.

Pisano:2021:CPE

[553] Raffaele Pisano, Emilio Marco Pellegrino, and Maxime Nagels. Conceptual polymorphism of entropy into the history: extensions of the second law of thermodynamics towards statistical physics and chemistry during nineteenth–twentieth centuries. Foundations of Chemistry, 23(3): 337–378, October 2021. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-021-09401-y.

Fortin:2021:PMS

[554] Sebastian Fortin and Olimpia Lombardi. Is the problem of molecular structure just the quantum measurement problem? Foundations of Chemistry, 23(3):379–395, October 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-021-09402-x.

Nelson:2021:STC

[555] Peter G. Nelson. A simple treatment of chemical equilibrium. Foundations of Chemistry, 23(3):397–405, October 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.

com/article/10.1007/s10698-021-09404-9.

Weber:2021:QMA

[556] Erik Weber, Merel Lefevere, and Kristian Gonzalez Barman. Quantum mechanical atom models, legitimate explanations and mechanisms. Foundations of Chemistry, 23(3):407–429, October 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.

springer.com/article/10.1007/s10698-021-09408-5.

Petrusevski:2021:BCM

[557] Vladimir M. Petrusevski. A brief comment on ‘Mendeleyev Revisited’ by Marks & Marks (foundations of chemistry, https://doi.org/10.1007/s10698-021-09398-4). Foundations of Chemistry, 23(3):431–432, October 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

https://link.springer.com/article/10.1007/s10698-021-09411-w.

See [542].

Taber:2021:MSP

[558] Keith S. Taber. Making sense of a pedagogic text. Foundations of Chemistry, 23(3):433–457, October 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-021-09405-8.

Aduriz-Bravo:2021:ICT

[559] Agustín Adúriz-Bravo. Improving chemistry teacher education with the philosophy of chemistry. Foundations of Chemistry, 23(3):459–463, October 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
021-09409-4.

Ramberg:2021:BRE

[560] Peter J. Ramberg. Book review: Eric Scerri and Elena Ghibaudi, eds: What is an element? A collection of essays by chemists, philosophers, historians, and educators. Foundations of Chemistry, 23(3):465–473, October 2021. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
021-09406-7.

Scerri:2022:EPI

[561] Eric Scerri. Editorial 70 (the platinum issue). Foundations of Chemistry, 24(1):1–2, April 2022. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-022-09429-8.

Nelson:2022:UE

[562] Peter G. Nelson. Understanding entropy. Foundations of Chemistry, 24(1):3–13, April 2022. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-021-09412-9.

Hisamatsu:2022:ONR

[563] Yoji Hisamatsu, Kazuhiro Egashira, and Yoshiteru Maeno. Ogawa’s nipponium and its re-assignment to rhenium. Foundations of Chemistry, 24(1):15–57, April 2022. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-021-09410-x.

Pucci:2022:DFT

[564] R. Pucci and G. G. N. Angilella. Density functional theory, chemical reactivity, and the Fukui functions. Foundations of Chemistry, 24(1):59–71, April 2022. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
022-09416-z.

Viswanathan:2022:OPS

[565] Balakrishnan Viswanathan and Mohamed Shajahan Gulam Razul. Orthogonality properties of states, configurations, and orbitals. Foundations of Chemistry, 24(1):73–86, April 2022. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-022-09417-y.

Raos:2022:STR

[566] Nenad Raos. Special theory of relativity in chemistry. Foundations of Chemistry, 24(1):87–95, April 2022. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-022-09420-3.

Blackie:2022:KBC

[567] Margaret A. L. Blackie. Knowledge building in chemistry education. Foundations of Chemistry, 24(1):97–111, April 2022. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.

com/article/10.1007/s10698-022-09419-w.

daSilva:2022:HPL

[568] Sérgio Luís da Silva. The habit of the pipe: a layperson’s view of the periodic table. Foundations of Chemistry, 24(1):113–120, April 2022. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https:/

/link.springer.com/article/10.1007/s10698-022-09422-1.

Goodwin:2022:CPS

[569] Alan Goodwin. Chemistry: progress since 1860 — reflections on chemistry and chemistry education triggered by reading Muspratt’s Chemistry. Foundations of Chemistry, 24(1):121–142, April 2022. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https:/

/link.springer.com/article/10.1007/s10698-022-09424-z.

Talpsepp:2022:IES

[570] Edit Talpsepp. Interview with Eric Scerri. Foundations of Chemistry, 24(1):143–153, April 2022. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-022-09418-x.

Atkins:2022:BB

[571] Peter Atkins. Brick by brick. Foundations of Chemistry, 24(1):155–157, April 2022. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
021-09415-6.

Scerri:2022:Ea

[572] Eric Scerri. Editorial 71. Foundations of Chemistry, 24(2):159–160, July 2022. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
022-09441-y.

Wray:2022:WHW

[573] K. Brad Wray. What happened when chemists came to classify elements by their atomic number? Foundations of Chemistry, 24(2):161–170, July 2022. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
022-09423-0.

Ferreira:2022:DCC

[574] Larissa Moreira Ferreira, Jean Pscheidt Weiss, and Marcelo Lambach. Disparities and conceptual connections regarding the concept of substance in general chemistry textbook glossaries. Foundations of Chemistry, 24 (2):171–187, July 2022. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-022-09425-y.

Imyanitov:2022:PTC

[575] Naum S. Imyanitov. Periodic tables for cations + 1, + 2, + 3 and anions — 1. Quantitative characteristics for manifestations of internal periodicity and kainosymmetry. Foundations of Chemistry, 24(2):189–219, July 2022. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

https://link.springer.com/article/10.1007/s10698-022-09421-2.

Paparazzo:2022:PC

[576] Ernesto Paparazzo. Plato on chemistry. Foundations of Chemistry, 24 (2):221–238, July 2022. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-022-09426-x.

Hartman:2022:ISS

[577] JudithAnn R. Hartman, Eric A. Nelson, and Paul A. Kirschner. Improving student success in chemistry through cognitive science. Foundations of Chemistry, 24(2):239–261, July 2022. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-022-09427-w.

Sharma:2022:QAS

[578] Smriti Sharma. Quantum algorithms for simulation of quantum chemistry problems by quantum computers: an appraisal. Foundations of Chemistry, 24(2):263–276, July 2022. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-022-09428-9.

Anooshe:2022:CEZ

[579] Nazila Farmani Anooshe and Aliyar Mousavi. The case of Zinjafr in the medical and mineralogical texts of medieval Persia: a puzzle created in the absence of the concept of chemical elements. Foundations of Chemistry, 24(2):277–284, July 2022. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-022-09436-9.

Scerri:2022:PT

[580] Eric R. Scerri. In praise of triads. Foundations of Chemistry, 24 (2):285–300, July 2022. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-022-09434-x.

Labinger:2022:BR

[581] Jay Labinger. Book review. Foundations of Chemistry, 24(2):301–310, July 2022. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
022-09438-7.

Scerri:2022:Eb

[582] Eric Scerri. Editorial 72. Foundations of Chemistry, 24(3):311–313, October 2022. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
022-09445-8.

Longo:2022:HFL

[583] Savino Longo. Hints for a formal language inspired by Lewis structures. Foundations of Chemistry, 24(3):315–330, October 2022. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https:/

/link.springer.com/article/10.1007/s10698-022-09439-6.

Imyanitov:2022:NPT

[584] Naum S. Imyanitov. Non-periodic table of periodicities and periodic table with additional periodicities: tetrad periodicity. Foundations of Chemistry, 24(3):331–358, October 2022. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-022-09437-8.

Sharma:2022:PSA

[585] Praveen Kumar Sharma. Prospective sustainable agriculture principles inspired by green chemistry. Foundations of Chemistry, 24(3):359–362, October 2022. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
022-09442-x.

Seeman:2022:CEB

[586] Jeffrey I. Seeman. A case for the engagement between the sciences and the humanities. Jay A. Labinger’s: Connecting Literature and Science. New York: Routledge, 2022. Foundations of Chemistry, 24(3):363–373, October 2022. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
022-09440-z.

Zambon:2022:CRC

[587] Alfio Zambon. Chemical reactivity: cause-effect or interaction? Foundations of Chemistry, 24(3):375–387, October 2022. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.

springer.com/article/10.1007/s10698-022-09430-1.

Scerri:2022:HCN

[588] Eric R. Scerri. Hasok Chang on the nature of acids. Foundations of Chemistry, 24(3):389–404, October 2022. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-022-09432-z.

Alassia:2022:POA

[589] Fiorela Alassia. A process ontology approach in biochemistry: the case of GPCRs and biosignaling. Foundations of Chemistry, 24(3):405–422, October 2022. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-

022-09443-w. See correction [603].

Banchetti-Robino:2022:RBR

[590] Marina Paola Banchetti-Robino. Robert Boyle and the relational and dispositional nature of chemical properties. Foundations of Chemistry, 24 (3):423–431, October 2022. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-022-09435-w.

Scerri:2023:Ea

[591] Eric Scerri. Editorial 73. Foundations of Chemistry, 25(1):1–4, April 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

https://link.springer.com/article/10.1007/s10698-023-09472-z.

Viswanathan:2023:EPR

[592] Balakrishnan Viswanathan and M. Shajahan Gulam Razul. Electronegativity provides the relationship between formal charge, oxidation state, and actual charge. Foundations of Chemistry, 25(1):5–28, April 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

https://link.springer.com/article/10.1007/s10698-022-09447-6.

Miskowiec:2023:NGNa

[593] Paweł Miśkowiec. Name game: the naming history of the chemical elements — part 1 — from antiquity till the end of 18th century. Foundations of Chemistry, 25(1):29–51, April 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-022-09448-5. See correction [594].

Miskowiec:2023:CNG

[594] Paweł Miśkowiec. Correction to: Name game: the naming history of the chemical elements — part 1 — from antiquity till the end of 18th century. Foundations of Chemistry, 25(1):53–55, April 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://

link.springer.com/article/10.1007/s10698-022-09462-7. See [593].

Wray:2023:TRD

[595] K. Brad Wray. Theodore Richards and the discovery of isotopes. Foundations of Chemistry, 25(1):57–66, April 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.

com/article/10.1007/s10698-022-09449-4. See response [625].

Cat:2023:ANI

[596] Jordi Cat and Nicholas W. Best. Atomic number and isotopy before nuclear structure: multiple standards and evolving collaboration of chemistry and physics. Foundations of Chemistry, 25(1):67–99, April 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

https://link.springer.com/article/10.1007/s10698-022-09450-x.

Scerri:2023:IOL

[597] Eric R. Scerri. Interview with Olimpia Lombardi. Foundations of Chemistry, 25(1):101–117, April 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-022-09453-8.

Anderson:2023:ESC

[598] G. M. Anderson. Entropy and sign conventions. Foundations of Chemistry, 25(1):119–125, April 2023. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-022-09463-6.

JaimesArriaga:2023:NQC

[599] Jesus Alberto Jaimes Arriaga. On the nature of quantum-chemical entities: the case of electron density. Foundations of Chemistry, 25 (1):127–139, April 2023. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-022-09431-0.

Ochiai:2023:PGD

[600] Hirofumi Ochiai. Philosophical grounds for designing invisible molecules. Foundations of Chemistry, 25(1):141–149, April 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.

springer.com/article/10.1007/s10698-022-09433-y.

Pergola:2023:MLR

[601] Martín Pérgola and Lydia Galagovsky. Models, languages and representations: philosophical reflections driven from a research on teaching and learning about cellular respiration. Foundations of Chemistry, 25 (1):151–166, April 2023. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-022-09444-9.

Headley:2023:NKC

[602] Clevis Headley. Natural kinds, chemical practice, and interpretive communities. Foundations of Chemistry, 25(1):167–187, April 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https:/

/link.springer.com/article/10.1007/s10698-022-09459-2.

Alassia:2023:CPO

[603] Fiorela Alassia. Correction to: A process ontology approach in biochemistry: the case of GPCRs and biosignaling. Foundations of Chemistry, 25(1):189–206, April 2023. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-022-09461-8. See [589].

Mao:2023:PUB

[604] Guoyong Mao, Runzhan Liu, and Ning Zhang. Predicting unknown binary compounds from the view of complex network. Foundations of Chemistry, 25(2):207–214, July 2023. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-022-09457-4.

Miskowiec:2023:NGNb

[605] Paweł Miśkowiec. Name game: the naming history of the chemical elements: part 2 — turbulent nineteenth century. Foundations of Chemistry, 25(2):215–234, July 2023. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-022-09451-w.

Miskowiec:2023:NGNc

[606] Paweł Miśkowiec. Name game: the naming history of the chemical elements: part 3 — rivalry of scientists in the twentieth and twenty-first centuries. Foundations of Chemistry, 25(2):235–251, July 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https:/

/link.springer.com/article/10.1007/s10698-022-09452-9.

Scerri:2023:CWC

[607] Eric R. Scerri. A commentary on Weisberg’s critique of the ‘structural conception’ of chemical bonding. Foundations of Chemistry, 25(2):253–264, July 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
022-09454-7.

DiGiacomo:2023:HSF

[608] Francesco Di Giacomo. On how some fundamental chemical concepts are correlated by arithmetic, geometric and harmonic means. Foundations of Chemistry, 25(2):265–268, July 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-023-09466-x.

Garditz:2023:PUP

[609] Klaus Ferdinand Gärditz. The poetry of the universe, the periodic table, and the scientific progress: a review of new studies on the periodic table of the elements. Foundations of Chemistry, 25(2):269–283, July 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

https://link.springer.com/article/10.1007/s10698-023-09468-9.

Eaton:2023:IBB

[610] Gareth R. Eaton. Interpreting the bonding of B_{2}H_{6} and the nature of the 3-center-2-electron bond: decisive test of theory of valency. Foundations of Chemistry, 25(2):285–298, July 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-023-09469-8.

Tantillo:2023:UTA

[611] Dean J. Tantillo and Jeffrey I. Seeman. On a unified theory of acids and bases: Hasok Chang, Eric R. Scerri, modern theoretical chemistry, and the philosophy of chemistry. Foundations of Chemistry, 25(2):299–320, July 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
022-09456-5.

Seeman:2023:RSR

[612] Jeffrey I. Seeman. Revolutions in science, revolutions in chemistry. Foundations of Chemistry, 25(2):321–335, July 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.

com/article/10.1007/s10698-023-09467-w.

Hanlon:2023:BRP

[613] Robert T. Hanlon. Book review of Paul Sen’s, “Einstein’s Fridge. How the difference between hot and cold explains the universe” ISBN: 978-1-5011-8130-6. Foundations of Chemistry, 25(2):337–338, July 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https:/

/link.springer.com/article/10.1007/s10698-022-09446-7. See cor-

rection [614].

Hanlon:2023:CPS

[614] Robert T. Hanlon. Correction: Paul Sen’s, “Einstein’s Fridge. How the difference between hot and cold explains the universe” ISBN: 978-1-5011-8130-6. Foundations of Chemistry, 25(2):339, July 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://

link.springer.com/article/10.1007/s10698-022-09465-4. See [613].

Scerri:2023:Eb

[615] Eric R. Scerri. Editorial 75. Foundations of Chemistry, 25(3):341–342, October 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
023-09489-4.

Friend:2023:ENM

[616] Michèle Indira Friend. Editor’s note by Michele Friend. Foundations of Chemistry, 25(3):343–344, October 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-023-09487-6.

Vancik:2023:CS

[617] Hrvoj Vancik. From complexity to systems. Foundations of Chemistry, 25 (3):345–358, October 2023. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-022-09455-6.

Ochiai:2023:WDE

[618] Hirofumi Ochiai. Why do prima facie intuitive theories work in organic chemistry? Foundations of Chemistry, 25(3):359–367, October 2023. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

https://link.springer.com/article/10.1007/s10698-023-09470-1.

Suarez:2023:RCP

[619] Mauricio Suárez and Pedro J. Sánchez Gómez. Reactivity in chemistry: the propensity view. Foundations of Chemistry, 25(3):369–380, October 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
023-09477-8.

Gomez:2023:SRS

[620] Pedro J. Sánchez Gómez. Scientific representation and science identity: the case of chemistry. Foundations of Chemistry, 25(3):381–391, October 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
023-09481-y.

Bano:2023:DPE

[621] Safia Bano. A defense of placeholder essentialism. Foundations of Chemistry, 25(3):393–404, October 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-023-09478-7.

Miller:2023:CRQ

[622] Ryan Miller. Chemical reduction and quantum interpretation: a case for Thomistic emergence. Foundations of Chemistry, 25(3):405–417, October 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
023-09479-6.

Sari:2023:MCT

[623] Rr. Lis Permana Sari, Heru Pratomo, Isti Yunita, Sukisman Purtadi, Mahesh Narayan, and Kristian Handoyo Sugiyarto. Misconception in chemistry textbooks: a case study on the concept of quantum number, electronic configuration and review for teaching material. Foundations of Chemistry, 25(3):419–437, October 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-023-09475-w.

Xu:2023:CEF

[624] Yachun Xu, Yichen Tong, and Jiangyang Yuan. Common empirical foundations, different theoretical choices: The Berthollet–Proust controversy and Dalton’s resolution. Foundations of Chemistry, 25(3):439–455, October 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
023-09471-0.

Eaton:2023:RCD

[625] Gareth R. Eaton. Response to the critique by Dr. K. Brad Wray, published in Foundations of Chemistry October 6, 2022. Foundations of Chemistry, 25(3):457–461, October 2023. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-023-09473-y. See [595].

Krasnodebski:2023:UBH

[626] Marcin Krasnodebski. An unlikely bifurcation: history of sustainable (but not green) chemistry. Foundations of Chemistry, 25(3):463–484, October 2023. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
023-09474-x.

Scerri:2024:Ea

[627] Eric R. Scerri. Editorial 76. Foundations of Chemistry, 26(1):1–2, April 2024. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
024-09505-1.

Wentrup:2024:JRG

[628] Curt Wentrup. Johann Rudolph Glauber: the royals’ alchemist and his secret recipes. Foundations of Chemistry, 26(1):3–13, April 2024. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https:/

/link.springer.com/article/10.1007/s10698-023-09493-8.

Vernon:2024:HHP

[629] René Vernon. Hydrogen over helium: a philosophical position. Foundations of Chemistry, 26(1):15–36, April 2024. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.

com/article/10.1007/s10698-023-09496-5.

Rocci:2024:CBD

[630] Alessio Rocci. Celebrating the birth of De Donder’s chemical affinity (1922–2022): from the uncompensated heat to his Ave Maria. Foundations of Chemistry, 26(1):37–73, April 2024. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.com/

article/10.1007/s10698-023-09488-5.

Wray:2024:CAC

[631] K. Brad Wray. Co-authorship in chemistry at the turn of the twentieth century: the case of Theodore W. Richards. Foundations of Chemistry, 26(1):75–88, April 2024. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-023-09491-w.

Viswanathan:2024:RSA

[632] Balakrishnan Viswanathan and M. Shajahan Gulam Razul. Relating screening to atomic properties and electronegativity in the Slater atom. Foundations of Chemistry, 26(1):89–113, April 2024. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.

springer.com/article/10.1007/s10698-023-09492-9.

Carpentieri:2024:IUV

[633] Maria Antonietta Carpentieri and Valentina Domenici. Introducing UVvisible spectroscopy at high school level following the historical evolution of spectroscopic instruments: a proposal for chemistry teachers. Foundations of Chemistry, 26(1):115–139, April 2024. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.

com/article/10.1007/s10698-024-09501-5.

Estrada:2024:WMD

[634] Ernesto Estrada. What is a mathematician doing … in a chemistry class? Foundations of Chemistry, 26(1):141–166, April 2024. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https:/

/link.springer.com/article/10.1007/s10698-023-09497-4.

Lang:2024:BOB

[635] Peter F. Lang. Bond order and bond energies. Foundations of Chemistry, 26(1):167–177, April 2024. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-023-09486-7.

Hanlon:2024:DPM

[636] Robert T. Hanlon. Deciphering the physical meaning of Gibbs’s maximum work equation. Foundations of Chemistry, 26(1):179–189, April 2024. CO-DEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

https://link.springer.com/article/10.1007/s10698-024-09503-3.

Scerri:2024:Eb

[637] Eric Scerri. Editorial 77. Foundations of Chemistry, 26(2):191–192, August 2024. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
024-09520-2.

Ruthenberg:2024:CWB

[638] Klaus Ruthenberg. Centenary workshop on the bifurcation of acidity — protonism vs. electronism. Foundations of Chemistry, 26(2):193–196, August 2024. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
023-09490-x.

Flechsig:2024:UNC

[639] Gerd-Uwe Flechsig. Usanovich and Nernst colliding: inconsistencies in the all-in-one acid–base concept? Foundations of Chemistry, 26(2):197–202, August 2024. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
023-09482-x.

Gerontas:2024:STC

[640] Apostolos K. Gerontas. “Sharp of taste”: the concept of acidity in the Greek system of natural explanation. Foundations of Chemistry, 26(2): 203–211, August 2024. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-023-09483-w.

Ruthenberg:2024:B

[641] Klaus Ruthenberg. Bifurcations. Foundations of Chemistry, 26(2): 213–224, August 2024. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-023-09484-9.

Thyssen:2024:ANK

[642] Pieter Thyssen. Are acids natural kinds? Foundations of Chemistry, 26 (2):225–253, August 2024. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-023-09485-8.

deSouza:2024:IIR

[643] Karina Aparecida de Freitas Dias de Souza and Paulo Alves Porto. Interaction, interpretation and representation: the construction and dissemination of chemical knowledge from a Peircean semiotics perspective. Foundations of Chemistry, 26(2):255–273, August 2024. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.springer.

com/article/10.1007/s10698-024-09506-0.

Posa:2024:CDE

[644] Mihalj Posa. Connecting De Donder’s equation with the differential changes of thermodynamic potentials: understanding thermodynamic potentials. Foundations of Chemistry, 26(2):275–290, August 2024. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https:/

/link.springer.com/article/10.1007/s10698-024-09507-z.

Vivas-Reyes:2024:CPK

[645] Ricardo Vivas-Reyes. Clashing perspectives: Kantian epistemology and quantum chemistry theory. Foundations of Chemistry, 26(2):291–300, August 2024. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10698-
024-09508-y.

Sharma:2024:RSP

[646] Kamna Sharma, Deepak Kumar Das, and Saibal Ray. Research status of the periodic table: a bibliometric analysis. Foundations of Chemistry, 26 (2):301–314, August 2024. CODEN FOCHFL. ISSN 1386-4238 (print),

1572-8463 (electronic). URL https://link.springer.com/article/10.
1007/s10698-024-09509-x.

RodriguezPena:2024:PSE

[647] Mario Rodríguez Peña and José Ángel García Guerra. The periodic spiral of elements. Foundations of Chemistry, 26(2):315–321, August 2024. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL

https://link.springer.com/article/10.1007/s10698-024-09510-4.

Shakhashiri:1983:CDH

[648] Bassam Z. Shakhashiri. Chemical demonstrations: a handbook for teachers of chemistry. University of Wisconsin Press, Madison, WI, USA, 1983–2011. ISBN 0-299-08890-1 (vol. 1), 0-299-10130-4 (vol. 2), 0-299-11950-5 (vol. 3), 0-299-12860-1 (vol. 4), 0-299-22650-6 (vol. 5). xxiv + 343 (vol. 1: 1983), xxxvi + 311 (vol. 2: 1983), xx + 401 (vol. 3: 1989) xxvii + 290 (vol. 4: 1992), xxxiv + 323 (vol. 5: 2011) pp. LCCN QD43 .S5 1983. US$25.00. URL http://www.

loc.gov/catdir/enhancements/fy0608/81070016-d.html;   http:/
/www.loc.gov/catdir/enhancements/fy0608/81070016-t.html;   http:
//www.loc.gov/catdir/enhancements/fy0609/81070016-b.html.

Grosberg:1997:GMH

[649] A. Yu. Grosberg and A. R. Khokhlov. Giant molecules: here, and there, and everywhere. Academic Press, New York, USA, 1997. ISBN 0-12-304130-9. xv + 244 pp. LCCN QD381 .G755 1997. URL http://

www.loc.gov/catdir/description/els032/96027593.html;   http:
//www.loc.gov/catdir/toc/els032/96027593.html.

Pullman:1998:AHH

[650] Bernard Pullman. The atom in the history of human thought. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1998. ISBN 0-19-511447-7 (hardcover). x + 403 pp. LCCN QC171.2 .P8513 1998.

URL http://www.loc.gov/catdir/enhancements/fy0635/97036040-
d.html;   http://www.loc.gov/catdir/enhancements/fy0723/97036040-

b.html. Translation from French by Axel Reisinger.

Albert:2000:GLN

[651] Mark Albert. Galen’s lectures: a novel about chemistry. Xlibris, ????, 2000. ISBN 0-7388-4196-X. 493 pp. LCCN PS3551.L259 G35 2000.

Djerassi:2000:IMS

[652] Carl Djerassi. An immaculate misconception: sex in an age of mechanical reproduction. Imperial College Press, London, UK, 2000. ISBN 1-86094-248-2 (paperback). 1-84816-010-0 (e-book). xviii + 134 pp. LCCN PS3554.J47 I46 2000.

Laszlo:2000:MC

[653] Pierre Laszlo. Miroir de la chimie. Science ouverte. Seuil, Paris, France, 2000. ISBN 2-02-035273-7. 328 pp. LCCN QD39 .L37 2000.

vanBrakel:2000:PCB

[654] J. (Jaap) van Brakel. Philosophy of chemistry: between the manifest and the scientific image, volume 15 of Louvain philosophical studies. Leuven University Press, Leuven, Belgium, 2000. ISBN 90-5867-063-5. xiv + 246 pp. LCCN QD6 .B75 2000.

Christie:2001:OLP

[655] Maureen Christie. The Ozone Layer: A Philosophy of Science Perspective. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-65072-0 (hardcover), 0-521-65908-6 (paperback). xii + 215 pp. LCCN QC881.2.O9

C48 2001. URL http://www.loc.gov/catdir/description/cam021/
00036295.html;   http://www.loc.gov/catdir/samples/cam031/00036295.
html;   http://www.loc.gov/catdir/toc/cam021/00036295.html.

Gillespie:2001:CBM

[656] Ronald J. (Ronald James) Gillespie and Paul L. A. Popelier. Chemical bonding and molecular geometry: from Lewis to electron densities.

Topics in inorganic chemistry. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-510495-1, 0-19-510496-X (paperback). xiii + 268 pp. LCCN QD461 .G537 2001. URL http://www.

loc.gov/catdir/enhancements/fy0610/00025404-d.html;   http:/
/www.loc.gov/catdir/enhancements/fy0610/00025404-t.html.

Basolo:2002:CIC

[657] Fred Basolo. From Coello to inorganic chemistry: a lifetime of reactions. Profiles in inorganic chemistry. Kluwer Academic/Plenum Publishers, New York, NY, USA, 2002. ISBN 0-306-46774-7. xxii + 245

pp. LCCN QD22.B258 A3 2002. URL http://www.loc.gov/catdir/
enhancements/fy0813/2002022220-t.html;   http://www.loc.gov/
catdir/enhancements/fy0820/2002022220-d.html.

Cobb:2002:MMM

[658] Cathy Cobb. Magick, mayhem, and mavericks: the spirited history of physical chemistry. Prometheus Books, Amherst, NY, USA, 2002. ISBN 1-57392-976-X. 420 pp. LCCN QD452 .C63 2002.

Habashi:2002:AAB

[659] Fathi Habashi. From alchemy to atomic bombs: history of chemistry, metallurgy, and civilization. Métallurgie Extractive Québec, Sainte Foy, QC, Canada, 2002. ISBN 2-922686-00-0. x + 357 pp. LCCN QD11 .H23 2002.

Hentschel:2002:MST

[660] Klaus Hentschel. Mapping the spectrum: techniques of visual representation in research and teaching. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-850953-7 (hardcover). xii + 362

pp. LCCN QD95 .H47 2002. URL http://www.e-streams.com/es0511/
es0511\_2165.html;   http://www.loc.gov/catdir/enhancements/
fy0613/2002072578-d.html;   http://www.loc.gov/catdir/toc/fy032/
2002072578.html.

Seligardi:2002:LIC

[661] Raffaella Seligardi. Lavoisier in Italia: la comunità scientifica italiana e la rivoluzione chimica, volume 46 of Biblioteca di storia della scienza. L. S. Olschki, Firenze, Italia, 2002. ISBN 88-222-5150-4. ISSN 0394-5065. 410 pp. LCCN QD18.I8 S45 2002.

Pohanish:2003:WGC

[662] Richard P. Pohanish and Stanley A. Greene. Wiley guide to chemical incompatibilities. John Wiley, New York, NY, USA, second edition, 2003. ISBN 0-471-23859-7. xvi + 1278 pp. LCCN T55.3.H3 P647 2003.

URL http://www.loc.gov/catdir/bios/wiley044/2002032959.html;
 http://www.loc.gov/catdir/description/wiley037/2002032959.html;
 http://www.loc.gov/catdir/toc/wiley031/2002032959.html.

Gordin:2004:WOT

[663] Michael D. Gordin. A well-ordered thing: Dmitrii Mendeleev and the shadow of the periodic table. Basic Books, New York, NY, USA, 2004. ISBN 0-465-02775-X. xx + 364 pp. LCCN QD22.M43

G67 2004. URL http://www.loc.gov/catdir/enhancements/fy0832/
2003025533-b.html;   http://www.loc.gov/catdir/enhancements/
fy0832/2003025533-d.html;   http://www.loc.gov/catdir/toc/ecip0411/
2003025533.html.

Scerri:2007:PTS

[664] Eric R. Scerri. The Periodic Table: Its Story and Its Significance. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-530573-6. xxii + 346 pp. LCCN QD467

.S345 2007. URL http://www.loc.gov/catdir/enhancements/fy0635/
2005037784-d.html;   http://www.loc.gov/catdir/enhancements/
fy0722/2005037784-b.html;   http://www.loc.gov/catdir/toc/ecip065/
2005037784.html.

Bensaude-Vincent:2008:CIS

[665] Bernadette Bensaude-Vincent and Jonathan Simon. Chemistry: the impure science. Imperial College Press, London, UK, 2008. ISBN 1-84816-225-1 (hardcover). xii + 268 pp. LCCN QD31.3 .B455 2008.

Brock:2008:WCC

[666] W. H. (William Hodson) Brock. William Crookes (1832–1919) and the commercialization of science. Science, technology, and culture. Ashgate, Aldershot, England, 2008. ISBN 0-7546-6322-1. ISSN 1700-1945. xxvii + 556 pp. LCCN Q143.C7 B76 2008.

Chalmers:2009:SAP

[667] A. F. (Alan Francis) Chalmers. The Scientist’s Atom and the Philosopher’s Stone: How Science Succeeded and Philosophy Failed to Gain Knowledge of Atoms, volume 279 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 90-481-2361-5, 90-481-2362-3. ISSN 0068-0346. xii + 287 pp. LCCN QD461 .C375 2009. URL https://link.springer.

com/book/10.1007/978-90-481-2362-9.

He:2009:TDX

[668] Bob B. He. Two-dimensional X-ray diffraction. John Wiley, New York, NY, USA, 2009. ISBN 0-470-22722-2. xiv + 426 pp. LCCN QC482.D5

H4 2009. URL http://www.loc.gov/catdir/enhancements/fy1012/
2009005619-b.html;   http://www.loc.gov/catdir/enhancements/
fy1012/2009005619-d.html;   http://www.loc.gov/catdir/enhancements/
fy1012/2009005619-t.html.

Klabunde:2009:NMC

[669] Kenneth J. Klabunde and Ryan Richards. Nanoscale materials in chemistry. John Wiley, New York, NY, USA, second edition, 2009. ISBN 0-470-22270-0 (hardcover). xiii + 777 + 12 pp. LCCN TA418.9.N35 N345 2009.

Olah:2009:BOG

[670] George A. (George Andrew) Olah, Alain Goeppert, and G. K. Surya Prakash. Beyond oil and gas: the methanol economy. Wiley-VCH, Weinheim, German, second updated and enlarged edition, 2009. ISBN 3-527-32422-4. xvi + 334 pp. LCCN HD9502.5.M472

O43 2009. URL http://www.loc.gov/catdir/enhancements/fy1106/
2010485132-b.html;   http://www.loc.gov/catdir/enhancements/
fy1106/2010485132-d.html;   http://www.loc.gov/catdir/enhancements/
fy1106/2010485132-t.html.

Pohanish:2009:WGC

[671] Richard P. Pohanish and Stanley A. Greene. Wiley guide to chemical incompatibilities. John Wiley, New York, NY, USA, third edition, 2009. ISBN 0-470-38763-7. xviii + 1110 pp. LCCN T55.3.H3

P647 2009. URL http://www.loc.gov/catdir/enhancements/fy1011/
2009009766-b.html;   http://www.loc.gov/catdir/enhancements/
fy1011/2009009766-d.html;   http://www.loc.gov/catdir/enhancements/
fy1011/2009009766-t.html.

Roesky:2009:EGS

[672] Herbert W. Roesky, Dietmar K. Kennepohl, and Jean-Marie Lehn, editors. Experiments in green and sustainable chemistry. Wiley-VCH, Weinheim, Germany, 2009. ISBN 3-527-32546-8 (hardcover). xxiv + 283 pp. LCCN TP155.2.E58 E96 2009.

Fisher:2010:MAA

[673] David E. Fisher. Much ado about (practically) nothing: a history of the noble gases. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-539396-1 (hardcover). x + 264 pp. LCCN QD162 .F57 2010.

Kean:2010:DSO

[674] Sam Kean. The disappearing spoon: and other true tales of madness, love, and the history of the world from the periodic table of the elements. Little, Brown and Company, Boston, Toronto, London, 2010. ISBN 0-316-05164-0. vi + 391 pp. LCCN QD466 .K37 2010.

Grosberg:2011:GMH

[675] A. Yu. Grosberg and A. R. Khokhlov. Giant molecules: here, there, and everywhere. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, second edition, 2011. ISBN 981-283-922-4 (hardcover). xxiv + 322 pp. LCCN QD381 .G755 2011. Foreword by Pierre-Gilles de Gennes.

WeidlerKubanek:2011:NLT

[676] Anne-Marie Weidler Kubanek. Nothing less than an adventure: Ellen Gleditsch and her life in science. Createspace Online, ????, 2011. ISBN 1-4528-4213-2. 185 pp. LCCN ????

Llored:2013:PCP

[677] Jean-Pierre Llored, editor. The philosophy of chemistry: practices, methodologies, and concepts. Cambridge Scholars Publishing, Newcastle upon Tyne, UK, 2013. ISBN 1-4438-4605-8 (hardcover). xix + 764 pp. LCCN QD6 .P49 2013.

Scerri:2013:SEMa

[678] Eric R. Scerri. 30-second elements: the 50 most significant elements, each explained in half a minute. Metro Books, New York, NY, USA, 2013. ISBN 1-4351-4521-6. 160 pp. LCCN QD466 .S34 2013.

Scerri:2013:SEMb

[679] Eric R. Scerri. 30-second elements: the 50 most significant elements, each explained in half a minute. Icon, London, UK, 2013. ISBN 1-84831-594-5. 160 pp. LCCN ???? With Hugh Aldersey-Williams.

Scerri:2013:SEMc

[680] Eric R. Scerri and Danian (Ying Yu) Li. 30-second elements. Miao gao dong hua xue yuan su: zhou qi biao mei ni xiang de na me nan, volume 1. Qi lin wen hua chu ban, Tai bei shi, Taiwan, 2013. ISBN 986-5893-54-1. ???? pp. LCCN QD466.

Scerri:2013:TE

[681] Eric R. Scerri. A tale of 7 elements. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-539131-4 (hardcover), 0-19-987503-0 (e-book). xxxiii + 270 pp. LCCN QD467 .S3475 2013. URL

http://catalogue.bnf.fr/ark:/12148/cb436625450.

Barton:1982:RBW

[682] Sir Derek Barton, FRS, editor. R. B. Woodward remembered: a collection of papers in honour of Robert Burns Woodward, 1917–1979. Pergamon Press, Oxford, UK, 1982. ISBN 0-08-029238-0. xx + 522 pp. LCCN QD255 .R23 1982.

Cornish-Bowden:1997:NBO

[683] Athel Cornish-Bowden, editor. New beer in an old bottle: Eduard Buchner and the growth of biochemical knowledge. Coll̈ecció oberta. Universitat de València, Valéncia, Spain, 1997. ISBN 84-370-3328-4. 252 pp. LCCN ????

URL http://www.gbv.de/dms/bs/toc/318369354.pdf.

Knight:1998:MCS

[684] David M. Knight and Helge Kragh, editors. The making of the chemist: the social history of chemistry in Europe, 1789–1914. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-58351-9 (hardcover). xxi + 353 pp. LCCN QD39.7 .M34 1998. URL http://

www.loc.gov/catdir/description/cam028/97043009.html;   http:
//www.loc.gov/catdir/toc/cam021/97043009.html.

Bhushan:2000:MMN

[685] Nalini Bhushan and Stuart M. Rosenfeld, editors. Of minds and molecules: new philosophical perspectives on chemistry. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-512834-6. xvi

+ 299 pp. LCCN QD6 .O34 2000. URL http://www.loc.gov/catdir/
enhancements/fy0604/99040329-d.html.

Lundgren:2000:CCT

[686] Anders Lundgren and Bernadette Bensaude-Vincent, editors. Communicating chemistry: textbooks and their audiences, 1789–1939, volume 3 of European studies in science history and the arts. Science History Publications/USA, Canton, MA, 2000. ISBN 0-88135-274-8. vii + 465

pp. LCCN QD8.5 .C66 2000. URL http://www.loc.gov/catdir/toc/
fy032/99052211.html.

Benfey:2001:RBW

[687] O. Theodor (Otto Theodor) Benfey and Peter John Turnbull Morris, editors. Robert Burns Woodward: architect and artist in the world of molecules. History of modern chemical sciences. Chemical Heritage Foundation, Philadelphia, PA, USA, 2001. ISBN 0-941901-25-4. ISSN 1069-2452. xxvii + 470 pp. LCCN QD22.W67 R63 2001. URL http://www.

loc.gov/catdir/enhancements/fy0630/00050876-d.html.

Klabunde:2001:NMC

[688] Kenneth J. Klabunde, editor. Nanoscale materials in chemistry. Wiley-Interscience, New York, NY, USA, 2001. ISBN 0-471-38395-3. xi + 292 pp. LCCN TA418.9.N35 N345 2001. URL http:/

/www.loc.gov/catdir/bios/wiley043/2001022371.html;   http://
www.loc.gov/catdir/description/wiley036/2001022371.html;   http:
//www.loc.gov/catdir/toc/onix07/2001022371.html.

Ruthenberg:2008:SNC

[689] Klaus Ruthenberg and J. van (Jaap) Brakel, editors. Stuff: the nature of chemical substances. Königshausen and Neumann, Würzburg, Germany, 2008. ISBN 3-8260-3704-9 (paperback). LCCN QD39 .S78 2008.

Scerri:2008:CPP

[690] Eric R. Scerri, editor. Collected papers on philosophy of chemistry. Imperial College Press, London, UK, 2008. ISBN 1-84816-137-9. x + 235 pp.

LCCN QD6 .S34 2008. URL http://www.loc.gov/catdir/toc/fy0903/
2008278137.html.

Scerri:2009:SPP

[691] Eric R. Scerri, editor. Selected papers on the periodic table. Imperial College Press, London, UK, 2009. ISBN 1-84816-425-4. ix + 143 pp. LCCN QD467 .S347 2009.

Gavroglou:2012:NPN

[692] Kostas Gavroglou and Ana Simões. Neither physics nor chemistry: a history of quantum chemistry. Transformations: studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2012. ISBN 0-262-01618-4 (hardcover). xiv + 351 pp. LCCN QD462 .G38 2012.

Lagerkvist:2012:PTM

[693] Ulf Lagerkvist, editor. The periodic table and a missed Nobel Prize. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2012. ISBN 981-4295-95-7 (paperback), 981-4295-96-5 (e-book). xii + 122 pp. LCCN AS911.N9 .L34 2012; QD11 .L34 2012; QD467

.L34 2012. URL http://www.worldscientific.com/worldscibooks/

10.1142/7658. Edited by Erling Norrby.

Scerri:2014:SEM

[694] Eric R. Scerri, editor. 30-second elements: the 50 most significant elements, each explained in half a minute. Pier 9, Crows Nest, NSW, Australia, 2014. ISBN 1-74336-084-3. 160 pp. LCCN ????